### v5

# Conformal dynamics at infinity for groups with contracting elements

Abstract.
This paper develops a theory of conformal density at infinity for groups with contracting elements. We start by introducing a class of convergence boundary encompassing many known hyperbolic-like boundaries, on which a detailed study of conical points and Myrberg points is carried out. The basic theory of conformal density is then established on the convergence boundary, including the Sullivan shadow lemma and a Hopf–Tsuji–Sullivan dichotomy. This gives a unification of the theory of conformal density on the Gromov and Floyd boundary for (relatively) hyperbolic groups, the visual boundary for rank-1 CAT(0) groups, and Thurston boundary for mapping class groups. Besides that, the conformal density on the horofunction boundary provides a new important example of our general theory. Applications include the identification of Poisson boundary of random walks, the co-growth problem of divergent groups, measure theoretical results for CAT(0) groups and mapping class groups.
Key words and phrases:
Patterson–Sullivan measures, contracting elements, Myrberg points, horofunction boundary, ergodicity
1991 Mathematics Subject Classification:
20F65,20F67,37D40,37B05,20F69,30F60
Partially supported by National Key R & D Program of China (SQ2020YFA070059) and National Natural Science Foundation of China (No. 12131009 and No. 12326601)
1. Introduction
Suppose that a group $G$ admits a proper and isometric action on a proper geodesic metric space $(\mathrm{Y},d)$. The group $G$ is assumed to be non-elementary: there is no finite index subgroup isomorphic to the integer group $\mathbb{Z}$ or the trivial group. This paper continues the investigation of such group actions with a contracting element from the point of dynamics at infinity. This provides a complementary view to several studies of growth problems carried out in [Yan19, Yan20, HY21, GY22].
The contracting property captures the key feature of quasi-geodesics in Gromov hyperbolic spaces, rank-1 geodesics in CAT(0) spaces, and thick geodesics in Teichmüller spaces, and many others. In recent years, this notion and its variants have been proven fruitful in the setup of general metric spaces.
Let $X$ be a closed subset of $\mathrm{Y}$, and $\pi_{X}:\mathrm{Y}\to X$ be the shortest projection (set-valued) map. We say that $X$ is $C$-contracting for $C\geq 0$ if
$\|\pi_{X}(B)\|\leq C$
[…]
  • hyperbolic elements in the relatively hyperbolic groups, cf. [GP16], [GP13];
  • rank-1 elements in proper CAT(0) spaces, cf. [Bal95], [BF09];
  • pseudo-Anosov elements in Teichmüller spaces with Teichmüller metric, cf. [Min96].
The first goal of this paper is aiming to develop a boundary theory for a proper action with contracting elements. In particular, our study is motivated by and attempts to axiomize the corresponding theory of the following
  • Gromov boundary of hyperbolic spaces;
  • Floyd boundary and ends boundary for locally finite graphs;
  • Visual boundary of proper CAT(0) spaces;
  • Thurston boundary, and Gardiner–Masur boundary of Teichmüller spaces.
One of the basic findings in the paper is that these boundaries are in a certain sense, quotients or of its variants, of the horofunction boundary of general metric spaces.
Our results comprise the following three parts.
In the first part, we define a class of compactification for metric spaces, called convergence boundary, reflecting the convergence behaviour of the contracting subsets to infinity. The main novelty is that we develop the notion of conical points and of more intrinsic Myrberg points in line with the corresponding notions in the theory of convergence groups.
The second part studies the conformal measures on the convergence compactifications via the (partial) shadow lemma. We prove a Hopf–Tsuji–Sullivan theorem, setting up the equivalences between the divergence action, and positive/full conformal measure supported on the Myrberg set.
The last part is devoted to applications to growth problems, random walks, measure vanishing results in CAT(0) spaces, and Thurston boundary of mapping class groups.
We start by introducing the so-called convergence compactification.
1.1. Convergence compactification
A bordification of a metric space $\mathrm{Y}$ is a metrizable Hausdorff topological space $\overline{\mathrm{Y}}$ in which $\mathrm{Y}$ embeds as an open and dense subset, and the boundary is $\partial{\mathrm{Y}}:=\overline{\mathrm{Y}}\setminus\mathrm{Y}$. We say that a sequence of subsets $X_{n}$ is escaping if $d(o,X_{n})\to\infty$ for some (thus any) basepoint $o\in\mathrm{Y}$, and converges to a point $\xi\in\partial{\mathrm{Y}}$ if any sequence of points $x_{n}\in X_{n}$ converges to $\xi$.
We are interested in a compactification of $\mathrm{Y}$: $\overline{\mathrm{Y}}$ and $\partial{\mathrm{Y}}$ are compact, though the language of bordification is convenient, as $\mathrm{Y}$ is bordified by any subset of the boundary.
A fundamental notion in this work is the so-called convergence bordification $\overline{\mathrm{Y}}$ so that the following two assumptions are true:
  1. (A)
    Any contracting geodesic ray $X$ converges to a boundary point $\xi\in\partial{\mathrm{Y}}$, and any sequence of $y_{n}\in\mathrm{Y}$ with escaping projections $\pi_{X}(y_{n})$ tends to the same $\xi$.
  2. (B)
    Let $X_{n}$ be any escaping sequence of $C$-contracting subsets. Then for any $x\in\mathrm{Y}$, a subsequence of their cones defined as follows
    $\Omega_{x}(X_{n}):=\{y\in\mathrm{Y}:\|[x,y]\cap N_{C}(X_{n})\|\geq 10C\}$
    converges to a boundary point in $\partial{\mathrm{Y}}$.
By definition, the convergence bordification persists under taking Hausdorff quotient: if we have a quotient map of $\partial{\mathrm{Y}}$ onto a Hausdorff metrizable space, denoted by []delimited-[][$\cdot$][ ⋅ ] as follows
$[\cdot]:\quad\partial{\mathrm{Y}}\to[\partial{\mathrm{Y}}]$
extending the identification $\mathrm{Y}\to\mathrm{Y}$, then $[\partial{\mathrm{Y}}]$ is a convergence bordification of $\mathrm{Y}$. As one-point compactification is always a convergence bordification, we say that a convergence bordification is nontrivial if the following assumption holds
  1. (C)
    the set $\mathcal{C}$ of non-pinched points $\xi\in\partial{\mathrm{Y}}$ is non-empty: if $x_{n},y_{n}\in\mathrm{Y}$ converge to $\xi$, then $[x_{n},y_{n}]$ is an escaping sequence.
See a schematic illustration of Assumptions (A) and (B) in Fig. 1.
Figure 1. Assumption A (left) and Assumption B (right)
Let us keep in mind the following examples, which motivate and satisfy the above two assumptions:
Examples 1.
  1. (1)
    Gromov boundary $\partial_{\mathrm{G}}\mathrm{Y}$ of a (possibly non-proper) Gromov hyperbolic space $\mathrm{Y}$ gives a convergence bordification, where all boundary points are non-pinched.
  2. (2)
    Bowditch boundary $\partial_{\mathrm{B}}\mathrm{Y}$ of (the Cayley graph of) a relatively hyperbolic group $\mathrm{Y}$ gives a convergence compactification ([Bow12, Ger12]). Non-pinched points are exactly conical points, while certain bi-infinite geodesics exist with two half-rays tending to the same parabolic point.
  3. (3)
    Floyd boundary $\partial_{\mathrm{F}}\mathrm{Y}$ of any locally finite graph $\mathrm{Y}$ forms a convergence compactification ([Flo80, Ger12]). Any pinched point admits a horocycle, so non-pinched points contain at least all conical points. See Lemma 10.10.
    The same holds for the end boundary of a locally finite graph (Corollary 10.11).
  4. (4)
    Visual boundary $\partial_{\mathrm{Vis}}\mathrm{Y}$ of a proper CAT(0) space $\mathrm{Y}$ is a convergence boundary, where all boundary points are non-pinched. See Lemma 11.1.
In spite of these examples, one main purpose of our investigation is to setup a unified theory for several boundaries of Teichmüller spaces (Thurston boundary and Gardiner–Masur boundary, etc), Roller boundary of CAT(0) cube complexes, and more generally, the horofunction boundary of any proper metric space. For boundaries of Teichmüller spaces, a typical phenomenon is that a geodesic ray may not admit a unique limit point in the boundary ([Len08]), but generic ones do have ([Mas82a]). Moreover, in pursuing the common feature, we found morally that all these boundaries have much like the horofunction boundary. In the end, we are able to formulate the above assumptions (A) and (B) for these boundaries, modulo a certain ambiguity of the convergence described as follows.
To absorb the pathological convergence, we endow an $\textrm{Isom}(\mathrm{Y})$-invariant partition []delimited-[][$\cdot$][ ⋅ ] on the boundary. This might be artificial at the first sight, but partitioning boundary could be found quite naturally in the study of boundary comparison. For instance, establishing the topological comparison map between Floyd boundary, Bowditch boundary and Martin boundary has generated many applications in the study of relatively hyperbolic groups and random walks ([Ger12, GP13, GP15, GGPY21]). Here we shall go one step further to allow that the corresponding quotient space might be non-Hausdorff.
Given such a partition []delimited-[][$\cdot$][ ⋅ ], the bordification $\overline{\mathrm{Y}}=\mathrm{Y}\cup\partial{\mathrm{Y}}$ is called a convergence bordification if the convergence to a point $\xi$ in the the assumptions (A),(B) and (C) is replaced by the convergence to some point in its equivalent class $[\xi]$. The above Examples 1 of convergence boundaries have been equipped with maximal partitions (i.e. $[\xi]$ are singletons). See Assump A, Assump B, Assump C in §3 for precise formulations.
From a different point of view, we could understand the convergence in the assumptions (A), (B) and (C) after passing to the quotient of $\partial{\mathrm{Y}}$ by identifying points in the same []delimited-[][$\cdot$][ ⋅ ]-class
$:\quad\partial{\mathrm{Y}}$ $\longrightarrow[\partial{\mathrm{Y}}]:=\partial{\mathrm{Y}}/[\sim]$
$\xi$ $\longmapsto[\xi]$
As []delimited-[][$\cdot$][ ⋅ ] may not necessarily form a closed relation, the quotient $[\partial{\mathrm{Y}}]$ might be non-Hausdorff or even not $T_{1}$ (that is, []delimited-[][$\cdot$][ ⋅ ]-classes are not necessarily closed subsets). We say that the partition []delimited-[][$\cdot$][ ⋅ ] is maximal if every []delimited-[][$\cdot$][ ⋅ ]-class is singleton, and it is trivial if $\partial{\mathrm{Y}}$ is a []delimited-[][$\cdot$][ ⋅ ]-class. We shall refer to $[\partial{\mathrm{Y}}]$ as the reduced convergence boundary.
Introducing the partition is indispensable in obtaining a hyperbolic-like boundary, as certain natural subsets in the boundary are distinguishable in a view of negative curvature. The following examples illustrate some aspects of such consideration.
Examples 2.
Let $\mathcal{T}_{g}$ be the Teichmüller space of a closed orientable surface of genus $g\geq 2$, endowed with the following two boundaries (recalled in §12.1 in details):
  1. (1)
    Thurston boundary $\partial_{\mathrm{Th}}\mathcal{T}_{g}$: it is the set $\mathscr{PMF}$ of projective measured foliations, which is a topological sphere of dimension $76g-$ 7 ([FLP79]). Kaimanovich–Masur [KM96] considered a partition []delimited-[][$\cdot$][ ⋅ ] on $\mathscr{PMF}$ which is induced by an intersection form. The restriction to the subset $\mathscr{MIN}$ of the minimal measured foliations is closed. Its quotient, the (Hausdorff) ending lamination space, can be identified with the Gromov boundary of the complex of curves on $\Sigma_{g}$ ([Kla99, Ham06]). Inside $\mathscr{MIN}$, the partition is maximal exactly on the subset $\mathscr{UE}$ of uniquely ergodic points: their equivalent classes are singletons.
  2. (2)
    Gardiner–Masur boundary $\partial_{\mathrm{GM}}\mathcal{T}_{g}$ ([GM91]): it is homeomorphic to the horofunction boundary of $\mathcal{T}_{g}$ with respect to the Teichmüller metric ([LS14]). It turned out that $\mathscr{PMF}$ embeds as a proper subset (with a different subspace topology) in $\partial_{\mathrm{GM}}\mathcal{T}_{g}$. The relation between $\partial_{\mathrm{Th}}\mathcal{T}_{g}$ and $\partial_{\mathrm{GM}}\mathcal{T}_{g}$ has been carefully investigated in [Miy08, Miy13, Miy14a, Wal19].
Examples 3.
The horofunction boundary $\partial_{\mathrm{H}}{\mathrm{Y}}$ was introduced by Gromov for any metric space $\mathrm{Y}$. Two horofunctions are said to be equivalent if they differ by a bounded amount. This defines a finite difference relation on $\partial_{\mathrm{H}}{\mathrm{Y}}$, and the quotient $[\partial_{\mathrm{H}}{\mathrm{Y}}]$ is usually called reduced horofunction boundary. Here are three examples we are interested in.
  1. (1)
    If $\mathrm{Y}$ is a hyperbolic space, the finite difference relation can be shown closed, for two horofunctions in a []delimited-[][$\cdot$][ ⋅ ]-class have a uniformly bounded difference. Thus, the quotient of the horofunction boundary is a Hausdorff space, which is homeomorphic to the Gromov boundary. A non-trivial finite difference relation arises naturally when taking a direct product of a hyperbolic graph with an interval, though the Gromov boundary is kept same.
  2. (2)
    Roller boundary of a locally finite CAT(0) cubical complex is homeomorphic to the horofunction boundary of $\mathrm{Y}$ with combinatorial metric ([FLM18]). Every boundary point is represented by a geodesic ray and the finite difference of horofunctions is equivalent to the finite symmetric difference of boundary points viewed as ultrafilters. The quotient is homeomorphic to the combinatorial boundary studied in [Gen20]. See Lemma 11.5.
  3. (3)
    Martin boundary of a finitely supported random walk on a non-amenable group is homeomorphic to the horofunction boundary defined using (generally non-geodesic) Green metric. The points on which the finite difference relation restricts to be a maximal partition are minimal harmonic functions. The resulting minimal Martin boundary is essential in representing all positive harmonic functions.
Our first result confirms that the horofunction boundary with the finite difference partition provides a nontrivial convergence compactification. The proof is presented in §5.
Theorem 1.1.
The horofunction boundary $\partial_{\mathrm{H}}{\mathrm{Y}}$ of a proper geodesic metric space $\mathrm{Y}$ with contracting subsets is a convergence boundary, with finite difference relation, so that all boundary points are non-pinched.
Gardiner–Masur boundary and Roller boundary are the horofunction boundary of Teichmüller spaces and of CAT(0) cube complexes endowed with Teichmüller metric and combinatorial metric respectively. Hence, they provide examples of convergence boundary with finite difference relation.
Our next result says that Thurston boundary is also a convergence boundary for Teichmüller space endowed with Teichmüller metric. By a result of [Wal19], Thurston boundary is actually the horofunction boundary of the Teichmüller space endowed with Thurston’s Lipschitz metric. The following result proved in Theorem 12.5 is thus not derived as a corollary from Theorem 1.1.
Theorem 1.2.
The Thurston boundary $\partial_{\mathrm{Th}}\mathcal{T}_{g}$ is a convergence boundary, with respect to the above Kaimanovich–Masur partition, so that the set of non-pinched points contains the set of uniquely ergodic points.
As fore-mentioned in Examples 3, the reduced horofunction boundary of a hyperbolic space is metrizable. In general, the reduced $[\partial_{\mathrm{H}}{\mathrm{Y}}]$ of a metric space may be neither Hausdorff nor second countable. This leads to technical issues in the topological and measure theoretical applications using convergence boundary.
A key observation in fixing the issues is that, if we consider a non-elementary group $G<\textrm{Isom}(\mathrm{Y},d)$ so that
(1) $G\text{ acts properly on a proper geodesic space }(\mathrm{Y},d)\text{ with a contracting element}$
then a smaller subset of the boundary consisting of the so-called Myrberg points would have better properties being Hausdorff, second countable and metrizable on the quotient space.
If $G\curvearrowright\mathrm{Y}$ is of divergent type, we shall establish in §1.3 that Myrberg points are generic in conformal measure, and furthermore, the push-forward measures to the quotient become unique and ergodic in §8.
Before moving on the measurable theoretical applications, let us begin a topological investigation of Myrberg points and their relatives conical points.
1.2. Conical and Myrberg limit sets
Conical points are an important sub-class of limit points, which have been well-studied in Kleinian groups and in a more general class of convergence group actions (see Section 10). Myrberg points form a strictly smaller subset of conical points, on which the dynamics is quite powerful, but seems to be under-estimated in literature. The main novelty here is to formulate the notion of Myrberg and conical limit points for a general proper action as in (1), which shares desirable properties as much as in a convergence group action (see Subsection 10.1).
Dynamics on the limit set
We start discussing the notion of limit sets for any group $G\curvearrowright\mathrm{Y}$ as in (1).
A main consequence of the convergence boundary allows us to develop a “good” theory of limit sets in a sense that the limit set is the unique and minimal invariant closed subset, up to taking []delimited-[][$\cdot$][ ⋅ ]-closure.
Define the limit set $\Lambda(Go)$ to be the set of accumulation points of a $G$-orbit $Go$ in $\partial{\mathrm{Y}}$. It may depend on the choice of $o\in\mathrm{Y}$ but the []delimited-[][$\cdot$][ ⋅ ]-closure $[\Lambda(Go)]$ does not by Lemma 3.8. By Assump A, if $h$ is a contracting element, the positive and negative half rays (i.e. $n>0$ and $n<0$) of its quasi-geodesic
$n\in\mathbb{Z}\longmapsto h^{n}o\in\mathrm{Y}$
determine two []delimited-[][$\cdot$][ ⋅ ]-classes of boundary points denoted by $[h^{+}],[h^{-}]$. The union $[h^{\pm}]=[h^{-}]\cup[h^{+}]$ are the []delimited-[][$\cdot$][ ⋅ ]-closure of the limit set of $\langle h\rangle\cdot o$. We say that $h$ is non-pinched if its fixed points $[h^{\pm}]$ are non-pinched in Assump C. Equivalently, this is amount to asking $[h^{-}]\neq[h^{+}]$ by Lemma 3.19. In a discrete group, any two non-pinched contracting elements have either the same fixed points or disjoint fixed points (Lemma 3.21).
In what follows, we always assume that the convergence boundary $\partial{\mathrm{Y}}$ under consideration is non-elementary for the action $G\curvearrowright\mathrm{Y}$ in (1). That is, $G$ contains a non-pinched contracting element, whose fixed points are distinct []delimited-[][$\cdot$][ ⋅ ]-classes in $\partial{\mathrm{Y}}$.
Lemma 1.3 (NS dynamics, see Lemma 3.27).
Let $h\in\textrm{Isom}(\mathrm{Y})$ be a non-pinched contracting isometry. Then, under Assump A and Assump C, the action of $\langle h\rangle$ on $\partial{\mathrm{Y}}\setminus[h^{\pm}]$ has the North–South dynamics:
  • for any two open sets $[h^{+}]\subseteq U$ and $[h^{-}]\subseteq V$ in $\partial{\mathrm{Y}}$, there exists an integer $n>0$ such that $h^{n}(\partial{\mathrm{Y}}\setminus V)\subseteq U$ and $h^{-n}(\partial{\mathrm{Y}}\setminus U)\subseteq V$.
Let us formulate the main conclusion we obtained for the limit set.
Lemma 1.4 (see Lemma 3.30).
Let $h\in G$ be a non-pinched contracting element with minimal fixed points $[h^{-}]=\{h^{-}\},[h^{+}]=\{h^{+}\}$ on a convergence boundary $\partial{\mathrm{Y}}$. Let $\Lambda$ be the closure of the fixed points of all conjugates of $h$ in $G$. Then
  1. (1)
    $\Lambda$ is the unique, minimal, and $G$-invariant closed subset.
  2. (2)
    $[\Lambda]=[\Lambda(Go)]$ for any $o\in\mathrm{Y}$.
  3. (3)
    The set of the fixed point pairs of all non-pinched contracting elements in $G$ is dense in the distinct pairs of $\Lambda\times\Lambda$.
All the boundaries in Examples (1, 2 and 3), except the horofunction boundary of a general metric space, contain a non-pinched contracting element with minimal fixed points. The list of properties as above are mostly known in these boundaries, but seem to be new for Roller and Gardiner–Masur boundaries, see Lemmas 11.5 and 12.6 for concrete formulations.
Myrberg points
We first introduce the notion of Myrberg points, which is more canonical than that of conical points. To that end, let us denote the set of distinct pairs of []delimited-[][$\cdot$][ ⋅ ]-classes:
(2) $\Lambda G\Join\Lambda G:=\{([\xi],[\eta])\in[\Lambda(Go)]\times[\Lambda(Go)]:[\xi]\neq[\eta]\}$
Recall that $\mathcal{C}$ is the subset of non-pinched boundary points in Assump C.
Definition 1.5 (Myrberg points).
A point $\xi\in\mathcal{C}\subseteq\partial{\mathrm{Y}}$ is called Myrberg point if for any $x\in\mathrm{Y}$, the set of $G$-translates of the ordered pair $(x,\xi)$ is dense in $\Lambda G\Join\Lambda G$ in the following sense:
  • For any $[\zeta]\neq[\eta]\in[\Lambda(Go)]$ there exists $g_{n}\in G$ so that $g_{n}x\to[\zeta]$ and $g_{n}\xi\to[\eta]$.
Remark 1.6.
Introduced by Myrberg [Myr31], this class of limit points was proved to be full Lebesgue measure in the ideal boundary $\mathbb{S}^{1}$ for finite co-volume Fuchsian groups. This was generalized later on to higher dimension [Aga83, Tuk94b].
In contrast with the reduced convergence boundary $[\partial{\mathrm{Y}}]$, the Myrberg subset enjoys quite good topological properties after passing to the quotient.
Lemma 1.7 (see Lemma 4.22).
The quotient $[\Lambda_{m}{(Go)}]$ of the Myrberg set in a non-elementary convergence boundary $\partial{\mathrm{Y}}$ is a Hausdorff, second countable, metrizable topological space.
Moreover, $\mathrm{Y}\cup[\Lambda_{m}{(Go)}]$ forms a convergence bordification, endowed with maximal partition, so that all boundary points are non-pinched.
If $\partial{\mathrm{Y}}$ is the horofunction boundary we prove in Lemma 5.7, via different means, that $[\Lambda_{m}{(Go)}]$ is a Hausdorff and second countable topological space. Lemma 1.7 is proved in Subsection §4.5 by embedding $[\Lambda_{m}{(Go)}]$ into the Gromov boundary of a quasi-tree of spaces. The above properties are therefore inherited from those of the Gromov boundary.
Conical points
Our notion of conical points is defined relative to a collection $\mathscr{F}$ of contracting subsets with bounded intersection. There are various (quantitatively) equivalent ways to define it. We give a quick definition in Introduction, and the other two elaborating ones are discussed in Section §4.
Let $F$ be a set of three (mutually) independent non-pinched contracting elements $f_{i}$ ($i=1,2,3$), which form a contracting system as follows
(3) =$\mathscr{F}=\{g\cdot\textrm{Ax}(f_{i}):g\in G\}$
where the axis $\textrm{Ax}(f_{i})$ defined in (11) depends on the choice of a basepoint $o\in\mathrm{Y}$.
One consequence of Assump A and Assump C together is as follows: the shortest projection map $\pi_{X}:\mathrm{Y}\to X$ to a contracting bi-infinite quasi-geodesic $X$ extends to the subset $\mathcal{C}$ of non-pinched boundary points (§3.4):
$\pi_{X}:\mathrm{Y}\cup\mathcal{C}\setminus[\Lambda X]\longrightarrow X$
which satisfies the bounded geodesic image property (for possibly larger $C$):
$d_{X}(x,y)\geq C\quad\Longrightarrow\quad[x,y]\cap N_{C}(X)\neq\emptyset$
where $d_{X}(x,y):=\|\pi_{X}(x)\cup\pi_{X}(y)\|$. See Lemma 3.22.
Definition 1.8 (Conical points, see Definition 4.5).
A non-pinched point $\xi\in\mathcal{C}$ is called $L$-conical point for $L>0$ relative to $\mathscr{F}$ if for some $x\in\mathrm{Y}$, there exists a sequence of $X_{n}\in\mathscr{F}$ such that $d_{X_{n}}(x,\xi)\geq L$. Denote by $\Lambda_{L}^{\mathscr{F}}(Go)$ the set of all such points.
An important ingredient in our argument is the axiomized construction of a projection complex $\mathcal{P}_{K}(\mathscr{F})$ from $\mathscr{F}$, developed by Bestvina-Bromberg-Fujiwara [BBF15]. Loosely speaking, fixing a constant $K$, $\mathcal{P}_{K}(\mathscr{F})$ is a graph with vertex set $\mathscr{F}$ so that two vertices $U\neq V\in\mathscr{F}$ are adjacent if and only if they are visually $K$-small from any third party $W\in\mathscr{F}\setminus\{U,V\}$:
$d_{W}(U,V)\leq K$
(see §2.6 for more details). The basic fact is that for any $K\gg 0$, the graph $\mathcal{P}_{K}(\mathscr{F})$ is a unbounded quasi-tree (with universal hyperbolicity constant).
Our next result, which is proved in Lemma 4.19, describes the above-defined conical points via the familiar Gromov boundary of the projection complex.
Lemma 1.9.
For any $L\gg 0$, there exists $K>0$ such that the Gromov boundary of the projection complex $\mathcal{P}_{K}(\mathscr{F})$ admits a topological embedding into the set of $L$-conical points relative to $\mathscr{F}$.
In [BBF15], a quasi-tree of spaces denoted by $\mathcal{C}(\mathscr{F})$ is built from the projection complex $\mathcal{P}_{K}(\mathscr{F})$ in the following way. Keeping the edge incidence, $\mathcal{C}(\mathscr{F})$ is obtained from $\mathcal{P}_{K}(\mathscr{F})$ by replacing each vertex with the corresponding subspace in $\mathrm{Y}$. The technical advantage is that the fixed points of $f\in F$ are incorporated into the Gromov boundary of $\mathcal{C}(\mathscr{F})$ (but not in $\mathcal{P}_{K}(\mathscr{F}))$. Consequently, we are able to embed the conical points $\Lambda_{r}^{F}(Go)$ into the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$ in Lemma 4.21. In particular, Myrberg points are thus embedded in the Gromov boundary as a subset of a separable metrizable space. This proves Lemma 1.7.
We close the discussion with a connection between Myrberg and conical points. Namely, Corollary 4.17 sets up an equivalent characterization of Myrberg points in terms of conical points.
Proposition 1.10.
The Myrberg set $\Lambda_{m}{(Go)}$ is the (countable) intersection of all $(L,\mathscr{F})$-conical sets $\Lambda_{L}^{\mathscr{F}}(Go)$, over all possible $\mathscr{F}$ given by (3) and all large integers $L\in\mathbb{N}$.
This result admits several applications, which shall be delivered in the forth-coming work. At the moment, let us mention the following the topological one (with measurable one in next subsection) concerning Bowditch boundary of relatively hyperbolic groups.
Theorem 1.11 (see Theorem 10.21).
Let $(G,\mathcal{P})$ be a relatively hyperbolic pair and $\partial_{\mathrm{B}}(G,\mathcal{P})$ be the associated Bowditch boundary. Then the Myrberg set for the action on $\partial_{\mathrm{B}}(G,\mathcal{P})$ is homeomorphic to the Myrberg set for the action the Floyd boundary.
We refer the reader to §10.3 for a brief introduction to relatively hyperbolic groups.
A finitely generated group may admit different peripheral structure $\mathcal{P}$ to be relatively hyperbolic (see [Yan14b] for their relations). Hence, this result says that the Myrberg set is invariant up to changing the peripheral structure on $G$. This could be compared with Floyd mapping theorem ([Flo80, Ger12]). A similar conclusion holds for Cannon–Thurston maps; see Theorem 10.23 for precise statement.
1.3. Conformal density on the convergence boundary
A family of conformal measures on the limit set of Fuchsian groups was firstly constructed by Patterson [Pat76] and studied extensively by Sullivan [Sul79] in higher dimensional Kleinian groups. Patterson’s construction is very robust and carries over verbatim to the general metric spaces with horofunction boundary (see [BM96]). The second part of our study is to give a vast generalization of these works in groups with contracting elements, endowed with a convergence compatification. Let us first introduce the setup for the Patterson’s construction.
Fix a basepoint $o\in\mathrm{Y}$. Consider the growth function of the ball of radius $R>0$:
$N(o,R):=\{v\in Go:d(o,v)\leq n\}.$
The critical exponent $\omega_{\Gamma}$ for a subset $\Gamma\subseteq G$:
(4) $\omega_{\Gamma}=\limsup\limits_{R\to\infty}\frac{\log\sharp(N(o,R)\cap\Gamma o)}{R},$
is independent of the choice of $o\in\mathrm{Y}$, and intimately related to the Poincaré series
(5) $s\geq 0,x,y\in\mathrm{Y},\quad\mathcal{P}_{\Gamma}(s,x,y)=\sum\limits_{g\in\Gamma}e^{-sd(x,gy)}$
as $\mathcal{P}_{\Gamma}(s,x,y)$ diverges for $s<\omega_{\Gamma}$ and converges for $s>\omega_{\Gamma}$. Thus, the action $G\curvearrowright\mathrm{Y}$ as in (1) is called of divergent type (resp. convergent type) if $\mathcal{P}_{G}(s,x,y)$ is divergent (resp. convergent) at $s=\omega_{G}$.
Consider a further assumption, Assump D, on a convergence compactification $\overline{\mathrm{Y}}=\mathrm{Y}\cup\partial{\mathrm{Y}}$, so that for non-pinched points $\xi\in\mathcal{C}$, as $z\to\xi,$ the Buseman cocycle
$\forall x,y,z\in\mathrm{Y}:\;B_{z}(x,y):=d(x,z)-d(y,z)$
extends coarsely to a well-defined Buseman quasi-cocycle $B_{\xi}(x,y)$. See Assump D for precise formulation and Remark 6.2 for further clarification.
Denote by $\mathcal{M}_{+}(\overline{\mathrm{Y}})$ the set of positive Radon measures on $\overline{\mathrm{Y}}$. If $G\curvearrowright\mathrm{Y}$ is of divergent type, the Patterson–Sullivan measures $\{\mu_{x}\}_{x\in\mathrm{Y}}$ are accumulation points in $\mathcal{M}_{+}(\overline{\mathrm{Y}})$ of the following family of measures $\{\mu_{x}^{s}\}_{x\in\mathrm{Y}}$ supported on $Go$
$\mu_{x}^{s}=\frac{1}{\mathcal{P}_{G}(s,o,o)}\sum\limits_{g\in G}e^{-sd(x,go)}\cdot{\mbox{Dirac}}{(go)},$
as $s>\omega_{G}$ tends to $\omega_{G}$. Similar construction works for groups of convergent type via a Patterson’s trick [Pat76] on the Poincaré series. The $G$-equivariance of such measures $\mu_{x}$ on basepoints $x$ and their conformality on switching basepoints are formulated in a general notion of (quasi)-conformal density (see Definition 6.1).
In [Sul79], Sullivan proved a shadow lemma describing the local nature of Patterson’s measure on the limit set of Kleinian groups. We extend it to our setting. Define the usual shadow $\Pi_{x}(y,r)$ as the topological closure in $\partial{\mathrm{Y}}$ of the cone as follows:
$\Omega_{x}(y,r)=\{z\in\mathrm{Y}:[x,z]\cap B(y,r)\neq\emptyset\}$
Analogous to the partial cone and shadow in [Yan22], we are actually working with a similar version of shadows via contracting segments. With a set $F$ of three independent contracting elements as in (3) and $r>0$, a segment $\alpha$ in $\mathrm{Y}$ contains an $(r,F)$-barrier at $go$ if for some $f\in F$, we have $go,\;gfo\in N_{r}([x,z]).$ For $x\in\mathrm{Y},y\in Go$ and $r>0$, the $(r,F)$-cone is defined as follows
$\Omega_{x}^{F}(y,r):=\{z\in\mathrm{Y}:y\text{ is an }(r,F)\text{-barrier for some geodesic }[x,z]\}$
whose topological closure in $\partial{\mathrm{Y}}$ gives the $(r,F)$-shadow $\Pi_{x}^{F}(y,r)$. See §3 for details. The following version of Shadow Lemma is proved in Lemma 6.4.
Lemma 1.12.
Let $G\curvearrowright\mathrm{Y}$ be as in (1). Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be a $\omega$-dimensional $G$-quasi-equivariant, quasi-conformal density on $\partial{\mathrm{Y}}$ for some $\omega>0$, charging positive measure to the non-pinched points $\mathcal{C}$. Then there exists $r_{0}>0$ such that
$\begin{array}[]{rl}e^{-\omega\cdot d(o,go)}\;\prec\;\mu_{o}(\Pi_{o}^{F}(go,r)\cap\mathcal{C})\;\leq\;\mu_{o}(\Pi_{o}(go,r)\cap\mathcal{C})\;\prec_{r}\;e^{-\omega\cdot d(o,go)}\\ \end{array}$
for any $g\in G$ and $r\geq r_{0}$, where $\prec_{r}$ means the inequality holds up to a multiplicative constant depending on $r$ (or a universal constant if no $r$ is displayed).
Remark 1.13.
By Theorem 1.1, we have $\mathcal{C}=\partial_{\mathrm{H}}{\mathrm{Y}}$: the complete version of shadow lemma holds in the horofunction boundary. This is the main new instance, see §1.5 for a short history of it.
In order to state the next result, we also introduce the usual version of conical points defined via the usual shadow without involving $\mathscr{F}$:
(6) $\Lambda_{c}(Go):=\bigcup_{r\geq 0}\bigcup_{x\in Go}\left(\limsup_{y\in Go}\Pi_{x}(y,r)\right).$
By definition, $L$-conical points and Myrberg points are conical. Note that every point in the visual boundary is conical in this sense for any geometric action on a CAT(0) space. This is partly the reason why we restrict our attention to the smaller subset of $L$-conical points and Myrberg points.
The main result of this investigation is relating the divergence of Poincaré series to the positive conformal measure charged on conical points. This type of results is usually refereed to as Hopf–Tsuji–Sullivan dichotomy by researchers. Note here that the item on the ergodicity of product measures is missing.
Theorem 1.14.
Let $G\curvearrowright\mathrm{Y}$ be as in (1). Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be a $\omega$-dimensional $G$-quasi-equivariant, quasi-conformal density on a non-elementary convergence boundary $\partial{\mathrm{Y}}$ for some $\omega>0$. Assume that $\mu_{o}(\mathcal{C})>0$. Then the following statements are equivalent: Let $\mathscr{F}$ be as in (3) and any $L\gg 0$.
  1. (1)
    The Poincaré series $\mathcal{P}_{G}(s,o,o)$ diverges at $s=\omega$;
  2. (2)
    The Myrberg set has the positive $\mu_{x}$-measure;
  3. (3)
    The Myrberg set has the full $\mu_{x}$-measure in $\mathcal{C}$;
  4. (4)
    The set of conical points (or $(L,\mathscr{F})$-conical points) has positive $\mu_{x}$-measure;
  5. (5)
    The set of conical points (or $(L,\mathscr{F})$-conical points) has full $\mu_{x}$-measure in $\mathcal{C}$.
If one of the above statements is true, then $\omega=\omega_{G}$.
Remark 1.15.
The theorem is proved via the routine: (1) $\Longleftrightarrow$ (4) $\Longrightarrow$ (5) $\Longrightarrow$ (3) $\Longrightarrow$ (2) $\Longrightarrow$ (4). We comment on the following directions while the others are trivial or classic.
  • (1) $\Longrightarrow$ (4): this is the most difficult direction, whose proof uses the projection complex to build a family of the visual sphere decomposition of group elements. With light source at the basepoint $o\in\mathrm{Y}$, each visual sphere is shadowed by the previous one and blocks the following one. Such a structure allows to re-arrange the series $\mathcal{P}_{G}(s,o,o)$ as a geometric-like series. See §7.3 for more details.
  • (5) $\Longrightarrow$ (3): This follows from Proposition 1.10, as $\Lambda_{m}{(Go)}$ is the countable intersection of $\mu_{o}$-full $(L,\mathscr{F})$-conical points, over all possible $\mathscr{F}$ in (3).
Under any statement in Theorem 1.14, the conformal density turns out to be unique when passed to the quotient of Myrberg set (which is Hausdorff and second countable by Lemma 1.7). This is proved in Theorem 8.5.
Theorem 1.16.
Let $G\curvearrowright\mathrm{Y}$ be as in (1) of divergent type, compactified by a non-elementary convergence boundary $\partial_{\mathrm{H}}{\mathrm{Y}}$. Then the $\omega_{G}$-dimensional $G$-quasi-equivariant, quasi-conformal density on the quotient $[\Lambda_{m}{(Go)}]$ of $\Lambda_{m}{(Go)}\subseteq\partial{\mathrm{Y}}$ is unique up to a bounded multiplicative constant, and ergodic.
By Rohklin’s theory, $[\Lambda_{m}{(Go)}]$ endowed with conformal measures are Lebesgue spaces ([Cou16, Lemma 15.4]), i.e. measurably isomorphic to the interval $[0,1]$ with Lebesgue measure.
1.4. More applications
We now present various applications to the growth problems, and measure theoretical results in CAT(0) groups and mapping class groups.
Cogrowth tightness
We start with the following co-growth results. This extends the work of [MYJ20] about divergent group actions on hyperbolic spaces, and answers positively [AC20, Questions 4.1 and 4.2]. The following is proved in Theorem 9.1.
Theorem 1.17.
Let $G\curvearrowright\mathrm{Y}$ be as in (1). Suppose that the group $G$ is of divergent type. Then for any infinite normal subgroup $H$ of $G$, we have
$\omega_{H}>\frac{\omega_{G}}{2}$
Moreover, if $H$ is of divergent type, then $\omega_{H}=\omega_{G}$.
Remark 1.18.
In [AC20], Arzhantseva-Cashen proved the same result under the assumption that $G\curvearrowright\mathrm{Y}$ has purely exponential growth. This already includes many interesting examples, but the divergent action is weaker than having purely exponential growth.
Myrberg set as Poisson boundary of random walks
We now give an application of previous results to random walks on groups with contracting elements. A probability measure $\mu$ on $G$ is called irreducible if its support generates the group $G$ as a semi-group. A $\mu$-random walk on $G$ is the stochastic process $(w_{n})$ defined as
$w_{n}:=g_{1}g_{2}\dots g_{n}$
where $(g_{n})$ is a sequence of i.i.d. random variables valued in $G$ with law $\mu$. The entropy is defined by
$H(\mu):=-\sum_{g\in\textrm{supp}(\mu)}\mu(g)\cdot|\log\mu(g)|.$
By the work of Maher–Tiozzo [MT18], the following statement holds for the random walk for acylindrical actions on hyperbolic spaces. The previous two results transfer the results to the horofunction boundary. More details on the proof is given in §8.4.
Theorem 1.19.
Let $G\curvearrowright\mathrm{Y}$ as in (1). Consider an irreducible $\mu$-random walk driven by a probability measure $\mu$ on $G$ with finite logarithmic moment on $\mathrm{Y}$:
$\sum_{g\in G}\mu(g)\cdot|\log d(o,go)|<\infty.$
Then almost every trajectory of $\mu$-random walk converges to the []delimited-[][$\cdot$][ ⋅ ]-class of a Myrberg point in $\partial_{\mathrm{H}}{\mathrm{Y}}$.
Moreover, if $H(\mu)<\infty$, then the reduced horofunction boundary $[\partial_{\mathrm{H}}{\mathrm{Y}}]$ is the Poisson boundary, with harmonic measure supported on the quotient $[\Lambda_{m}{(Go)}]$ of Myrberg set.
Remark 1.20.
If $\mathrm{Y}$ is a CAT(0) space with rank-1 elements, this completes the theorem of Karlsson–Margulis [KM99] by showing the positive drift. For the proper essential action on CAT(0) cube complex, our result recovers the main results [FLM18, Thm 1.2] [Fer18] that almost every trajectory tends to a point in the Roller boundary. See Lemma 11.5 for the relevant facts on Roller boundary.
Using the recent work [CFFT22], the finite logarithmic moment condition could be removed in the above statement, still under finite entropy assumption. See the remark 8.8.
Conformal measure on the contracting boundary
Contracting boundary for CAT(0) spaces was introduced by Charney–Sultan [CS15] as quasi-isometric invariant, and has attracted active research interests in recent years. It is observed that the contracting boundary is measurably negligible in harmonic measures. We derive the same result in conformal measures from a more general Theorem 11.6. Here let us state it in CAT(0) spaces.
The underlying set of the contracting boundary consists of the endpoints of contracting geodesic rays in the visual boundary.
Theorem 1.21.
Let $G\curvearrowright\mathrm{Y}$ in (1) be co-compact, where $\mathrm{Y}$ is a proper CAT(0) space. Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be the $\omega_{G}$-dimensional conformal density on the visual boundary $\partial_{\mathrm{Vis}}\mathrm{Y}$. Then the underlying set of contracting boundary is $\mu_{o}$-null if and only if $G$ is not a hyperbolic group.
Hopf–Tsuji–Sullivan Dichotomy for mapping class groups and cubical groups
At last, for subgroups of mapping class groups, we can complete the Hopf–Tsuji–Sullivan Dichotomy 12.4 with ergodicity of product measures.
Theorem 1.22.
Consider any non-elementary subgroup $G$ of $\textrm{Mod}(\Sigma_{g})$ ($g\geq 2$) with pseudo-Anosov elements. Let $\{\mu_{x}\}_{x\in\mathcal{T}_{g}}$ be a $\omega$-dimensional $G$-equivariant conformal density on $\partial_{\mathrm{Th}}\mathcal{T}_{g}=\mathscr{PMF}$. Then the following are equivalent:
  1. (1)
    The Poincaré series $\mathcal{P}_{G}(s,o,o)$ of $G$ diverges at $\omega$;
  2. (2)
    The Myrberg set $\Lambda_{m}{(Go)}$ has either full or positive $\mu_{x}$-measure;
  3. (3)
    The conical set $\Lambda_{c}{(Go)}$ has either full or positive $\mu_{x}$-measure;
  4. (4)
    The diagonal action on $\mathscr{PMF}\times\mathscr{PMF}$ is ergodic with respect to the product measure $\mu_{o}\times\mu_{o}$.
Remark 1.23.
If $G=\textrm{Mod}(\Sigma_{g})$ the uniqueness of $\omega_{G}$-dimensional conformal density was proved [Ham06] [LM08], and the ergodicity was known long before by work of Masur [Mas82a] and Veech [Vee82]. The above statement, however, seems new for non-elementary proper subgroups.
At last, we note the following corollary to the non-elementary action on a CAT(0) cube complex for further reference. Such an action is called essential if no half-space contains an orbit in its fixed finite neighborhood (See [CS11]). In this setup, a Myrbert limit point is minimal, so we do not need take the quotient to obtain the uniqueness of PS measures.
Theorem 1.24.
Let $G\curvearrowright\mathrm{Y}$ in (1) be an essential action of divergent type on a CAT(0) cube complex compactified by the Roller boundary $\partial_{\mathrm{R}}\mathrm{Y}$. Then the $\omega_{G}$-dimensional $G$-quasi-equivariant, quasi-conformal density on $\partial_{\mathrm{R}}\mathrm{Y}$ is unique up to a bounded multiplicative constant, and ergodic.
1.5. Historical remarks and further questions
In a series of works [Yan19, Yan20, HY21, GY22] without using ergodic theory, the elementary and geometric arguments are employed to establish coarse counting results for groups with contracting elements acting on general metric spaces.
Analogous to Anosov shadowing / closing property, a geometric tool called extension lemma was used (recalled here in Lemma 2.17): roughly speaking, any two geodesics $\alpha,\beta$ can be connected by a short arc $c$ to form a quasi-geodesic $\alpha\cdot c\cdot\beta$. The short arc is provided by a contracting segment, whose existence facilities the extensive use of Gromov-hyperbolic geometry. In particular, if $\alpha=\beta$, then a repeated application of extension lemma concatenates copies of $\alpha\cdot c$ to get a periodic quasi-geodesic. In other words, this gives a closed geodesic nearby, with length coarsely equal $\ell(\alpha)$ on the quotient $\mathrm{Y}/G$. On account of our previous works, the complementary view from the ergodic theory is anticipated, and actually forms the primary goal of the present paper.
The conformal density on the limit set of Fuchsian groups was famously constructed by Patterson [Pat76], and further developed by Suillvan [Sul79] in Kleinian groups with a number of applications. Particularly, their work sets the right track for the further generalization of (quasi)-conformal density on boundaries of hyperbolic spaces [Coo93], CAT(-1) spaces [BM96][Rob03], and CAT(0) spaces [Ric17, Lin18]. The Sullivan Shadow Lemma [Sul79] provides a very useful tool in applications, and thus is most desirable in any “good” theory of conformal density. Although the Sullivan’s proof carries over to Gromov hyperbolic spaces, its generalization in CAT(0) manifolds and spaces follows a different argument given by Knieper ([Kne98],[Kne97, Remark on p.781]) and Link [Lin18]. In Teichmüller space, the Shadow lemma is obtained in [TY21] with applications to fundamental inequality of random walks (see [Gek, Lemma 5.1] for a special case).
Attempting to formulate a unified framework for the (quasi-)conformal density in the above examples motivates the investigation of this paper.
In [Yan22], the quasi-conformal density on Floyd and Bowditch boundary of relatively hyperbolic groups is used to study growth problems in word metrics. Like Teichmüller space, the Cayley graph of a relatively hyperbolic group is generally neither Gromov hyperbolic nor CAT(0) spaces. Moreover, the lack of Buseman cocyles at all boundary points forces us to define Buseman (quasi-)cocyles only at conical points. To handle non-conical points, we prove the constructed Patterson–Sullivan measures are fully supported on conical points. This strategy is axiomized in the present paper as the Assump D. As another instance, the conformal density obtained on the Thurston boundary [ABEM12] follows a similar route detailed in §12.3.
Concerning the convergence boundary, it would be interesting to compare with Karlsson’s theory of stars at infinity [Kar05]. Defined using half-spaces, the stars describe some incidental relation on boundary, but do not form a partition in general. A nice feature of his theory allows to formulate an analogue of convergence group property. Some partial convergence property also exists on our conical points but is not included as we do not see an application.
Further questions
We discuss some questions we found interesting about the convergence boundary. The notion of contracting subsets in this paper is usually called strongly contracting in literature: there are many co-existing notions of contracting subsets (see [ACGH17]). We are working with proper geodesic metric spaces for simplicity, but the framework proposed here could be adapted to non-geodesic and non-symmetric metric spaces with strongly contracting elements. Here are two concrete examples in potential applications:
  1. (1)
    Martin boundary of groups with nontrivial Floyd boundary, that is the horofunction boundary of Green metric (see [GGPY21]).
  2. (2)
    Horofunction boundary of outer space for $\textrm{Out}(\mathbb{F}_{n})$ with Lipschitz (asymmetric) metric.
On the other hand, a challenging question would be the following.
Question 1.25.
Does there exist a “good” theory of the convergence boundary for groups with weakly contracting elements / Morse elements, and the corresponding conformal density at infinity?
Boundary comparison is an interesting research topic, and has been well-studied for convergence group actions, see the work of Floyd and Gerasimov [Flo80, Ger12]. Recently, a surjective continuous map from Martin boundary to Floyd boundary was discovered in [GGPY21], where the Martin boundary can be seen as the horofunction boundary of the Green metric. As the dynamical notions of conical points and Myrberg points witness, the convergence boundary shows many similarities with convergence group actions. In view of these results, we are interested in understanding.
Question 1.26.
Which group action on a convergence compactification $G\curvearrowright\overline{\mathrm{Y}}$ admit a quotient with a nontrivial convergence group action?
All known examples with this property have non-trivial Floyd boundary. It is easy to check that the horofunction boundary for word metric surjects to the Floyd boundary (cf. Lemma 10.9). It is still open that whether such groups are necessarily relatively hyperbolic.
The last question concerns about the Myrberg set. We prove in Lemma 10.21 that Myrberg set in Floyd boundary persists in quotients: there exists a homeomorphism from the Myrberg set in the Floyd boundary onto the Myrberg set in any of its quotient with a non-trivial convergence group action. This result seems providing a positive evidence to the following.
Conjecture 1.27.
Suppose a group $G$ acts geometrically on two proper CAT(0) spaces $X$ and $Y$. Then there exists a $G$-equivariant homeomorphism between the corresponding Myrberg sets.
As the two CAT(0) actions share the same set of rank-1 elements, it still has some chance to have a positive answer. Compare with the famous examples of Croke–Kleiner [CK00] and the quasi-isometric invariant boundaries recently studied in [CS15, Cor17, QR22].
At the completion of writing of this paper, the author received a preprint of Gekhtman–Qing–Rafi [GQR] which contains similar results as Theorems 1.19 and 1.21 for sublinearly Morse directions. On this regard, it would be interesting to compare sublinearly Morse boundary with Myrberg set.
Recently, R. Coulon independently developed a Patterson–Sullivan theory on the horofunction boundary in [Cou], in a large overlap with ours, including Shadow lemma 6.4, Theorems 1.14, 1.16, and 1.17. However, our proof of Theorem 1.14 via the projection complex machinery are very different from his approach, and is actually a novelty of our study. The measurable quotient of horofunction boundary in Theorem 1.16 is a Lebesgue space, which is not known in [Cou].
Organization of the paper
The paper is organized into three parts. The first part from Sections §2 – §5 presents the basic theory of convergence boundary. After the preparatory section §2, we introduce the notion of convergence boundary in §3 and derive basic facts for later use. The study of conical points and Myrberg points is carried out in §4, with Proposition 1.10 and Theorem 1.9 proved among others. Theorem 1.1 about convergence of horofunction boundary is shown in §5.
The second part from Sections §6 – §8 develops the theory of conformal density on the convergence boundary. Shadow Lemma 6.4 and Shadow Principle 6.11 are established in §6, and the Hopf–Tsuji–Sullivan Theorem 1.14 in §7. Section §8 contains two different but related results: the unique conformal density Theorem 1.16 for divergent groups, and the reduced Myrberg set as Poisson boundary stated in Theorem 1.19.
The final Sections 9, 10, 11 and 12 collect various applications of this study; in particular, the last three Sections are devoted accordingly to Case Studies of Relatively hyperbolic groups, CAT(0) groups and Mapping class groups. In §9, we establish a co-growth tightness for normal subgroups in a general divergent action, answering questions in [AC20]. Among others, Theorem 1.11 about homeomorphic Myrberg sets is proved in §11 for relatively hyperbolic groups, and Theorem 1.21 on nullity of contracting boundary is derived as a corollary of a general theorem 11.6 for CAT(0) groups in §11, and Hopf–Tsuji-Sullivan Dichotomy 1.22 for mapping class groups is proved in §12.
Acknowledgments
This work started in December 2015, where the author was visiting Université de Lille 1 under a 3-month CNRS research fellowship. He wishes to thank Professor Leonid Potyagailo for the invitation and the hospitality of Math department. During the visit, the proof of shadow lemma was obtained after discussions with Ilya Gekhtman and was explained to Fanny Kassel. The author thanks Ilya for his interests since then, and Rémi Coulon for related conversations in August 2018 and May 2019. Thanks also to Hideki Miyachi for sharing his prints, and Anthony Genevois, Weixu Su and Giulio Tiozzo for many helpful discussions. Part of the writing was completed during the visit in 2022 to the Institute for Advanced Study in Mathematics in Zhejiang University.
2. Preliminary
2.1. Notation and Convention
Let $(\mathrm{Y},d)$ be a proper geodesic metric space. The shortest projection of a point $y\in\mathrm{Y}$ to a closed subset $X\subseteq\mathrm{Y}$ is given by
$\pi_{X}(y)=\{x\in X:d(y,x)=d(y,X)\},$
and of a subset $A\subseteq\mathrm{Y}$, $\pi_{X}(A)=\bigcup_{a\in A}\pi_{X}(a).$ Denote the diameter of a set $A$:
$\|A\|=\sup\{d(x,y):x,y\in A\}$
If $F$ is a finite set of isometries and $o\in\mathrm{Y}$ is a base point, we also denote
$\|Fo\|=\max\{d(o,fo):f\in F\}$
$\|Fo\|_{\min}=\min\{d(o,fo):f\in F\}$
Let $\alpha:[s,t]\to\mathrm{Y}$ be an arc-length parameterized path from the initial point $\alpha^{-}=\alpha(s)$ to the terminal point $\alpha^{+}=\alpha(t)$. Denote by $[x,y]_{\alpha}$ the parametrized subpath of $\alpha$ going from $x$ to $y$. Precisely, if $x=\alpha(s_{0})$ and $y=\alpha(t_{0})$, $[x,y]_{\alpha}$ denotes the restriction of $\alpha$ on $[s_{0},t_{0}]$. Without the sub-index, $[x,y]$ denotes a choice of a geodesic between $x,y\in\mathrm{Y}$.
Let $\alpha:\mathbb{R}\to\mathrm{Y}$ be a bi-infinite path. The restriction of $\alpha$ to $[a,+\infty)$ for $a\in\mathbb{R}$ is referred to as a positive ray, and its complement a negative ray. By abuse of language, we often denote them by $\alpha^{+}$ and $\alpha^{-}$. In particular, if $\mathrm{Y}$ is compactified with boundary $\partial{\mathrm{Y}}$ and two half ways $\alpha^{\pm}$ converge in the compactification, $\alpha^{+}$ and $\alpha^{-}$ acquire the usual sense as the end points of $\alpha$ in the compact space $\mathrm{Y}\cup\partial{\mathrm{Y}}$.
A path $\alpha$ is called a $c$-quasi-geodesic for $c\geq 1$ if for any subpath $\beta$,
$\ell(\beta)\leq c\cdot d(\beta^{-},\beta^{+})+c$
where $\ell(\beta)$ denotes the length of $\beta$.
Denote by $\alpha\cdot\beta$ (or simply $\alpha\beta$) the concatenation of two paths $\alpha,\beta$ provided that $\alpha^{+}=\beta^{-}$.
Let $f,g$ be real-valued functions. Then $f\prec_{c_{i}}g$ means that there is a constant $C>0$ depending on parameters $c_{i}$ such that $f<Cg$. The symbol $\succ_{c_{i}}$ is defined similarly, and $\asymp_{c_{i}}$ means both $\prec_{c_{i}}$ and $\succ_{c_{i}}$ are true. The constant $c_{i}$ will be omitted if it is a universal constant.
The group action on $\mathrm{Y}$ is always assumed by isometry and the identity in a group $G$ shall be denoted by $1$.
2.2. Contracting subsets
Definition 2.1 (Contracting subset).
For given $C\geq 1$, a subset $U$ in $\mathrm{Y}$ is called $C$-contracting if for any geodesic $\gamma$ with $d(\gamma,U)\geq C$, we have
$\|\pi_{U}(\gamma)\|\leq C.$
A collection of $C$-contracting subsets is referred to as a $C$-contracting system.
Contracting property has several equivalent characterizations. When speaking about $C$-contracting property, the constant $C$ shall be assumed to satisfy the following three statements.
Lemma 2.2.
Let $U$ be a contracting subset. Then there exists $C>0$ such that
  1. (1)
    If $d(\gamma,U)\geq C$ for a geodesic $\gamma$, we have $\|\pi_{U}(\gamma)\|\leq C.$
  2. (2)
    If $\|\pi_{U}(\gamma)\|\geq C$ then $d(\pi_{U}(\gamma^{-}),\gamma)\leq C,\;d(\pi_{U}(\gamma^{+}),\gamma)\leq C$.
  3. (3)
    For a metric ball $B$ disjoint with $U$, we have $\|\pi_{U}(B)\|\leq C$.
$\leq C$$U$$B$$U$$\leq C$$x$$y$$C\geq$$\pi_{U}(x)$$\pi_{U}(y)$
Figure 2. Contracting property in Lemma 2.2
We continue to list a few more consequences used through out this paper.
A subset $U\subseteq\mathrm{Y}$ is called $\sigma$-Morse for a function $\sigma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ if given $c\geq 1$, any $c$-quasi-geodesic with endpoints in $U$ lies in $N_{\sigma(c)}(U)$.
Lemma 2.3.
Let $U\subseteq\mathrm{Y}$ be a $C$-contracting subset for $C>0$.
  1. (1)
    There exists $\sigma=\sigma(C)$ such that $U$ is $\sigma$-Morse.
  2. (2)
    For any $r>0$, there exists $\hat{C}=\hat{C}(C,r)$ such that a subset $V\subseteq\mathrm{Y}$ within $r$-Hausdorff distance to $U$ is $\hat{C}$-contracting.
  3. (3)
    If a geodesic $\gamma$ intersects $N_{r}(U)$ only at the endpoint $\gamma^{-}$ for given $r\geq C$, then
    $\pi_{U}(\gamma^{+})\subseteq B(\gamma^{-},r+C).$
  4. (4)
    For any geodesic $\gamma$, we have
    $\big{|}d_{U}(\gamma^{-},\gamma^{+})-\|\gamma\cap N_{C}(U)\|\big{|}\leq 4C.$
  5. (5)
    There exists $\hat{C}=\hat{C}(C)$ such that $d_{U}(y,z)\leq d(y,z)+\hat{C}$ for any $y,z\in\mathrm{Y}$.
Proof.
All the statements are well-known or straightforward. We leave them as exercises to the interested reader. ∎
In this paper, we are interested in a contracting system $\mathscr{F}$ with $\tau$-bounded intersection property for a function $\tau:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ so that the following holds
$\forall U,V\in\mathscr{F}:\quad\|{N_{r}(U)\cap N_{r}(V)}\|\leq\tau(r)$
for any $r\geq 0$. This is, in fact, equivalent to a bounded projection property of $\mathscr{F}$: there exists a constant $B>0$ such that the following holds
$\|\pi_{U}(V)\|\leq B$
for any $U\neq V\in\mathscr{F}$. See [Yan22, Lemma 2.3] for a proof.
We now state two elementary lemmas used later on, which precises the same idea: the projection of a geodesic to a contracting subset equals the intersection up to a bounded amount. This could be thought of as a general version of the assertion (4) in Lemma 2.3.
Consider a $C$-contracting subset $U\subseteq\mathrm{Y}$ and $r\geq 100C$. Let $\alpha$ be a geodesic so that $\|N_{r}(U)\cap\alpha\|>3r$. Recall that the entry and exit points, $x,y\in\alpha\cap N_{r}(U)$, of $\alpha$ in $N_{r}(U)$ satisfy $d(x,y)=\|N_{r}(U)\cap\alpha\|$.
Lemma 2.4.
The entry and exit points of $\alpha$ in $N_{C}(U)$ are $(3r/2)$-close to the entry and exit points of $\alpha$ in $N_{r}(U)$ respectively. In particular,
(7) $\|N_{C}(U)\cap\alpha\|\geq\|N_{r}(U)\cap\alpha\|-3r$
(8) $|d_{U}(\alpha^{-},\alpha^{+})-\|N_{r}(U)\cap\alpha\||\leq 4r$
Proof.
By hypothesis, we have $d(x,y)>3r$. First notice that $N_{C}(U)\cap\alpha\neq\emptyset$. Otherwise, if $N_{C}(U)\cap\alpha=\emptyset$, the contracting property gives
$d(x,y)\leq d(x,U)+d(y,U)+\|\pi_{U}(\alpha)\|\leq 2r+C<3r.$
This is a contradiction, confirming the above observation.
Let us thus choose $u,v\in[x,y]\cap N_{C}(U)$ such that $d(u,v)=\|N_{C}(U)\cap\alpha\|$. Using again contracting property of $U$, we obtain
$d(x,u)\leq d(x,U)+d(u,U)+\|\pi_{U}([x,u])\|\leq r+2C\leq 3r/2.$
Similarly, $d(v,y)\leq r+2C\leq 3r/2$. Thus, $d(u,v)\geq d(x,y)-3r$, so the (7) follows.
To show the inequality (8), note that $\pi_{U}(\alpha^{-})$ (resp. $\pi_{U}(\alpha^{+})$) is $2C$-close to $u$ (resp. $v$). Hence, (8) follows from (7). The proof is complete. ∎
Lemma 2.5.
Let $\beta$ be a geodesic such that $d_{U}(\alpha^{-},\beta^{-})\leq 10C$ and $d_{U}(\alpha^{+},\beta^{+})\leq 10C$. Then the entry and exit points of $\beta$ in $N_{r}(U)$ are $(4r)$-close to $x$ and $y$ respectively.
Proof.
We first prove that $\beta\cap N_{C}(U)\neq\emptyset$. Indeed, assume to the contrary that $\beta\cap N_{C}(U)=\emptyset$, so $\|\pi_{U}(\beta)\|\leq C$ by the contracting property.
Since $[x,\alpha^{-}]\cap N_{r}(U)=\emptyset$ and $[\alpha^{-},y]\cap N_{r}(U)=\emptyset$, the contracting property again shows $d_{U}(x,\alpha^{-})\leq C$ and $d_{U}(\alpha^{-},y)\leq C$. We compute
$\begin{array}[]{rl}d(x,y)\leq&d(x,U)+d_{U}(x,\alpha^{-})+d_{U}(\alpha^{-},\beta^{-})\\ &+\|\pi_{U}(\beta)\|+d_{U}(\alpha^{+},\beta^{+})+d_{U}(\alpha^{-},y)+d(y,U)\\ \leq&14C+2r<3r.\end{array}$
This contradicts the assumption that $d(x,y)>3r$. Thus, $\beta\cap N_{C}(U)\neq\emptyset$ is proved, so we can choose $u,v\in\beta\cap N_{C}(U)$ such that $d(u,v)=\|N_{C}(U)\cap\beta\|$. Using the contracting property of $U$, we obtain
$\begin{array}[]{rl}d(x,u)\leq&d(x,U)+d_{U}(x,\alpha^{-})+d_{U}(\alpha^{-},\beta^{-})+d_{U}(\beta^{-},u)+d(u,U)\\ \leq&2r+12C\leq(5/2)r.\end{array}$
Similarly, we have $d(v,y)\leq 2r+12C\leq(5/2)r$.
By Lemma 2.4, the enter and exit points of $\beta$ in $N_{r}(U)$ are $(3r/2)$-close to $u$ and $v$ respectively, so are $(4r)$-close to the enter and exit points $x,y$ of $\alpha$ respectively. The proof of the lemma is proved. ∎
2.3. Contracting elements
A group $H$ is called contracting if for some (hence any) $o\in\mathrm{Y}$, the subset $Ho$ is contracting in $\mathrm{Y}$.
An isometry $h\in\textrm{Isom}(\mathrm{Y})$ is called contracting if it is of infinite order and the subgroup $\langle h\rangle$ is contracting. The set of contracting isometries is preserved under conjugacy. We say that $h$ has the quasi-isometric image property if the orbital map
(9) $n\in\mathbb{Z}\to h^{n}o\in\mathrm{Y}$
is a quasi-isometric embedding. We then define the axis of $h$ as the following concatenating path:
(10) $\textrm{Ax}(h)=\cup_{n\in\mathbb{Z}}h^{n}[o,ho].$
Let $G<\textrm{Isom}(\mathrm{Y})$ be a discrete group, that is, which acts properly on $\mathrm{Y}$.
Lemma 2.6.
[Yan19, Lemma 2.11] Let $h\in G$ be a contracting element with quasi-isometric image. Then $h$ is contained in a maximal elementary subgroup defined as follows
$E(h)=\{g\in G:\exists n\in\mathbb{N}_{>0},(\;gh^{n}g^{-1}=h^{n})\;\lor\;(gh^{n}g^{-1}=h^{-n})\}.$
Keeping in mind the basepoint $o\in\mathrm{Y}$, the axis of $h$ is defined as the following quasi-geodesic
(11) $\textrm{Ax}(h)=\{fo:f\in E(h)\}.$
Notice that $\textrm{Ax}(h)=\textrm{Ax}(k)$ and $E(h)=E(k)$ for any contracting element $k\in E(h)$.
Remark 2.7 (Definition of axis).
For most results in the paper, contracting elements are contained in a discrete group, for which we make use of the axis in (11). As $\langle h\rangle$ is of finite index, $E(h)\cdot o$ has a finite Hausdorff distance to $\langle h\rangle\cdot o$. Hence, there exists no essential difference in these two definitions.
An element $g\in G$ preserves the orientation of a bi-infinite quasi-geodesic $\gamma$ if $\alpha$ and $g\alpha$ has finite Hausdorff distance for any half ray $\alpha$ of $\gamma$. Let $E^{+}(h)$ be the subgroup of $E(h)$ with possibly index 2 whose elements preserve the orientation of their axis. Then we have
$E^{+}(h)=\{g\in G:\exists n\in\mathbb{N}_{>0},\;gh^{n}g^{-1}=h^{n}\}.$
and $E^{+}(h)$ contains all contracting elements in $E(h)$. In the sequel, unless explicitly mentioned, we always assume contracting elements to satisfy quasi-isometrically embedded image property (9).
Two contracting elements $h_{1},h_{2}\in G$ are called independent if the collection $\{g\textrm{Ax}(h_{i}):g\in G;\ i=1,2\}$ is a contracting system with bounded intersection. Note that two conjugate contracting elements with disjoint fixed points are not independent in our sense. This is slightly different from the use of independence by other researchers (cf. [MT18]).
Definition 2.8.
Fix $r>0$ and a set $F$ in $G$. A geodesic $\gamma$ contains an $(r,f)$-barrier for $f\in F$ if there exists an element $g\in G$ so that
(12) $\max\{d(g\cdot o,\gamma),\;d(g\cdot fo,\gamma)\}\leq r.$
By abuse of language, the point $go$ or the axis $g\textrm{Ax}(f)$ is called $(r,F)$-barrier of type $f$ on $\gamma$.
Recall that $\|Fo\|$ denotes the diameter of $Fo$, and $\|Fo\|_{\min}:=\min\{d(o,fo):f\in F\}$ the minimal length of elements in $F$. As $F$ is usually chosen as a finite set (of three elements), their distinction does not really matter and could be ignored on a first reading.
Lemma 2.9.
For any $r>0$ there exists $\hat{r}=\hat{r}(r)>0$ with the following property.
Let $f$ be a contracting element with $d(o,fo)>3r$. Suppose that a geodesic $\alpha$ contains an $(r,f)$-barrier $go$. Set $X=g\textrm{Ax}(f)$. Let $\beta$ be a geodesic such that $d_{X}(\alpha^{-},\beta^{-})\leq 10C$ and $d_{X}(\alpha^{+},\beta^{+})\leq 10C$. Then $go$ is an $(\hat{r},f)$-barrier for $\beta$.
Proof.
Let $x,y\in\alpha\cap N_{r}(X)$ be the corresponding entry and exit points. The $(r,f)$-barrier $go$ for $\alpha$ implies that $d(go,[x,y]_{\alpha})\leq r$ and $d(gfo,[x,y]_{\alpha})\leq r$. By Lemma 2.5, the entry and exit points $u,v$ of $\beta$ in $N_{r}(X)$ are $(4r)$-close to those $x,y$ respectively.
As $X$ is a contracting quasi-geodesic, any geodesic subsegment (say $[u,v]$) in its $r$-neighborhood is also uniformly contracting by [Yan20, Prop. 2.2]. By Morse property, there exists $\tau=\tau(r)>0$ such that $[x,y]$ is contained in a $\tau$-neighborhood of $[u,v]$. Setting $\hat{r}=4r+\tau$ then completes the proof. ∎
Let $h$ be a contracting element with a $C$-contracting axis $\gamma:=\textrm{Ax}(h)$ in (11). Here we understand $\gamma$ as a quasi-geodesic path for simplicity. Let $f\in E^{+}(h)$ be an orientation-preserving element which must be contracting. By the Morse property, there exist constants $r,L$ depending only on $C$ such that any geodesic segment in $N_{r}(\gamma)$ of length $L+d(o,fo)$ contains $[o,fo]$. We record this fact as the following.
Lemma 2.10.
There exist $r=r(C),L=L(C)>0$ with the following property.
Let $x,y\in\gamma\cap N_{r}(\alpha)$ be the entry and exit points of $\gamma$ in a geodesic segment $\alpha$. If $z\in[x,y]_{\gamma}$ is a point such that $\min\{d(z,x),d(z,y)\}\geq L+d(o,fo)$, then $z$ is an $(r,F)$-barrier for $\alpha$, where $F:=\{f,f^{-1}\}$
Lastly, the following result will be used later on.
Lemma 2.11.
[Yan14a, Lemma 4.6] Let $G\curvearrowright\mathrm{Y}$ be as in (1) and $\Gamma<G$ be an infinite normal subgroup. Then $\Gamma$ must be non-elementary and contains infinitely many independent contracting elements.
2.4. Admissible paths
Let $\mathscr{F}$ be a $C$-contracting system in the space $\mathrm{Y}$; that is, a collection of $C$-contracting subsets for a common $C>0$. A class of admissible paths introduced in [Yan14a] is of great importance in our study, which are quasi-geodesics with fellow travel property relative to such a system $\mathscr{F}$. In most cases, this paper considers the $C$-contracting system $\mathscr{F}$ as in (3).
[…]
Figure 3. Admissible path
Definition 2.12 (Admissible Path).
Given $L,B\geq 0$, a path $\gamma$ is called $(L,B)$-admissible relative to $\mathscr{F}$, if the path $\gamma=p_{0}q_{1}p_{1}\ldots q_{n}p_{n}$ is a piece-wise geodesic path satisfying the Long Local and Bounded Projection properties:
  1. (LL)
    Each $p_{i}$ for $0\leq i\leq n$ has the two endpoints in $X_{i}\in\mathscr{F}$ and length bigger than $L$ unless $i=0$ or $i=n$,
  2. (BP)
    For each $X_{i}$, we have $X_{i}\neq X_{i+1}$ and
    $\|\pi_{X_{i}}(q_{i+1})\|\leq B,\quad\|\pi_{X_{i}}(q_{i})\|\leq B$
    whenever the previous and next paths $q_{i},q_{i+1}$ are defined. See Fig. 3 for an illustration.
The indexed collection of $X_{i}\in\mathscr{F}$ as above, denoted by $\mathscr{F}(\gamma)$, will be referred to as the saturation of $\gamma$.
In the case that $p_{n}$ (or $q_{n}$ if $p_{n}$ is trivial) is a geodesic ray, we say that $\gamma$ is an admissible ray.
As the properties (LL) and (BP) are local conditions, we can produce admissible paths via the following operations:
  • Subpath A subpath of an $(L,B)$-admissible path is $(L,B)$-admissible.
  • Concatenation Let $\alpha,\beta$ be two $(L,B)$-admissible paths so that the last contracting subset associated to the geodesic $p$ of $\alpha$ is the same as the first contracting subset to the geodesic $q$ of $\beta$. If $d(p^{-},q^{+})>L$, the path $\gamma$ by concatenating paths
    $\gamma:=[\alpha^{-},p^{-}]_{\alpha}\cdot[p^{-},q^{+}]\cdot[q^{+},\beta^{+}]_{\beta}$
    is $(L,B)$-admissible.
A sequence of points $x_{i}$ in a path $p$ is called linearly ordered if $x_{i+1}\in[x_{i},p^{+}]_{p}$ for each $i$.
Definition 2.13 (Fellow travel).
Let $\gamma=p_{0}q_{1}p_{1}\cdots q_{n}p_{n}$ be an $(L,B)-$admissible path. We say $\gamma$ has $r$-fellow travel property for some $r>0$ if for any geodesic $\alpha$ with the same endpoints as $\gamma$, there exists a sequence of linearly ordered points $z_{i},w_{i}$ ($0\leq i\leq n$) on $\alpha$ such that
$d(z_{i},p_{i}^{-})\leq r,\quad d(w_{i},p_{i}^{+})\leq r.$
In particular, $\|N_{r}(X_{i})\cap\alpha\|\geq L$ for each $X_{i}\in\mathscr{F}(\gamma)$.
Remark 2.14.
Note that $\|N_{r}(X_{i})\cap\alpha\|\geq L$ is basically amount to saying that $X_{i}$ is an $(r,F)$-barrier for $\alpha$. Thus, the $R$-fellow travel property is closely related to the existence of $(r,F)$-barriers. These cumbersome terminologies may have their own convenience in practice.
The following result says that a local long admissible path enjoys the fellow travel property.
Proposition 2.15.
[Yan14a] For any $B>0$, there exist $L,r>0$ depending only on $B,C$ such that any $(L,B)$-admissible path has $r$-fellow travel property.
The constant $B$ in an $(L,B)$-admissible path can be made uniform by the following operation.
  • Truncation By the contracting property, a geodesic issuing from a $C$-contracting subset diverges in an orthogonal way from its $C$-neighborhood. At the expense of decreasing $L$, we can thus truncate these diverging parts to obtain the constant $B$ depending only on $\mathscr{F}$.
We explain this procedure in a more general form.
[…]
Figure 4. Truncating a geodesic around $N_{C}(X_{i})$.
Lemma 2.16.
Assume that $\|Fo\|_{\min}>3r$ for some $r>0$. There exist $L=L(\|Fo\|_{\min},r),B=B(\mathscr{F})>0$ with the following property.
Let $\alpha$ be a geodesic with a set of distinct $(r,F)$-barriers $X_{i}\in\mathscr{F}$. Then there exists an $(L,B)$-admissible path $\hat{\alpha}$ with the same endpoints as $\alpha$ with saturation $\{X_{i}\}$.
The path $\hat{\alpha}$ is called truncation of $\alpha$. If $\alpha$ is a fellow travel geodesic for the admissible path $\gamma$ in Proposition 2.15, then $\alpha$ has $(r,F)$-barriers $X\in\mathscr{F}(\gamma)$. The truncation $\hat{\alpha}$ will also be referred as truncation of $\gamma$ and has a uniform constant $B$.
Proof.
By definition, $\|\alpha\cap N_{r}(X_{i})\|\geq\|Fo\|_{\min}$. Consider the entry and exit points $z_{i}^{\prime},w_{i}^{\prime}\in\alpha$ in $N_{C}(X_{i})$ by Lemma 2.4 such that
$d(z_{i}^{\prime},w_{i}^{\prime})=\|\alpha\cap N_{C}(X_{i})\|\geq\|Fo\|_{\min}-3r.$
Let $z_{i},w_{i}\in X_{i}$ such that $d(z_{i},z_{i}^{\prime})=d(w_{i},w_{i}^{\prime})=C$, where $z_{0}:=\alpha^{-}$ and $w_{n}:=\alpha^{+}$. See Fig. 4. Since $\mathscr{F}$ has $\tau$-bounded intersection and then bounded projection, one deduces
$\|\pi_{X_{i}}([w_{i},z_{i+1}])\|\leq B$
for some $B$ depending only on $\tau$ and $C$. Note that $d(z_{i},w_{i})\geq L:=\|Fo\|_{\min}-3r-2C$. Denoting $\hat{p}_{i}=[z_{i},w_{i}]$ and $\hat{q}_{i+1}=[w_{i},z_{i+1}]$, the path
$\hat{\alpha}:=\hat{p}_{0}\hat{q}_{1}\hat{p}_{1}\ldots\hat{q}_{n}\hat{p}_{n}$
is an $(L,B)$-admissible path with the same endpoints as $\alpha$. ∎
2.5. Extension lemma
Admissible paths have demonstrated strong fellow travel properties. We now introduce a basic tool to build admissible paths.
Lemma 2.17 (Extension Lemma).
There exist $L,r,B>0$ depending only on $C$ with the following property.
Let $h_{1},h_{2},h_{3}\in G$ be three independent contracting elements. Choose any element $f_{i}\in\langle h_{i}\rangle$ for each $1\leq i\leq 3$ to form the set $F$ satisfying $\|Fo\|_{\min}\geq L$. Let $g\in G$ and $\alpha$ be a geodesic starting at $\alpha^{-}=o$.
  1. (1)
    There exists an element $f\in F$ such that the path
    $\gamma:=[o,go]\cdot(g[o,fo])\cdot(gf\alpha)$
    is an $(L,B)$-admissible path relative to $\mathscr{F}$ (defined in (3)).
  2. (2)
    The point $go$ is an $(r,f)$-barrier for any geodesic $[\gamma^{-},\gamma^{+}]$.
Remark 2.18.
  1. (1)
    In [Yan19], the geodesic $\alpha=[o,ho]$ is chosen to be ending at $Go$ for some $h\in G$. However, the proof given there works for a geodesic starting from $o$ possibly ending at any point in $\mathrm{Y}$.
  2. (2)
    Since admissible paths are local conditions, we can connect via $F$ any number of elements $g\in G$ to satisfy (1) and (2). We refer the reader to [Yan19] for precise formulation.
2.6. Projection complex
In this subsection, we briefly recall the work of Bestvina-Bromberg-Fujiwara [BBF15] on constructing a quasi-tree of spaces.
Definition 2.19 (Projection axioms).
Let $\mathscr{F}$ be a collection of metric spaces equipped with (set-valued) projection maps
$\{\pi_{U}:\mathscr{F}\setminus\{U\}\to U\}_{U\in\mathscr{F}}.$
Denote $d_{U}(V,W):=\|\pi_{U}(V)\cup\pi_{U}(W)\|$ for $V\neq U\neq W\in\mathscr{F}$. The pair (,$(\mathscr{F},\{\pi_{U}\}_{U\in\mathscr{F}})$ satisfies projection axioms for a constant $\kappa\geq 0$ if
  1. (1)
    $\|\pi_{U}(V)\|\leq\kappa$ when $U\neq V$.
  2. (2)
    if $U,V,W$ are distinct and $d_{V}(U,W)>\kappa$ then $d_{U}(V,W)\leq\kappa$.
  3. (3)
    the set $\{U\in\mathscr{F}\,:\,d_{U}(V,W)>\kappa\}$ is finite for $V\neq W$.
By definition, the following triangle inequality holds
(13) $d_{Y}(V,W)\leq d_{Y}(V,U)+d_{Y}(U,W).$
It is well-known that the projection axioms hold for a contracting system with bounded intersection (cf. [Yan14a, Appendix]).
Lemma 2.20.
Let $h$ be any contracting element in a proper action of $G\curvearrowright\mathrm{Y}$. Then the collection =$\mathscr{F}=\{g\textrm{Ax}(h):g\in G\}$ with shortest projection maps $\pi_{U}(V)$ satisfies the projection axioms with constant $\kappa=\kappa(\mathscr{F})>0$.
In [BBF15, Def. 3.1], a modified version of $d_{U}$ is introduced so that it is symmetric and agrees with the original $d_{U}$ up to an additive amount $2\kappa$. So, the axioms (1)-(3) are still true for $3\kappa$, and the triangle inequality in (13) holds up to a uniform error. In what follows, we actually need to work with this modified $d_{U}$ to define the complex projection, but for sake of simplicity, we stick on the above definition of $d_{U}$.
We consider the interval-like set for $K>0$ and $V,W\in\mathscr{F}$ as follows
$\mathscr{F}_{K}(V,W):=\{U\in\mathscr{F}:d_{U}(V,W)>\kappa\}.$
Denote $\mathscr{F}_{K}[V,W]:=\mathscr{F}_{K}(V,W)\bigcup\{V,W\}$. It possesses a total order described in the next lemma.
Lemma 2.21.
[BBF15, Theorem 3.3.G] There exist constants $D=D(\kappa),K=K(\kappa)>0$ for the above $\mathscr{F}$ such that the set $\mathscr{F}_{K}[V,W]$ with order “$<$” is totally ordered with least element $V$ and great element $W$, such that given $U_{0}$, $U_{1}$, $U_{2}\in\mathscr{F}_{K}[V,W]$, if $U_{0}<U_{1}<U_{2}$, then
$d_{U_{1}}(V,W)-D\leq d_{U_{1}}(U_{0},U_{2})\leq d_{U_{1}}(V,W),$
and
$d_{U_{0}}(U_{1},U_{2})\leq D\quad and\quad d_{U_{2}}(U_{0},U_{1})\leq D.$
We now give the definition of a projection complex.
Definition 2.22.
The projection complex $\mathcal{P}_{K}(\mathscr{F})$ for $K$ satisfying Lemma 2.21 is a graph with the vertex set consisting of the elements in $\mathscr{F}$. Two vertices $u$ and $v$ are connected if $\mathscr{F}_{K}(U,V)=\emptyset$. We equip $\mathcal{P}_{K}(\mathscr{F})$ with a length metric $d_{\mathcal{P}}$ induced by assigning unit length to each edge.
The projection complex $\mathcal{P}_{K}(\mathscr{F})$ is connected since by [BBF15, Proposition 3.7] the interval set $\mathscr{F}_{K}[U,V]$ gives a connected path between $U$ and $V$ in $\mathcal{P}_{K}(\mathscr{F})$ called standard path: the consecutive elements directed by the total order are adjacent in $\mathcal{P}_{K}(\mathscr{F})$. By [BBFS19, Corollary 3.7], standard paths are $(2,1)$-quasi-geodesics. The structural result about the projection complex is the following.
Theorem 2.23.
[BBF15] For $K\gg 0$ as in Lemma 2.21, the projection complex $\mathcal{P}_{K}(\mathscr{F})$ is a quasi-tree, on which $G$ acts co-boundedly.
Proof.
It is proved there that the standard path from $U$ to $V$ lies in the $2$-neighborhood of any path with the same endpoints. This implies the conclusion via (a variant of) Manning’s bottleneck criterion of quasi-trees in [Man05]. ∎
A triangle formed by standard paths in the projection complex looks almost like a tripod.
Lemma 2.24.
[BBFS19, Lemma 3.6] For every $U,V,W\in\mathscr{F}$, the path $\mathscr{F}_{K}[U,V]$ is contained in $\mathscr{F}_{K}[U,W]\cup\mathscr{F}_{K}[W,V]$ except for at most two consecutive vertices. Precisely, the (possibly empty) initial and terminal subpaths of $\mathscr{F}_{K}[U,V]$ are contained in $\mathscr{F}_{K}[U,W]\cup\mathscr{F}_{K}[W,V]$, with an exception of at most two vertices in the middle.
For any two points $x\in X$ and $z\in Z$, we often need to lift a standard path to $\mathrm{Y}$. It is a piece-wise geodesic path (admissible path) as concatenation of the normal paths between two consecutive vertices and geodesics contained in vertices. This is explained by the following lemma proved in [HLY20, Lemma 4.5]; we include the proof for completeness.
Lemma 2.25.
For any $K>0$, there exist a constant $L=L(K,\kappa)\geq 0$ with $L\to\infty$ as $K\to\infty$, and a constant $B=B(\kappa)>0$ with the following property.
For any two points $u\in U,v\in V$ there exists an $(L,B)$-admissible path $\gamma$ in $\mathrm{Y}$ from $u$ to $v$ with saturation $\mathscr{F}_{K}[U,V]$.
Proof.
List $\mathscr{F}_{K}[U,V]=\{S_{0}=U,S_{1},S_{2},\cdots,S_{k},S_{k+1}=V\}$ by the total order in Lemma 2.21. The admissible path is constructed by connecting projections between $S_{i}$ and $S_{i+1}$. Namely, choose a sequence of points $x_{i}\in\pi_{S_{i}}(S_{i+1}),y_{i}\in\pi_{S_{i+1}}(S_{i})$ for $0\leq i\leq k$. We connect consecutively the points in $\{u,x_{0},y_{0},\cdots,x_{k},y_{k},v\}$ to get a piecewise geodesic path $\gamma$. See Fig. 5. By bounded projection $\|\pi_{S_{i}}(S_{i+1})\|\leq B$, we have
$d_{S_{i}}(x_{i-1},y_{i-1}),\;d_{S_{i}}(x_{i},y_{i})\leq B.$
By Lemma 2.21, we have
$d_{S_{i}}(S_{i-1},S_{i+1})\geq d_{S_{i}}(U,V)-D\geq K-D,$
for a constant $D=D(\kappa)$. Thus, $d(y_{i-1},x_{i})\geq d_{S_{i}}(S_{i-1},S_{i+1})-2B\geq K-D-2B:=L(K,\kappa)$. So, $\gamma$ is an $(L,B)$-admissible path relative to $\mathscr{F}_{K}[U,V]$. The lemma is then proved. ∎
[…]
Figure 5. Schematic illustration of lifting standard paths
In practice, we always assume that $K$ is large enough so that $L$ satisfies Proposition 2.15 and the path $\gamma$ is a $c$-quasi-geodesic for some $c>1$.
Before moving on, let us put into perspective what we have proved so far about projection complex.
Consider the union $\widetilde{X}:=\bigcup\{X\in\mathscr{F}\}$ of $\mathscr{F}$ with the restricted metric $d$ from $\mathrm{Y}$. Then the natural projection map collapsing $X\in\mathscr{F}$ to the corresponding vertex $\bar{x}\in\mathcal{P}_{K}(\mathscr{F})$ defined as follows
(14) $\Phi_{1}:(\widetilde{X},d)$ $\longrightarrow(\mathcal{P}_{K}(\mathscr{F}),d_{\mathcal{P}})$
$X$ $\longmapsto\bar{x}$
is a coarsely Lipschitz map: there exists a constant $c>1$ so that for any $x,y\in\widetilde{X}$,
$d_{\mathcal{P}}(\bar{x},\bar{y})\leq cd(x,y)+c$
Indeed, this follows from the fact that the lifted path $\gamma$ in Lemma 2.25 is a uniform quasi-geodesic. Moreover, by Proposition 2.15, as $\gamma$ fellow travels any geodesic $\alpha$ with the same endpoints as $\gamma$, we could deduce that the map $\Phi$ sends the geodesic $\alpha$ to a unparametrized quasi-geodesic (standard path) in $\mathcal{P}_{K}(\mathscr{F})$: the gaps between two consecutive axis in the saturation $\mathscr{F}(\gamma)$ are sent to the corresponding edges. A map between two metric spaces with the property is usually called shadowing map in literature.
We now recall the following result to be used later on.
Lemma 2.26.
[BBF15, Lemma 3.18] For any $K>0$ there exists $\hat{K}>0$ such that $\mathscr{F}_{\hat{K}}(X,Z)$ is contained in the geodesic from $X$ to $Z$ in $\mathcal{P}_{K}(\mathscr{F})$.
Quasi-tree of spaces.
Fix a positive number $L$ so that $1/2K\leq L\leq 2K$. We now define a blowup version, $\mathcal{C}(\mathscr{F})$, of the projection complex $\mathcal{P}_{K}(\mathscr{F})$ by remembering the geometry of each $U\in\mathscr{F}$. Precisely,
We replace each $U\in\mathscr{F}$, a vertex in $\mathcal{P}_{K}(\mathscr{F})$, by the corresponding subspace $U\subset\mathrm{Y}$, and keep the adjacency relation in $\mathcal{P}_{K}(\mathscr{F})$: if $U$ and $V$ are adjacent in $\mathcal{P}_{K}(\mathscr{F})$ (i.e. $d_{\mathcal{P}}(U,V)=1$), then we attach an edge of length $L$ from every point $u\in\pi_{U}V$ to $v\in\pi_{V}U$. This choice of $L$ by [BBF15, Lemma 4.2] ensures that $U\subseteq\mathrm{Y}$ are geodesically embedded in $\mathcal{C}(\mathscr{F})$ (for simplicity, we mention $K$ rather than $L\in[K/2,2K]$).
Since $h$ is contracting, the infinite cyclic subgroup $\langle h\rangle$ is of finite index in $E(h)$ by Lemma 2.6, so $\textrm{Ax}(h)=E(h)o$ is quasi-isometric to a line $\mathbb{R}$. Thus, $\mathscr{F}$ consists of uniform quasi-lines. By [BBF15, Theorem B], we have the following.
Theorem 2.27.
[BBF15] The quasi-tree of spaces $\mathcal{C}(\mathscr{F})$ is a quasi-tree of infinite diameter, with each $U\in\mathscr{F}$ totally geodesically embedded into $\mathcal{C}(\mathscr{F})$. Moreover, the shortest projection from $U$ to $V$ in $\mathcal{C}(\mathscr{F})$ agrees with the projection $\pi_{U}V$ up to a uniform finite Hausdorff distance.
From the construction, there exists a natural map which collapses each totally geodesically embedded subspace $U\subseteq\mathcal{C}(\mathscr{F})$ as a point and identifies edges of length $L$ to edges of unit length in $\mathcal{P}_{K}(\mathscr{F})$. In a reverse direction, the connected path in $\mathcal{P}_{K}(\mathscr{F})$ given by $\mathscr{F}_{K}[X,Y]$ lifts to a standard path in $\mathcal{C}(\mathscr{F})$ described as follows. The notion of standard paths plays a key role to establish Theorem 2.27.
Definition 2.28.
A path $\gamma$ from $u\in U$ to $v\in V$ in the space $\mathcal{C}(\mathscr{F})$ is called a $K$-standard path if it passes through the set of vertices in $\mathscr{F}_{K}[U,V]$ in a natural order given in Lemma 2.21 and, within each vertex, the path is a geodesic.
Hence, standard paths in $\mathcal{C}(\mathscr{F})$ are also uniform quasi-geodesics.
Summarizing the above discussion, we also have the coarsely Lipschitz map:
(15) $\Phi_{2}:(\mathcal{C}(\mathscr{F}),d)$ $\longrightarrow(\mathcal{P}_{K}(\mathscr{F}),d)$
$X$ $\longmapsto\bar{x}$
which sends a standard path in $\mathcal{C}(\mathscr{F})$ to the corresponding standard path in $\mathcal{P}_{K}(\mathscr{F})$.
As each $U\in\mathscr{F}$ is isometrically embedded into $\mathcal{C}(\mathscr{F})$, the identification
(16) $\Phi:(\widetilde{X},d)$ $\longrightarrow(\mathcal{C}(\mathscr{F}),d)$
$x$ $\longmapsto x$
is a coarsely Lipschitz map, where we recall $\widetilde{X}=\bigcup\{X\in\mathscr{F}\}$.
We shall continue the investigation on the Gromov boundary of projection complex in §4.4.
2.7. Horofunction boundary
We recall the definition of horofunction boundary and setup the notations for further reference.
Fix a basepoint $o\in\mathrm{Y}$. For each $y\in\mathrm{Y}$, we define a Lipschitz map $b_{y}^{o}:\mathrm{Y}\to\mathbb{R}$ by
$\forall x\in\mathrm{Y}:\quad b^{o}_{y}(x)=d(x,y)-d(o,y).$
This family of $1$-Lipschitz functions sits in the set of continuous functions on $\mathrm{Y}$ vanishing at $o$. Endowed with the compact-open topology, the Arzela-Ascoli Lemma implies that the closure of $\{b^{o}_{y}:y\in\mathrm{Y}\}$ gives a compactification of $\mathrm{Y}$. The complement, denoted by $\partial_{\mathrm{H}}{\mathrm{Y}}$, of $\mathrm{Y}$ is called the horofunction boundary.
A Buseman cocycle $B_{\xi}:\mathrm{Y}\times\mathrm{Y}\to\mathbb{R}$ at $\xi\in\partial_{\mathrm{H}}{\mathrm{Y}}$ (independent of $o$) is given by
$\forall x_{1},x_{2}\in\mathrm{Y}:\quad B_{\xi}(x_{1},x_{2})=b_{\xi}^{o}(x_{1})-b_{\xi}^{o}(x_{2}).$
The topological type of horofunction boundary is independent of the choice of basepoints, since if $d(x,y_{n})-d(o,y_{n})$ converges as $n\to\infty$, then so does $d(x,y_{n})-d(o^{\prime},y_{n})$ for any $o^{\prime}\in\mathrm{Y}$. Moreover, the corresponding horofunctions differ by an additive amount:
$b^{o}_{\xi}(\cdot)-b_{\xi}^{o^{\prime}}(\cdot)=b_{\xi}^{o}(o^{\prime}),$
so we will omit the upper index $o$. Every isometry $\phi$ of $\mathrm{Y}$ induces a homeomorphism on $\partial_{\mathrm{H}}{\mathrm{Y}}$:
$\forall y\in\mathrm{Y}:\quad\phi(\xi)(y):=b_{\xi}(\phi^{-1}(y))-b_{\xi}(\phi^{-1}(o)).$
According to the context, both $\xi$ and $b_{\xi}$ are used to denote the boundary points.
Finite difference relation
Two horofunctions $b_{\xi},b_{\eta}$ have $K$-finite difference for $K\geq 0$ if the $L^{\infty}$-norm of their difference is $K$-bounded:
$\|b_{\xi}-b_{\eta}\|_{\infty}\leq K.$
The locus of $b_{\xi}$ consists of horofunctions $b_{\eta}$ so that $b_{\xi},b_{\eta}$ have $K$-finite difference for some $K$ depending on $\eta$. The loci $[b_{\xi}]$ of horofunctions $b_{\xi}$ form a finite difference equivalence relation []delimited-[][$\cdot$][ ⋅ ] on $\partial_{\mathrm{H}}{\mathrm{Y}}$. The locus $[\Lambda]$ of a subset $\Lambda\subseteq\partial_{\mathrm{H}}{\mathrm{Y}}$ is the union of loci of all points in $\Lambda$. We say that $\Lambda$ is saturated if $[\Lambda]=\Lambda$.
The following fact follows directly by definition.
Lemma 2.29.
Let $x_{n}\in\mathrm{Y}\to\xi\in\partial_{\mathrm{H}}{\mathrm{Y}}$ and $y_{n}\in\mathrm{Y}\to\eta\in\partial_{\mathrm{H}}{\mathrm{Y}}$ as $n\to\infty$. If $\sup_{n\geq 1}d(x_{n},y_{n})<\infty$, then $[\xi]=[\eta]$.
We say that the partition []delimited-[][$\cdot$][ ⋅ ] restricted on a saturated subset $\Lambda$ is called $K$-finite for $K\geq 0$ if any []delimited-[][$\cdot$][ ⋅ ]-class in $\Lambda$ consists of horofunctions with pairwise $K$-finite difference.
At last, we mention the following topological criterion to obtain second countable, Hausdorff quotient spaces.
Lemma 2.30.
Let $\Lambda\subseteq\partial_{\mathrm{H}}{\mathrm{Y}}$ be a saturated subset so that the restricted relation []delimited-[][$\cdot$][ ⋅ ] on $\Lambda$ is $K$-finite for some $K\geq 0$. Then the quotient map
$[\cdot]:\quad\xi\in\Lambda\longmapsto[\xi]\in[\Lambda]$
is a closed map from $\Lambda$ to $[\Lambda]$ with compact fibers. Moreover, $[\Lambda]$ has a metrizable and second countable quotient topology.
Proof.
A closed continuous surjective map $f:M\to N$ between two topological spaces with compact fibers is called perfect in [Mun00, Ex §31.7]. A prefect map preserves the Hausdorff, regular, and second countable properties of $\Lambda$, so by Urysohn’s metrization theorem, $[\Lambda]$ is metrizable. Hence, we only need to prove that the quotient map []delimited-[][$\cdot$][ ⋅ ] is a closed map.
Note first that the restricted $K$-finite difference relation []delimited-[][$\cdot$][ ⋅ ] on $\Lambda$ is closed. It suffices to recall that $b_{\xi_{n}}\to b_{\xi}$ is amount to the uniform convergence of $b_{\xi_{n}}(\cdot)\to b_{\xi}(\cdot)$ on any compact subset of $\mathrm{Y}$. Thus, $|b_{\xi_{n}}-b_{\eta_{n}}|_{\infty}\leq K$ implies $|b_{\xi}-b_{\eta}|_{\infty}\leq K$. Moreover, the []delimited-[][$\cdot$][ ⋅ ]-class of a point in $\Lambda$ is a closed subset, so is compact in $\partial_{\mathrm{H}}{\mathrm{Y}}$. Thus, the map []delimited-[][$\cdot$][ ⋅ ] on $\Lambda$ has compact fibers.
We now show the quotient map []delimited-[][$\cdot$][ ⋅ ] is closed: the locus $[Z]$ (the preimage of the map) of a closed subset $Z\subseteq\Lambda$ is closed in $\Lambda$. Indeed, let $w_{n}\in[Z]$ tend to a limit point $\xi\in\Lambda$. Take a sequence $z_{n}\in Z$ with $w_{n}\in[z_{n}]$, and for $\partial{\mathrm{Y}}$ is compact, let $\eta\in\partial{\mathrm{Y}}$ be any accumulation point of $z_{n}$. As the relation []delimited-[][$\cdot$][ ⋅ ] on $\Lambda$ is closed, we have $[\xi]=[\eta]$ and then $\eta\in\Lambda$ for $\Lambda$ is saturated. As $Z$ is a closed subset, we have $\eta\in Z$, so $\xi\in[Z]$. This implies that $[Z]$ is a closed subset in $\Lambda$. Hence, the quotient map is closed, and the lemma is proved. ∎
3. Convergence compactification
We develop a general theory of a convergence compactification. As in (1), let $(\mathrm{Y},d)$ be a proper metric space admitting an isometric action of a non-elementary countable group $G$ with a contracting element. Consider a metrizable compactification $\overline{\mathrm{Y}}:=\partial{\mathrm{Y}}\cup\mathrm{Y}$, so $\mathrm{Y}$ is open and dense in $\overline{\mathrm{Y}}$. We also assume that the action of $\textrm{Isom}(\mathrm{Y})$ extends by homeomorphism to $\partial{\mathrm{Y}}$.
3.1. Partition on a boundary
We equip $\partial{\mathrm{Y}}$ with an $\textrm{Isom}(\mathrm{Y})$-invariant partition []delimited-[][$\cdot$][ ⋅ ]: $[\xi]=[\eta]$ implies $[g\xi]=[g\eta]$ for any $g\in\textrm{Isom}(\mathrm{Y})$. The locus $[Z]$ of a subset $Z\subseteq\partial{\mathrm{Y}}$ is the union of all []delimited-[][$\cdot$][ ⋅ ]-classes of $\xi\in Z$. We say that $\xi$ is minimal if $[\xi]=\{\xi\}$.
Partitions on the boundary might look unnatural. It is, however, amount to descending the action to the quotient $[\partial{\mathrm{Y}}]$ of $\partial{\mathrm{Y}}$ induced by the relation []delimited-[][$\cdot$][ ⋅ ] via the map:
$\xi\in\partial{\mathrm{Y}}\quad\longmapsto\quad[\xi]\in[\partial{\mathrm{Y}}]$
so the open subsets of $[\partial{\mathrm{Y}}]$ are precisely the images of saturated open sets $U$ in $\partial{\mathrm{Y}}$ with $U=[U]$. In general, $[\partial{\mathrm{Y}}]$ may not be Hausdorff or $T_{1}$.
Say that []delimited-[][$\cdot$][ ⋅ ] is a closed partition if $x_{n}\to\xi\in\partial{\mathrm{Y}}$ and $y_{n}\to\eta\in\partial{\mathrm{Y}}$ are two sequences with $[x_{n}]=[y_{n}]$, then $[\xi]=[\eta]$. Equivalently, the relation $\{(\xi,\eta):[\xi]=[\eta]\}$ is a closed subset in $\partial{\mathrm{Y}}\times\partial{\mathrm{Y}}$, so the quotient space $[\partial{\mathrm{Y}}]$ is Hausdorff.
Lemma 3.1.
Assume that the partition []delimited-[][$\cdot$][ ⋅ ] is closed. For any given $Z\subseteq\partial{\mathrm{Y}}$, we have $\overline{[Z]}\subseteq[\overline{Z}]$. In particular, the quotient map is a closed map: if $Z$ is a closed subset, then $[Z]$ is closed.
Proof.
Let $w_{n}\in[Z]$ tend to $\xi\in\partial{\mathrm{Y}}$. Then there exists $z_{n}\in Z$ with $z_{n}\in[w_{n}]$, and for $\partial{\mathrm{Y}}$ is compact, let $\eta\in\bar{Z}$ be any accumulation point of $z_{n}$. As []delimited-[][$\cdot$][ ⋅ ] is closed, we have $[\xi]=[\eta]$. Thus, we proved $\xi\in[\overline{Z}]$ and the proof is complete. ∎
Recall that a point $\xi$ is an accumulation point of a sequence $\{x_{n}\in\mathrm{Y}:n\geq 1\}$, if there exists a subsequence denoted by $y_{n}$ so that $y_{n}$ converges to $\xi$. We say that $x_{n}$ tends (resp. accumulates) to $[\xi]$ if the limit point (resp. any accumulate point) is contained in $[\xi]$. This implies that $[x_{n}]$ tends or accumulates to $[\xi]$ in the quotient space $[\partial{\mathrm{Y}}]$. So, an infinite ray $\gamma$ terminates at a point in $[\xi]\in\partial{\mathrm{Y}}$ if any sequence of points on $\gamma$ accumulates in $[\xi]$.
Let $X$ be a subset in $\mathrm{Y}$. The limit set of $X$, denoted by $\Lambda X$, is the topological closure of $X$ in $\partial{\mathrm{Y}}$.
Convention 3.2.
We write “$[Z]\subseteq\partial{\mathrm{Y}}$” to emphasize $[Z]$ as a subset in $\partial{\mathrm{Y}}$, and “$[Z]\subseteq[\partial{\mathrm{Y}}]$” as a subset in $[\partial{\mathrm{Y}}]$. In the latter case, the convergence to $[\xi]$ makes the usual sense in the quotient topology.
3.2. Assumptions A and B
The first two assumptions relate the contracting property of subsets in the space to the boundary. Recall that a sequence of subsets $X_{n}$ is escaping if $d(o,X_{n})\to\infty$ for some basepoint $o\in\mathrm{Y}$.
Assump A.
Any contracting geodesic ray $\gamma$ accumulates into a closed subset $[\xi]$ for some $\xi\in\partial{\mathrm{Y}}$. Moreover, any sequence of $y_{n}\in\mathrm{Y}$ with escaping projections $\pi_{\gamma}(y_{n})$ tends to $[\xi]$.
Note that, by assumption, $[\xi]$ is requested to be a closed subset.
Lemma 3.3.
Under Assump A, any contracting quasi-geodesic ray $\gamma$ and its finite neighborhood accumulates in only one []delimited-[][$\cdot$][ ⋅ ]-class, denoted by $[\gamma^{+}]\in[\partial{\mathrm{Y}}]$.
Proof.
By Morse property in Lemma 2.3, a contracting quasi-geodesic ray is contained in a finite neighborhood of a contracting geodesic ray. ∎
Let $\gamma:\mathbb{R}\to\mathrm{Y}$ be a contracting quasi-geodesic, with the positive and negative half rays denoted by $\gamma^{+},\gamma^{-}$ respectively. Following [GP13], $\gamma$ is called horocycle if $\gamma^{+},\gamma^{-}$ accumulate into the same []delimited-[][$\cdot$][ ⋅ ]-class: $[\gamma^{+}]=[\gamma^{-}]$. This is opposite to a class of non-pinched points, which shall be consider in Assump C. If the points $[\xi]$ in Assump A are non-pinched, we show that the only possible way for $x_{n}$ tending to $[\xi]$ is that their projection to $\gamma$ is escaping. See Lemma 3.16 for a precise statement.
Definition 3.4.
Let $h\in\textrm{Isom}(\mathrm{Y})$ be a contracting isometry. The attracting fixed set $[h^{+}]$ and repelling fixed set $[h^{-}]$ in $\partial{\mathrm{Y}}$ are the locus of all accumulation points of $\{h^{n}o:n>0\}$ and $\{h^{-n}o:n>0\}$ respectively.
Abusing language, we also refer $[h^{-}],[h^{+}]$ as the fixed points of $h$, though they are usually non-singleton. Write $[h^{\pm}]:=[h^{+}]\cup[h^{-}]$.
Note that $h$ is not assumed to be in a discrete group, so the axis is defined as in (10) to be the path $\textrm{Ax}(h)=\cup_{n\in\mathbb{Z}}h^{n}[o,ho]$.
By Lemma 3.3, the sets $[h^{-}],[h^{+}]$ do not depend on $o\in\mathrm{Y}$. If $\gamma$ is a quasi-geodesic axis, $[h^{-}],[h^{+}]$ are also represented by the negative/positive half rays $\gamma^{+},\gamma^{-}$ (in the obvious parametrization).
Our second assumption deals with an escaping sequence of contracting subsets, which roughly says that their cones sub-converge to a boundary point. Recall that give $r\geq 0$, the $r$-cone of a subset $X\subseteq\mathrm{Y}$ relative to $o\in\mathrm{Y}$ is defined as follows:
$\Omega_{o}(X,r):=\{y\in\mathrm{Y}:d_{X}(o,y)\geq r\}$
Assump B.
Let $\{X_{n}\subseteq\mathrm{Y}:n\geq 1\}$ be an escaping sequence of $C$-contracting quasi-geodesics for some $C>0$. For given $o\in\mathrm{Y}$, consider the union $Z_{n}=X_{n}\cup\Omega_{o}(X_{n},10C)$. Then there exist a subsequence of $Z_{n}$ (still denoted by $Z_{n}$) and a point $\xi\in\partial{\mathrm{Y}}$ such that $Z_{n}$ accumulates into $[\xi]$:
  • Any convergent sequence of points $z_{n}\in Z_{n}$ has the limit point in $[\xi]$.
Remark 3.5.
In the above assumption, we could replace $Z_{n}=\Omega_{o}(X_{n},10C)$ with the following simpler defined set
$Z_{n}=\{y\in\mathrm{Y}:[o,y]\cap Z_{n}\neq\emptyset\}$
for hyperbolic and CAT(0) spaces by Lemma 5.4, and for Teichmüller space by Lemma 12.4, .
The following weaker assumption suffices in most results below, which asserts the sub-convergence of $X_{n}$ instead of the union with $10C$-cones.
Assump B’.
Let $\{X_{n}\subseteq\mathrm{Y}:n\geq 1\}$ be an escaping sequence of $C$-contracting quasi-geodesics for some $C>0$. Then there exists $\xi\in\partial{\mathrm{Y}}$ so that a subsequence of $X_{n}$ (still denoted by $X_{n}$) accumulates into $[\xi]$:
  • Any convergent sequence of points $x_{n}\in X_{n}$ has the limit point in $[\xi]$.
The next two results under Assump B are particularly useful in defining conical points later on. The first result says that if the escaping sequence of contracting subsets intersect a geodesic ray in a large diameter, then it converges to a boundary point without passing to subsequence.
Lemma 3.6.
Assume that $X_{n}$ is an escaping sequence of $C$-contracting quasi-geodesics with bounded projection by a constant $B$. Let $\gamma$ be a geodesic ray in $\mathrm{Y}$ so that $\|X_{n}\cap\gamma\|\geq 14C+B$ for each $n\geq 1$. Then under Assump B, $\gamma$ and the sequence $X_{n}$ accumulate in one []delimited-[][$\cdot$][ ⋅ ]-class, denoted by $[\gamma^{+}]$.
Proof.
According to Assump B, assume that the cones of a subsequence $X_{n_{i}}$ converges to a []delimited-[][$\cdot$][ ⋅ ]-class, denoted by $[\gamma^{+}]$. We first prove that any unbounded sequence $z_{n}$ of points on a geodesic ray $\gamma$ also converges to $[\gamma^{+}]$. Indeed, fixing $X_{n_{i}}$, we obtain from the assumption that $z_{n}\in\Omega_{o}(X_{n_{i}},10C)$ for any $n\gg 0$ and $o:=\gamma_{-}$. Thus, Assump B implies that $z_{n}\to[\gamma^{+}]$.
It now remains to prove that any subsequence of $X_{n}$ converges to $[\gamma^{+}]$. Fix $X_{n_{i}}$ and a point $x\in X_{n}\cap\gamma$ for any given $n\gg n_{i}$. By Lemma 2.4, we have $d_{X_{n_{i}}}(o,x)\geq\|N_{C}(X_{n_{i}})\cap\gamma\|-4C\geq 10C+B$. As $X_{n}$ projects to a set on $X_{n_{i}}$ with diameter at most $B$, we see that $d_{X_{n_{i}}}(o,y)\geq 10C$ for any $y\in X_{n}$, so $X_{n}\subseteq\Omega_{o}(X_{n_{i}},10C)$. By the above assumption, the cones of a subsequence $X_{n_{i}}$ converges to $[\gamma^{+}]$, hence $X_{n}$ does so. The proof is complete. ∎
Via Proposition 2.15, we can have the convergence of an admissible path and the sequence of contracting subsets in its saturation.
Lemma 3.7.
Let $\gamma$ be an $(L,B)$-admissible ray in $\mathrm{Y}$ where the saturation $\mathscr{F}(\gamma)$ is an escaping sequence of $C$-contracting quasi-geodesics with $B$-bounded projection. Then under Assump B, for $L\gg C+B$, the ray $\gamma$ and any infinite sequence $X_{n}\in\mathscr{F}(\gamma)$ accumulate in one []delimited-[][$\cdot$][ ⋅ ]-class, denoted by $[\gamma^{+}]$.
Proof.
By Proposition 2.15, if $L\gg C+B$, any geodesic $\alpha$ from $o:=\gamma^{-}$ to a point $y$ on $\gamma$ $r$-fellow travel the subpath of $\gamma$ between $o$ and $y$. Let $d(o,y)\to\infty$, we obtain a limiting geodesic ray $\alpha$ so that $\|X_{n}\cap\alpha\|\geq L-2r$. If $L$ is sufficiently large relative to $B+10C$, then the proof is completed by Lemma 3.6. ∎
In what follows, we apply Assump B’ to study the limit set of a discrete group action.
Consider a proper action $G\curvearrowright\mathrm{Y}$ as in (1). The limit set $\Lambda(Go)$ of a $G$-orbit $Go$ may depend on the basepoint $o\in\mathrm{Y}$, though $[h^{\pm}]$ is the locus of the limit set of $\langle h\rangle o$, which is independent of $o$ by Assump A.
Lemma 3.8.
Assume that $G$ contains a contracting element. Under Assump B’, for any two points $o,o^{\prime}\in\mathrm{Y}$, we have $[\Lambda(Go)]=[\Lambda(Go^{\prime})]$.
Proof.
Fix a contracting element $h$ with $C$-contracting axis $\textrm{Ax}(h)=E(h)\cdot o$. Let $g_{n}o\to\xi\in\Lambda(Go)$. After passage to a subsequence, it suffices to prove that $g_{n}o^{\prime}\to\eta\in[\xi]$. Indeed, consider the sequence $g_{n}o\in X_{n}:=g_{n}\textrm{Ax}(h)$. If $X:=X_{n}$ is eventually the same, then we are done by Lemma 3.3. Otherwise, assume that $X_{n}$ is escaping. For $r:=d(o,o^{\prime})$, $N_{r}(X_{n})$ is an escaping sequence of $(C+2r)$-contracting subsets containing $g_{n}o^{\prime}$ and $X_{n}$. By Assump B’, any convergent subsequence of $g_{n}o^{\prime}$ has the limit point in $[\xi]$. ∎
We now derive that the normal subgroup shares the limit set with the ambient group, which crucially uses Assump B instead of Assump B’.
Lemma 3.9.
Let $H<G$ be an infinite normal subgroup. Then under Assump A and Assump B, $[\Lambda(Ho)]=[\Lambda(Go)]$.
Proof.
By Lemma 2.11, we can choose a set of three independent contracting elements $F\subseteq H$. Let $g_{n}\in G$ such that $g_{n}o\to\xi\in\Lambda(Go)$. For each pair $(g_{n},g_{n}^{-1})$, after passage to subsequences, there exists $f\in F$ such that $h_{n}:=g_{n}fg_{n}^{-1}\in H$ satisfies Lemma 2.17, so $[o,h_{n}o]\cap N_{C}(g_{n}\textrm{Ax}(f))\neq\emptyset$. According to whether $X_{n}=g_{n}\textrm{Ax}(f)$ is escaping or not, we apply Assump B or Assump A to see that $h_{n}o$ accumulates into $[\xi]$. We thus obtain $\Lambda(Go)\subseteq[\Lambda(Ho)]$. The converse inclusion is obvious, so the lemma is proved. ∎
Remark 3.10.
(alternative proof) If $H$ contains a contracting element $h$ with minimal fixed point (i.e.: $[h^{+}],[h^{-}]$ being singletons), we provide an alternative argument under Assump B’.
Indeed, consider a sequence of elements $h_{n}\in\langle h\rangle$ so that $h_{n}o\to\xi\in\Lambda(Ho)$. On one hand, $[\Lambda(Go)]=[\overline{G\xi}]\subseteq[\Lambda(Go)]$ by Lemma 3.11 below. We now prove that $G\xi\subseteq[\Lambda(Ho)]$. For any given $g\in G$, as the conjugates $gh_{n}g^{-1}$ are contained in $H$ and $d(gh_{n}g^{-1}o,gh_{n}o)=d(o,go)$, Assump B’ implies that $gh_{n}g^{-1}o$ and $gh_{n}o$ converge to the same []delimited-[][$\cdot$][ ⋅ ]-class $[g\xi]$ which are singleton by assumption, so we obtain $g\xi\in\Lambda(Ho)$. Thus $G\xi\subseteq[\Lambda(Ho)]$ is proved.
The following lemma says that for any contracting element $h\in G$, the set of all $G$-translates of $[h^{\pm}]$ is dense in the limit set $[\Lambda(Go)]$ in the quotient $[\partial{\mathrm{Y}}]$.
Lemma 3.11.
Fix a contracting element $h\in G$ and a basepoint $o\in\mathrm{Y}$. Take any point $\xi\in\Lambda(E(h)o)$. Under Assump A and Assump B’, we have $\overline{G\xi}\subseteq\Lambda(Go)\subseteq[\overline{G\xi}]$. In particular, $[\Lambda(Go)]=[\overline{G\xi}]$.
Remark 3.12.
If the partition []delimited-[][$\cdot$][ ⋅ ] is closed, we can obtain $[\overline{G\xi}]=[\overline{G\eta}]$ from $[\xi]=[\eta]$. In this case, the above conclusion works for any $\xi\in[h^{\pm}]$.
Proof.
Let $\eta\in\Lambda(Go)$, and $x_{n}:=g_{n}o\to\eta$ a sequence of points. Set $\gamma_{n}:=g_{n}\textrm{Ax}(h)$. Up to taking a subsequence of $\gamma_{n}$, we consider two cases as follows.
Case 1. Assume that $d(o,\gamma_{n})\to\infty$ as $n\to\infty$. By Assump B’, as $x_{n}\in\gamma_{n}\to\eta$, any convergent sequence of $y_{n}\in\gamma_{n}$ tends to $\eta^{\prime}\in[\eta]$. By assumption, for each $n\geq 1$, the point $g_{n}\xi\in g_{n}[h^{\pm}]$ is an accumulation point of the set $\gamma_{n}$. The compactification $\mathrm{Y}\cup\partial{\mathrm{Y}}$ is assumed to be metrizable, so let us fix a compatible metric $\delta$. Choose $y_{n}\in\gamma_{n}$ so that $\delta(y_{n},g_{n}\xi)\leq 1/n$. Thus, $g_{n}\xi\to\eta^{\prime}\in[\eta]$.
Case 2. $\gamma:=\gamma_{n}$ for all $n\geq 1$, as the $G$-translated axis of $h$ forms a discrete sequence by the proper action. Note that $\gamma$ is a translated axis of $h$ on which $E(h)$ acts cocompactly. As a non-elementary $G$ contains infinitely many distinct conjugates of $h$, we can thus choose a sequence of their axes $\alpha_{n}$ (i.e. copies of $\gamma$) so that $\pi_{\gamma}(\alpha_{n})$ is uniformly close to $x_{n}$ independent of $n$. Pick up any sequence $y_{n}\in\alpha_{n}$. Since $d(o,x_{n})\to\infty$ we have $y_{n}\to[\eta]$ by Assump A. Considering the escaping sequence $\alpha_{n}$, we are therefore reduced to the setup of the Case 1, so the proof of the lemma is finished. ∎
We now prove the double density of fixed points of contracting elements. Recall that two contracting elements $h,k$ (in a discrete group $G$) are independent if their $G$-translated axes have bounded projection / intersection.
Lemma 3.13.
Let $h,k\in G$ be two independent contracting elements. Then under Assump A, for any large $n\gg 0$, the element $g_{n}:=h^{n}k^{n}$ is a contracting element. Moreover, there exist $\xi_{n}\in[g_{n}^{+}]$ and $\eta_{n}\in[g_{n}^{-}]$ such that $\xi_{n},\eta_{n}$ tend to $[h^{+}],[k^{-}]$ respectively.
Remark 3.14.
We explain some subtlety about existence of $\xi_{n}\in[g_{n}^{+}]$ and $\eta_{n}\in[g_{n}^{-}]$. In the proof, it suffices to choose $\xi_{n}$ and $\eta_{n}$ to be accumulation points of $\langle g_{n}\rangle o$. However, the []delimited-[][$\cdot$][ ⋅ ]-locus of $[g_{n}^{+}]$ may contain points which are not accumulation points of $\langle g_{n}\rangle o$.
Proof.
By assumption, $\textrm{Ax}(h)$ and $\textrm{Ax}(k)$ have $B$-bounded projection for some $B>0$. Consider the admissible path relative to $\mathbb{X}=\{g\textrm{Ax}(k):g\in G\}$. For large enough $n>0$, the element $g_{n}:=h^{n}k^{n}$ is contracting: indeed, its orbital points form a bi-infinite $(L,B)$-admissible path $\gamma_{n}$ as follows:
$\gamma_{n}=\bigcup_{i\in\mathbb{Z}}g_{n}^{i}\big{(}[o,h^{n}o]h^{n}[o,k^{n}o]\big{)}$
where $L:=d(o,k^{n}o)$. By [Yan19, Proposition 2.9], $\gamma_{n}$ is a uniformly contracting quasi-geodesic independent of $n$, so by definition, $g_{n}$ is a contracting element. The positive (resp. negative) ray $\gamma_{n}^{+}$ (resp. $\gamma_{n}^{-}$) denotes the half ray with positive (resp. negative) indices $i$ in $\gamma_{n}$.
Let $\xi_{n}\in[g_{n}^{+}]$ be an accumulation point of $\gamma_{n}^{+}$. We now only prove that $\xi_{n}$ tends to $[h^{+}]$. The case for repelling points $\eta_{n}\to[k^{-}]$ is symmetric.
If denote $\alpha:=\bigcup_{i\in\mathbb{Z}}h^{i}[o,ho]$ and $\beta:=\bigcup_{i\in\mathbb{Z}}k^{i}[o,ko]$, the positive ray of $\alpha$ accumulates into $[h^{+}]$, and the negative ray of $\beta$ into $[k^{-}]$. Since $\gamma_{n}$ has the overlap $[o,h^{n}o]$ with $\alpha$, any point $y_{n}$ on the positive ray $\gamma_{n}^{+}$ that is sufficiently far from $o$ projects to a fixed neighborhood of $h^{n}o$, whose radius depends only on the contracting constant. Applying Assump A with the contracting quasi-geodesic $\alpha$, we obtain $y_{n}\to[h^{+}]$ for $x_{n}\to[h^{+}]$. Since $y_{n}$ can be chosen arbitrarily close to $\xi_{n}$, we conclude that the attracting fixed point $\xi_{n}$ tends to $[h^{+}]$. ∎
It is convenient to understand the next result in the quotient space $[\partial{\mathrm{Y}}]$.
Lemma 3.15.
Let $\Lambda G\Join\Lambda G$ be the set of distinct []delimited-[][$\cdot$][ ⋅ ]-pairs in $[\Lambda(Go)]$ with quotient topology defined in (2). Then under Assump A and Assump B’, the set of (ordered) fixed point pairs $([h^{+}],[h^{-}])$ of all contracting elements $h\in G$ with $[h^{+}]\neq[h^{-}]$ is dense in $\Lambda G\Join\Lambda G$.
Proof.
Let $\xi,\eta\in\Lambda(Go)$ so that $[\xi]\neq[\eta]$. Consider any []delimited-[][$\cdot$][ ⋅ ]-saturated open neighborhoods $[\xi]\subseteq U=[U]$ and $[\eta]\subseteq V=[V]$. By Lemma 3.11, there exist two contracting elements $h,k$ so that $[h^{+}]\subseteq U,[k^{-}]\subseteq V$. By Lemma 3.13, there exist contracting elements $g_{n}=h^{n}k^{n}$ for $n\gg 0$ such that $[g_{n}^{+}]\subseteq U$ and $[g_{n}^{-}]\subseteq V$. This implies that $([g_{n}^{+}],[g_{n}^{-}])$ converges to $([\xi],[\eta])$ in the quotient topology. ∎
3.3. Assumption C
The third assumption demands the boundary to contain enough interesting points. This is a non-trivial condition to get a useful theory, as the one-point compactification satisfies the assumptions (A) and (B).
Assump C.
Assume that the following set $\mathcal{C}$ of non-pinched points $\xi\in\partial{\mathrm{Y}}$ is non-empty. If $x_{n},y_{n}\in\mathrm{Y}$ are two sequence of points converging to $[\xi]$, then $[x_{n},y_{n}]$ is an escaping sequence of geodesic segments.
[…]
Under Assump C, we first give a converse statement to Assump A, and two useful corollaries of it.
Lemma 3.16.
Assume that a $C$-contracting geodesic ray $\gamma$ accumulates into $[\xi]$ by Assump A, where $\xi\in\mathcal{C}$ is some non-pinched point. If $x_{n}\in\mathrm{Y}\to[\xi]$, then $\pi_{\gamma}(x_{n})$ forms an unbounded sequence.
Proof.
Indeed, the contracting property of $\gamma$ implies that
$d(\pi_{\gamma}(x_{n}),[x_{n},y])\leq 2C$
for any $y\in\gamma$. Choose $y_{n}\in\gamma$ so that $y_{n}\to[\xi]$. If $\pi_{\gamma}(x_{n})$ stays bounded, then we found a non-escaping sequence of $[x_{n},y_{n}]$ so that $x_{n}\to[\xi]$ and $y_{n}\to[\xi]$. This contradicts Assump C. ∎
As a corollary, we obtain.
Lemma 3.17.
Given $\xi\in\mathcal{C}$, let $\alpha$ be a contracting geodesic ray ending at $[\xi]$. If $\beta$ is a geodesic ray also ending at $[\xi]$, then $\beta$ is contained in a finite neighborhood of $\alpha$.
Proof.
Take a sequence $y_{n}\in\beta$ tending to $[\xi]$, where $y_{0}=\beta^{-}$. By Lemma 3.16 for $\alpha$, $\pi_{\alpha}(y_{n})$ is escaping, then applying Assump A, any sequence of $x_{n}\in\pi_{\alpha}(y_{n})$ tends to $[\xi]$. By $C$-contracting property of $\alpha$, the segment $[\beta_{-},y_{n}]_{\beta}$ intersects the two $C$-balls at $x_{0}$ and $x_{n}$ respectively. As $d(o,x_{n})\to\infty$, we see that $\beta$ intersects $N_{C}(\alpha)$ in a unbounded set $\{x_{n}\}$, so the Morse property of $\alpha$ implies that $\beta$ is contained in a finite neighborhood of $\alpha$. ∎
Another corollary is that non-pinched boundary points at which a contracting geodesic ray terminates are visual. This is a baby version of the more complicated Lemma 4.11.
Lemma 3.18.
Let $\alpha,\beta$ be two $C$-contracting geodesic rays ending at $[\xi]\neq[\eta]\subseteq\mathcal{C}$ respectively. Then there exists a bi-infinite geodesic $\gamma$ such that two half ways eventually lies in a $\tilde{C}$-neighborhood of $\alpha$ and $\beta$, where $\tilde{C}$ depends only on $C$.
Proof.
Let $y_{n}\in\beta$ tend to $[\eta]$. By Lemma 3.16, as $[\xi]\neq[\eta]$, the projection $\pi_{\alpha}(y_{n})$ stays in a bounded set $A\subseteq\alpha$. Similarly, for any $x_{n}\in\alpha\to[\xi]$, $\pi_{\alpha}(y_{n})$ lies in a bounded set $B\subseteq\beta$. Thus, $[x_{n},y_{n}]$ intersects $N_{C}(A)$ and $N_{C}(B)$ by the contracting property. By Ascoli-Arzela Lemma, we could extract a limiting bi-infinite geodesic $\gamma$ from the sequence of $[x_{n},y_{n}]$, so that the two half rays of $\gamma$ lies in a finite neighborhood of $\alpha$ and $\beta$. ∎
The set $\mathcal{C}$ of non-pinched points allows us to introduce non-pinched contracting elements, which are crucial in the following discussion.
A contracting isometry $h\in\textrm{Isom}(\mathrm{Y})$ is called non-pinched if their fixed points $[h^{+}],[h^{-}]$ are contained in $\mathcal{C}$; otherwise, it is pinched.
Note that $h^{n}o\to[h^{+}]$ and $h^{-n}o\to[h^{-}]$ as $n\to\infty$. Thus, it follows that $[h^{+}]\neq[h^{-}]$ for a non-pinched $h$. It turns out that distinct fixed points property characterizes the non-pinched contracting elements.
$\left[h^{+}\right]=h^{-m}\left[h^{+}\right]$$\left[h^{-}\right]$$A\mathrm{x}(h)$$x_{n}$$y_{n}$$\pi_{\textrm{Ax}(h)}(y_{n})=w_{n}$$h^{-m}w_{n}$$h^{-m}x_{n}$$h^{-m}y_{n}$
Figure 6. Disjoint fixed points are non-pinched by Lemma 3.19.
Lemma 3.19.
Let $h\in\textrm{Isom}(\mathrm{Y})$ be a contracting isometry so that $[h^{-}]\cap[h^{+}]=\emptyset$. Then $h$ is non-pinched.
Proof.
We shall show that if $[h^{+}]$ is pinched, then $[h^{-}]=[h^{+}]$. By definition, as $[h^{+}]$ is pinched, there exist $x_{n},y_{n}\in\mathrm{Y}$ so that $x_{n},y_{n}\to[h^{+}]$ and $\sup_{n\geq 1}d(o,[x_{n},y_{n}])\leq M<\infty$ for some $M$. The geodesic $\alpha_{n}:=[x_{n},y_{n}]$ has length tending to $\infty$. The same property holds for the pair $(h^{m}x_{n},h^{m}y_{n})$ for any $m\neq 0$, since $h$ fixes $[h^{+}]$.
Here is the key observation. Note that $\alpha_{n}$ intersects a $M$-neighborhood of $\textrm{Ax}(h)$. So if we project the two endpoints $x_{n},y_{n}$ to $u_{n},v_{n}$ on $\textrm{Ax}(h)$ respectively, it follows by Lemma 2.2 that $d(u_{n},\alpha_{n}),d(v_{n},\alpha_{n})\leq C$ for any $n\gg 0$. We claim that there exists $D>0$ so that $d(o,w_{n})\leq D$ for some $w_{n}\in\{u_{n},v_{n}\}$ and each $n\geq 1$. Indeed, if not, we have $d(o,\{u_{n},v_{n}\})\to\infty$ as $n\to\infty$. Recalling that $\sup_{n\geq 1}d(o,\alpha_{n})\leq M$ for some $M<\infty$, this implies that the geodesic segment $\alpha_{n}$ stays within a fixed finite neighborhood (depending on $C,M$) of the subpaths $[u_{n},o]_{\textrm{Ax}(h)}$ and $[o,v_{n}]_{\textrm{Ax}(h)}$ of the quasi-geodesic $\textrm{Ax}(h)$. As $n\to\infty$ and $\ell(\alpha_{n})\to\infty$, this contradicts that $\alpha_{n}$ is a geodesic segment. Thus, $d(o,w_{n})\leq D$ for some $w_{n}\in\{u_{n},v_{n}\}$. See Fig. 6.
For any fixed large $m>0$, we apply $h^{-m}$ to the sequence $(x_{n},y_{n})$: $h^{-m}x_{n},h^{-m}y_{n}\to[h^{+}]$ as $n\to\infty$, and $d(h^{-m}w_{n},h^{-m}o)\leq D$. Letting $m\to\infty$, we can choose a sequence of points $z_{m}$ in $\{h^{-m}x_{n},h^{-m}y_{n}:n\geq 1\}$, so that $z_{m}$ tends to $[h^{+}]$, and $z_{m}$ projects into the $D$-neighborhood of $h^{-m}o$. Assump A shows that $z_{m}$ tends to $[h^{-}]$ as $m\to\infty$. Hence, $[h^{+}]=[h^{-}]$ follows. This concludes the proof of lemma. ∎
Unless explicitly stated, the statements from now on until the end of this section are under Assump A and Assump C, but without Assump B.
As expected, we deduce below that two non-pinched contracting isometries have either the same or disjoint fixed points. Recall that if $h$ is not assumed to be in a discrete group, the axis is defined as in (10) to be the path $\textrm{Ax}(h)=\cup_{n\in\mathbb{Z}}h^{n}[o,ho]$.
Lemma 3.20.
Let $h,k\in\textrm{Isom}(\mathrm{Y})$ be two contracting isometries so that $\textrm{Ax}(h)$ and $\textrm{Ax}(k)$ have bounded intersection. If $h$ is non-pinched, then $[h^{\pm}]\cap[k^{\pm}]=\emptyset$.
Proof.
Up to taking inverses, assume by contradiction that $\xi\in[h^{+}]\cap[k^{+}]$.
Take a sequence of points $x_{n}\in\textrm{Ax}(h)\to\xi$. By $C$-contracting property, we have $[o,x_{n}]$ intersects the $C$-ball centered at $\pi_{\textrm{Ax}(k)}(x_{n})$. Let us take $u_{n}\in[o,x_{n}]$ so that $d(u_{n},\textrm{Ax}(k))\leq C$. If $\{u_{n}\}$ is unbounded, the infinite intersection of $N_{C}(\textrm{Ax}(h))$ and $N_{C}(\textrm{Ax}(k))$ then follows from their Morse property, contradicting the hypothesis. Thus, $\{u_{n}\}$ is a bounded set.
Let $y_{n}\in\textrm{Ax}(h)$ tend to $\xi$. The contracting property of $\textrm{Ax}(k)$ implies that $[x_{n},y_{n}]$ intersects the $C$-neighborhood of $\pi_{\textrm{Ax}(k)}(x_{n})$, that is contained in a fixed compact ball by boundedness of $\{u_{n}\}$. This contradicts Assump C for $[\xi]\subseteq\mathcal{C}$. Thus the lemma is proved. ∎
Corollary 3.21.
Let $h,k$ be two non-pinched contracting elements in a discrete group $G$. Then either $[h^{\pm}]\cap[k^{\pm}]=\emptyset$ or $[h^{\pm}]=[k^{\pm}]$.
Proof.
Assume to the contrary that $[h^{+}]=[k^{+}]$ and $[h^{-}]\neq[k^{-}]$. By Lemma 3.20, $\textrm{Ax}(h)$ has infinite intersection with $\textrm{Ax}(k)$. We shall get a contradiction from the proper action. We present the proof for the convenience of the reader.
As $[h^{-}]\neq[k^{-}]$, the Morse property of $\textrm{Ax}(h)$ and $\textrm{Ax}(k)$ shows that $\{h^{n}o:n\geq 1\}$ and $\{k^{n}o:n\geq 1\}$ are $R$-close for some finite $R>0$. In particular, there are two unbounded sequences of integers $n_{i},m_{i}$ so that $d(h^{n_{i}}o,k^{m_{i}}o)\leq R$. The proper action implies $\{h^{-n_{i}}k^{m_{i}}\}$ is a finite set. Passing to a subsequence, we may assume $\{h^{-n_{i}}k^{m_{i}}=\{h^{-n_{j}}k^{m_{j}}$ for any $i\neq j$, and thus $h^{n_{i}-n_{j}}=k^{m_{j}-m_{i}}$ where $n_{i}\neq n_{j}$ and $m_{i}\neq m_{j}$. As a consequence, we obtain $\textrm{Ax}(h)=\textrm{Ax}(k)$, contradicting the assumption $[h^{-}]\neq[k^{-}]$. ∎
Note, however, that two independent pinched elements may share the same fixed point.
Example.
Let us consider the free group $\mathbb{F}_{3}=\mathbb{Z}\star\mathbb{F}_{2}$ of rank 3, which is hyperbolic relative to the factor $\mathbb{F}_{2}=\langle g,h\rangle$. The Bowditch boundary is then obtained from the visual boundary ($\simeq$ Cantor set) of $\mathbb{F}_{3}$ by collapsing equivariantly the limit set of $\mathbb{F}_{2}$ to one point. It is easy to verify that Bowditch boundary is a convergence boundary, where the two pinched contracting elements $g,h$ are independent, but share the same unique fixed point.
3.4. Shortest projection for boundary points
We introduce two ways of defining the shortest projection points of a boundary points to a contracting subset. Before describing the definitions, let us repeat the convention that given a quasi-geodesic $\gamma$, we denote by $\gamma^{+}$ either a positive half ray of $\gamma$ or the endpoint in $\partial{\mathrm{Y}}$ determined by Assump A. In the later use, we often write $[\gamma^{+}]$ for the []delimited-[][$\cdot$][ ⋅ ]-class, since any unbounded sequence on the half ray $\gamma^{+}$ is assumed to accumulate into the same []delimited-[][$\cdot$][ ⋅ ]-class.
In what follows, let $\gamma$ be a $C$-contracting bi-infinite quasi-geodesic for $C>0$. We shall define a shortest projection map
$\pi_{\gamma}:\partial{\mathrm{Y}}\setminus[\gamma^{\pm}]$ $\longrightarrow\gamma$
$\xi$ $\longmapsto\pi_{\gamma}(\xi)$
with the following properties:
  1. (1)
    $\pi_{\gamma}(\xi)$ is a bounded set on $\gamma$ (possibly of diameter depending on $\xi$).
  2. (2)
    the map extends the shortest projection map $\pi_{\gamma}:\mathrm{Y}\to{\gamma}$ in the following sense: if $y_{n}\in\mathrm{Y}\to\xi$, then $\pi_{\gamma}(y_{n})\subseteq\pi_{\gamma}(\xi)$ for all $n\gg 0$.
  3. (3)
    the map is coarsely equivariant: $\pi_{g\gamma}(g\xi)=g\pi_{\gamma}(\xi)$ for any $g\in\textrm{Isom}(\mathrm{Y})$.
We first define the map for non-pinched points $\xi$ in $\mathcal{C}$ as in Assump C. In this case, the diameter of $\pi_{\gamma}(\xi)$ is bounded by a constant depending on the contracting constant of $\gamma$ but independent of $\xi$.
Lemma 3.22.
For any $\xi\in\mathcal{C}\setminus[\gamma^{\pm}]$, there exists a bounded set denoted by $\pi_{\gamma}(\xi)$ of diameter at most $2C$ with the following property.
Let $y_{n}\in\mathrm{Y}\to[\xi]$. Then $\pi_{\gamma}(y_{n})$ is contained in $\pi_{\gamma}(\xi)$ for all $n\gg 0$.
Proof.
We claim that for some $N>0$, the following set has diameter at most $C$:
$U:=\{\pi_{\gamma}(y_{n}):n\geq N\}$
Indeed, if not, there exists two subsequences $\{x_{n}\},\{z_{n}\}$ of points from $\{y_{n}\}$ such that $d_{\gamma}(x_{n},z_{n})>C$ for each $n\geq 1$. By Lemma 2.2, $[x_{n},z_{n}]$ intersects $N_{C}(\pi_{X}(x_{n}))$. By assumption, $[\xi]$ is disjoint with $[\gamma^{\pm}]$, so all projection points $\{\pi_{\gamma}(y_{n}):n\geq 1\}$ are contained in a fixed ball $Z$ of finite radius by Assump A. Consequently, $[x_{n},z_{n}]$ intersects $N_{C}(Z)$, but $x_{n},z_{n}\to[\xi]\subseteq\mathcal{C}$, so we get a contradiction with Assump C. Setting $\pi_{\gamma}(\xi):=N_{C}(U)$ completes the proof. ∎
The following fact obtained in the above proof will be used later on.
Corollary 3.23.
Let $x_{n},y_{n}\in\mathrm{Y}$ be two sequences of points tending to $[\xi]\subseteq\mathcal{C}\setminus[\gamma^{\pm}]$. Then $d_{\gamma}(x_{n},y_{n})\leq C$ for all $n\gg 0$.
We now define, without Assump C, the shortest projection $\pi_{\gamma}(\xi)$ for any boundary point $\xi\in\partial{\mathrm{Y}}\setminus[\gamma^{\pm}]$.
Lemma 3.24.
Assume that $K\subseteq\partial{\mathrm{Y}}\setminus[\gamma^{+}]$ is a closed subset which may contain $[\gamma^{-}]$. Then there is a negative half ray $\gamma^{-}$ (i.e.: a restriction to some $(-\infty,c)$ of $\gamma:(-\infty,\infty)\to\mathrm{Y}$) depending on $K$ such that for any $\xi\in K\setminus[\gamma^{-}]$ and any $y_{n}\in\mathrm{Y}\to\xi$, we have $\pi_{\gamma}(y_{n})\subseteq\gamma^{-}$.
In particular, if $K\cap[\gamma^{\pm}]=\emptyset$, then there exists a bounded subset $Z$ of $\gamma$ such that for any $y_{n}\in\mathrm{Y}\to\xi\in K$, $\pi_{\gamma}(y_{n})$ is contained in $Z$.
Proof.
We proceed by way of contradiction. Let us assume that for some $\xi\in K$ and a sequence $y_{n}\in\mathrm{Y}\to\xi$, so that $\{\pi_{\gamma}(y_{n}):n\geq 1\}$ is not contained in any negative half way of $\gamma$. By Assump A, (a subsequence of) $\pi_{\gamma}(y_{n})$ is escaping and tends to $[\gamma^{+}]$. Let us fix a metric $\delta$ compatible with the topology $\mathrm{Y}\cup\partial{\mathrm{Y}}$. We thus have $\delta(y_{n},[\gamma^{+}])\to 0$ on one hand, and $\delta(y_{n},\xi)\to 0$ on the other hand. This contradicts the assumption $\delta(K,[\gamma^{+}])>0$. ∎
As a consequence, we can define a projection $\pi_{\gamma}(\xi)$ to $\gamma$ for any boundary point $\xi\in\partial{\mathrm{Y}}\setminus[\gamma^{\pm}]$ as the bounded set $Z$ for a sufficiently small compact neighborhood $K$ of $\xi$.
Lemma 3.25.
For any bounded subset $Z\subseteq\gamma$, the preimage of $Z$ under $\pi_{\gamma}:\mathrm{Y}\cup\partial{\mathrm{Y}}\setminus[\gamma^{\pm}]\to\gamma$ has the closure disjoint with $[\gamma^{\pm}]$.
Proof.
Let $K$ be the topological closure of $\pi_{\gamma}^{-1}(Z)$ in $\partial{\mathrm{Y}}\cup\mathrm{Y}$. If $K\cap[\gamma^{\pm}]\neq\emptyset$, let $\xi_{n}\in\pi_{\gamma}^{-1}(Z)\to[\gamma^{+}]$ for definiteness. By definition of $\pi_{\gamma}$, choose $x_{n}\in\mathrm{Y}$ in a small neighborhood of $\xi_{n}$ such that $x_{n}\to[\gamma^{+}]$, and $\pi_{\gamma}(x_{n})\cap Z\neq\emptyset$. Choose $y_{n}\in\gamma$ so that $y_{n}\to[\gamma^{+}]$. By the $C$-contracting property of $\gamma$, each $[x_{n},y_{n}]$ intersects the $C$-neighborhood of $\pi_{\gamma}(x_{n})$, and thus of a compact subset as $\pi_{\gamma}(x_{n})\cap Z\neq\emptyset$. This contradicts Assump C with the assumption $[\gamma^{\pm}]\subseteq\mathcal{C}$. ∎
Lemma 3.26.
Let $\xi\in\mathcal{C}$. Then for any $\eta\in[\xi]$, we have $d_{\gamma}(\xi,\eta)\leq 6C$.
Proof.
Assume by contradiction that $d_{\gamma}(\xi,\eta)>6C$. If $x_{m}\in\mathrm{Y}\to\xi$ and $z_{m}\in\mathrm{Y}\to\eta$, the triangle inequality for $d_{\gamma}$ with Lemma 3.22 implies
$\forall m\gg 1:\quad d_{\gamma}(x_{m},z_{m})\geq d_{\gamma}(\xi,\eta)-4C\geq 2C$
so by Lemma 2.2, $[x_{m},z_{m}]$ intersects the $C$-ball around $\pi_{\gamma}(\xi)$. As $\pi_{\gamma}(\xi)$ is a fixed subset of diameter at most $2C$, the non-escaping sequence of $[x_{m},z_{m}]$ contradicts the assumption of $\xi\in\mathcal{C}$. So the lemma is proved. ∎
3.5. North–South dynamics of non-pinched contracting elements
Let $h\in\textrm{Isom}(\mathrm{Y})$ be a non-pinched contracting isometry, where the fixed points $[h^{\pm}]$ are non-pinched. This is equivalent to have $[h^{-}]\neq[h^{+}]$ by Lemma 3.19. We shall prove that the action of $\langle h\rangle$ on the complement to $[h^{\pm}]$ in $\partial{\mathrm{Y}}$ has so-called North–South dynamics.
Such type of dynamics can be found in many contexts, but surprisingly, our proof is quite general and only uses the above Assump A and Assump C. This property shall be often used later on.
Lemma 3.27 (NS dynamics).
The action of $\langle h\rangle$ on $\partial{\mathrm{Y}}\setminus[h^{\pm}]$ has the North–South dynamics:
for any two open sets $[h^{+}]\subseteq U$ and $[h^{-}]\subseteq V$ in $\partial{\mathrm{Y}}$, there exists an integer $n>0$ such that $h^{n}(\partial{\mathrm{Y}}\setminus V)\subseteq U$ and $h^{-n}(\partial{\mathrm{Y}}\setminus U)\subseteq V$.
Proof.
Since $h$ is a contracting element, the path $\gamma:=\bigcup_{i\in\mathbb{Z}}h^{i}[o,ho]$ is a contracting quasi-geodesic. Let $K$ be the set of points $x\in\mathrm{Y}$ such that $[o,ho]\cap\pi_{\gamma}(x)\neq\emptyset$, so the following holds
$\bigcup_{g\in G}gK=\mathrm{Y}.$
Denote by $\widetilde{K}$ the topological closure of $K$ in $\mathrm{Y}\cup\partial{\mathrm{Y}}$, and $\Lambda K:=\partial{\mathrm{Y}}\cap\widetilde{K}$.
Let $\pi_{\gamma}:\partial{\mathrm{Y}}\setminus[h^{\pm}]\to\gamma$ be the shortest projection map given by Lemma 3.24. By definition, $[h^{\pm}]=[\gamma^{\pm}]$. By Lemma 3.25 we have $\widetilde{K}\cap[h^{\pm}]=\emptyset$.
Observe that
(17) $\partial{\mathrm{Y}}\cup\mathrm{Y}=\Big{(}\bigcup_{g\in G}g\widetilde{K}\Big{)}\bigcup{[h^{\pm}]}.$
Indeed, if $\xi\in\partial{\mathrm{Y}}\setminus[h^{\pm}]$, any sequence of points $x_{n}\in\mathrm{Y}\cup\partial{\mathrm{Y}}\to\xi$ has bounded projection to $\gamma$ by Assump A. Then $\{x_{n}\}$ are contained in finitely many copies of $K$, so the inclusion $\subseteq$ follows. The other direction follows by definition $\widetilde{K}$ and $\cup_{g\in G}gK=\mathrm{Y}$.
By a similar argument, Lemma 3.24 also implies that $\pi_{\gamma}(\widetilde{K})$ lies in a finite neighborhood of $[o,ho]$. Consequently, the projection map $\pi_{\gamma}$ coarsely commutes two actions of $\langle h\rangle$ on $\mathrm{Y}\cup\partial{\mathrm{Y}}\setminus[h^{\pm}]$ and on $\gamma$: for any $h\in G$ and $\eta\in\partial{\mathrm{Y}}\setminus[h^{\pm}]$,
(18) $\pi_{\gamma}(h\eta)\simeq h\pi_{\gamma}(\eta)$
where $\simeq$ means one of the two sets being contained into a uniform neighborhood of the other.
As $\langle h\rangle$ acts by translation on $\gamma$, we see from (18) that $\langle h\rangle$ acts properly on $\mathrm{Y}\cup\partial{\mathrm{Y}}\setminus[h^{\pm}]$: for any compact subset $L\subseteq\partial{\mathrm{Y}}\setminus[h^{\pm}]$, the set $\{g\in G:gL\cap L\neq\emptyset\}$ is finite. Moreover, $\{h^{i}L:i\in\mathbb{Z}\}$ is a locally finite family of closed subsets, so the union of any subfamily is a closed subset in $\mathrm{Y}\cup\partial{\mathrm{Y}}\setminus[h^{\pm}]$.
If $W\subseteq\partial{\mathrm{Y}}\cup U$ is a compact neighborhood of $\widetilde{K}$ disjoint with $[h^{\pm}]$, then
$U_{N}:=\partial{\mathrm{Y}}\setminus\left(\bigcup_{i<N}h^{i}W\cup[h^{-}]\right)$
for $N\geq 1$ forms a neighborhood basis for $[h^{+}]$. Indeed, let $[h^{+}]\subseteq U$ be any open subset. By Lemma 3.24 applied to $U^{c}$, there exists $N>0$ such that $U^{c}\subseteq\bigcup_{i<N}h^{i}W\cup[h^{-}]$. That is to say, $U_{N}\subseteq U$.
Similarly, $V_{M}:=\partial{\mathrm{Y}}\setminus(\bigcup_{i>-M}h^{i}W\cup[h^{+}])$ for $M>0$ is a neighborhood basis for $[h^{-}]$. Taking large enough $n>N+M$, we obtain $h^{n}(\partial{\mathrm{Y}}\setminus V_{M})\subseteq U_{N}$, so the North–South dynamics follows. ∎
Let $\gamma=\textrm{Ax}(h)$ be the axis of the non-pinched contracting element. The following result is implicit in the proof.
Corollary 3.28.
For any $L\geq 0$ there exist open neighborhoods $[h^{-}]\subseteq V$ and $[h^{+}]\subseteq U$ in $\mathrm{Y}\cup\partial{\mathrm{Y}}$ with the following property. For any $x\in V\setminus[h^{-}]$ and $y\in U\setminus[h^{+}]$ we have
$d_{\gamma}(x,y)\geq L.$
Conversely, if $d_{\gamma}(x_{n},y_{n})\to\infty$ for $x_{n},y_{n}\in\mathrm{Y}\cup\partial{\mathrm{Y}}\setminus[h^{\pm}]$, we have $(x_{n},y_{n})$ converges to $([h^{-}],[h^{+}])$.
Proof.
Keep the notions as in the proof of Lemma 3.27. As $\widetilde{K}\subseteq W$ is a compact set disjoint with $[h^{\pm}]$, $\pi_{\gamma}(\widetilde{K})\subseteq\pi_{\gamma}(W)$ is bounded by Lemma 3.24. By (18), we can choose $U=U_{N}$ and $V=V_{M}$ for large $M,N\gg L$ so that $d_{\gamma}(x,y)>L$ for $x\in U,y\in V$. The converse direction also follows by the North–South dynamics. ∎
We remark that the following result could be used to give another proof of Shadow Lemma 6.4 (as given in [TY21]). It is recorded here only for further applications.
Lemma 3.29.
Let $F$ be a set of non-pinched contracting elements given by Lemma 2.17. For any $f\in F$ and $r\gg 0$, there exist open neighborhoods $U\subseteq V$ of the attracting fixed point $[f^{+}]$ so that for any $g\in G$, we have
$gU\subseteq\Pi_{o}^{F}(go,r)\subseteq gV.$
Proof.
Set $\gamma:=\textrm{Ax}(f)$. Choose a shortest projection point $w\in\pi_{g\gamma}(o)$ from $o$ to $g\gamma$, so $d(o,go)\geq d(o,w)$ for $go\in g\gamma$. Let $U$ be an open neighborhood of $[f^{+}]$ by Corollary 3.28 such that
$\forall z\in U\cap\mathrm{Y}:\quad d_{\gamma}(g^{-1}o,z)=d_{g\gamma}(o,gz)\gg d(o,fo)$
This is to say, the distance from $go\in g\gamma$ to the exit point of $[o,gz]$ at $N_{C}(g\gamma)$ is sufficiently larger than $d(o,fo)$. For the constant $r$ given by Lemma 2.10, $go$ is an $(r,f)$-barrier for $[o,gz]$. Thus, the open set $gU\cap\partial{\mathrm{Y}}$ is contained in $\Pi_{o}^{F}(go,r)$.
The definition of $z\in\Pi_{o}^{F}(go,r)$ implies $\|N_{r}(g\gamma)\cap[o,z]\|\geq\|Fo\|_{\min}$, so by Lemma 2.4, we have $d_{g\gamma}(o,z)\geq\|Fo\|_{\min}-4r$. For large enough $\|F\|\gg 0$, Corollary 3.28 gives an open neighborhood $V$ of $[f^{+}]$ such that $z\in gV$. That is, we proved that $\Pi_{o}^{F}(go,r)\subseteq gV$. ∎
By Lemma 3.8 (under Assump B’), the limit set $\Lambda(Go)$ is well-defined up to taking the []delimited-[][$\cdot$][ ⋅ ]-locus. One desirable property for limit sets is their topological minimality and uniqueness. We describe a typical situation where this happens. Recall that a point in $\partial{\mathrm{Y}}$ is called minimal if its []delimited-[][$\cdot$][ ⋅ ]-locus contains only one point.
Lemma 3.30.
Fix a non-pinched contracting element $h\in G$ with minimal fixed points. Let $\Lambda=\overline{\{gh^{\pm}:g\in G\}}$ be the closure of the fixed points of all conjugates. Then under Assump A, Assump B’ and Assump C,
  1. (1)
    $\Lambda$ is the unique, minimal, $G$-invariant closed subset.
  2. (2)
    $[\Lambda]=[\Lambda(Go)]$ for any $o\in\mathrm{Y}$.
  3. (3)
    The fixed point pairs $([g^{+}],[g^{-}])$ of all non-pinched contracting elements $g$ in $G$ are dense in the distinct pairs of $\Lambda\times\Lambda$ in the following sense:
    • For any $\xi\neq\eta\in\Lambda$, there exists $\xi_{n}\in[f_{n}^{+}],\eta_{n}\in[f_{n}^{-}]$ for some contracting element $f_{n}\in G$ so that $\xi_{n}\to\xi$ and $\eta_{n}\to\eta$.
Proof.
We only need to prove the assertion (1). The assertion $[\Lambda]=[\Lambda(Go)]$ follows by Lemma 3.11. Assuming assertion (1), the double density follows the same proof of Lemma 3.15 (without taking the []delimited-[][$\cdot$][ ⋅ ]-closure of fixed points of $h$).
Note that the conjugates of $h$ are also non-pinched with minimal fixed points. A non-elementary group $G$ thus contains infinitely many conjugates of $h$ with pairwise bounded intersection, so the set $Z$ of their fixed points are infinite by Lemma 3.20. Let $A$ be any $G$-invariant closed subset in $\partial{\mathrm{Y}}$. First of all, $A$ contains at least three points. Otherwise, if $G$ fixes a point or a pair of points, this contradicts the North–South dynamics of conjugates of $h$, as the set $Z$ of their fixed points is infinite. As $[h^{+}]=\{h^{+}\}$ is minimal, we can choose $\xi\in A\setminus\{h^{+}\}$. By Lemma 3.27, $h^{n}\xi\in A$ converges to $h^{+}$. This implies $h^{+}\in\Lambda$, since $A$ is closed and $G$-invariant. Thus, $A$ contains $Z$ and $\Lambda\subseteq A$ is proved. ∎
Corollary 3.31.
Assume that all contracting elements in $G$ are non-pinched with minimal fixed points. Let $\Lambda$ be as in Lemma 3.30. Then the set of the fixed point pairs of all contracting element in $G$ is dense in the distinct pairs of $\Lambda\times\Lambda$.
Remark 3.32.
The above two results are a generalization of the known results on CAT(0) groups with rank-1 elements ([Bal95, Theorem III.3.4][Ham09b]) and dynamically irreducible subgroups of mapping class groups ([MP89]). The applications to Roller and Gardiner–Masur boundary seem new, and are recorded in Lemmas 11.5 and 12.6. The assumption in Corollary 3.31 is satisfied for all these examples.
4. Conical and Myrberg points
In this section, we introduce the notion of conical and Myrberg limit points of a proper action with contracting elements on the convergence boundary. This shall be central in the study of the conformal density performed in Section §6.
We begin with some non-elementary assumption on the group action under consideration, throughout the remainder of this paper.
Let $G$ be a non-elementary group acting properly on a proper geodesic space $\mathrm{Y}$, as in (1), with a contracting element. Assume further that $\mathrm{Y}$ is equipped with a non-elementary convergence boundary $\partial{\mathrm{Y}}$: $G$ contains a non-pinched contracting element with fixed points being non-pinched in a sense in Assump C. Equivalently, this is plain to say that the repelling and attracting fixed points of some contracting element are distinct []delimited-[][$\cdot$][ ⋅ ]-classes in $\partial{\mathrm{Y}}$ (Lemma 3.19).
To define conical limit points, we need to choose three independent non-pinched contracting elements $F=\{f_{1},f_{2},f_{3}\}$ and form the $C$-contracting system for some $C\geq 0$ as in (3)
=$\mathscr{F}=\{g\textrm{Ax}(f):f\in F,g\in G\}$
Convention 4.1.
Fix a constant $B$ satisfying Lemma 2.16. The constant $L$ in Definition 2.12 of an admissible path is chosen sufficiently large with the following properties:
  1. (1)
    Any $(L,B)$-admissible path has $r$-fellow travel relative to $\mathscr{F}$, where the constant $r$ satisfies Proposition 2.15 and Lemma 2.17.
  2. (2)
    If $\alpha$ contains a $(r,f)$-barrier with $d(o,fo)>L$, then $\beta$ with endpoints at most $10C$-apart contains $(\hat{r},f)$-barrier. The constant $\hat{r}$ depends only on $r$ by Lemma 2.9.
  3. (3)
    An intersection $\alpha\cap N_{r}(X)$ of diameter at least $L$ for $X\in\mathscr{F}$ produces an $(r,f)$-barrier for any geodesic $\alpha$. See Lemma 2.10.
In particular, all the constants depend only on the contracting constant $C$ of $\mathscr{F}$.
4.1. Conical points using shadows
From different perspectives, we are going to present three definitions of conical limit points, which are quantitatively equivalent. We first give the main definition of conical points formulated using shadows.
Fix a basepoint $o\in\mathrm{Y}$, so everything we talk about below depends on $o$. First of all, define the usual cone and shadow:
$\Omega_{x}(y,r)\;:=\;\{z\in\mathrm{Y}:\exists[x,z]\cap B(y,r)\neq\emptyset\}$
and $\Pi_{x}(y,r)\subseteq\partial{\mathrm{Y}}$ be the topological closure in $\partial{\mathrm{Y}}$ of $\Omega_{x}(y,r)$.
We are actually working with a technically involving but very useful notion of partial cones, which are defined relative to $\mathscr{F}$.
Definition 4.2 (Partial cone and shadow).
For $x\in\mathrm{Y},y\in Go$, the $(r,F)$-cone $\Omega_{x}^{F}(y,r)$ is the set of elements $z\in\mathrm{Y}$ such that $y$ is a $(r,F)$-barrier for some geodesic $[x,z]$.
The $(r,F)$-shadow $\Pi_{x}^{F}(y,r)\subseteq\partial{\mathrm{Y}}$ is the topological closure in $\partial{\mathrm{Y}}$ of the cone $\Omega_{x}^{F}(y,r)$.
We give the first definition of a conical point. Recall that $\mathcal{C}\subseteq\partial{\mathrm{Y}}$ is the set of non-pinched boundary points in Assump C.
Definition 4.3.
A point $\xi\in\mathcal{C}$ is called $(r,F)$-conical point if for some $x\in Go$, the point $\xi$ lies in infinitely many $(r,F)$-shadows $\Pi_{x}^{F}(y_{n},r)$ for $y_{n}\in Go$. We denote by $\Lambda_{r}^{F}(Go)$ the set of $(r,F)$-conical points.
Remark 4.4.
  1. (1)
    If a geodesic ray $\gamma$ starting at $x\in Go$ contains infinitely many $(r,F)$-barriers $y_{n}\in Go$, then by definition, any accumulation points of $\gamma$ are $(r,F)$-conical.
  2. (2)
    According to our convention, $F$ consists of non-pinched contracting elements $f$. The fixed points $[f^{\pm}]$ of $f$ are the []delimited-[][$\cdot$][ ⋅ ]-locus of the accumulation points of the axes $\textrm{Ax}(f)$. It is obvious that the accumulation points of the axes $\textrm{Ax}(f)$ are contained in infinitely many shadows $\Pi_{x}^{F}(y_{n},r)$ over $y_{n}\in\textrm{Ax}(f)$, so are $(r,\{f,f^{-1}\})$-conical points for $r\gg 0$. Thus, fixed points $[f^{+}]$ are $(r,F)$-conical points.
By definition, $\Lambda_{r}^{F}(Go)$ is a $G$-invariant set, and depends on the choice of $o\in\mathrm{Y}$. It is the limit superior of the family of partial shadows:
(19) $\Lambda_{r}^{F}(Go)=\bigcup_{x\in Go}\Big{(}\limsup_{y\in Go}\Pi_{x}^{F}(y,r)\Big{)}.$
Recall that $\Lambda_{c}(Go)$ is the usual conical limit set given in (6), defined similarly as above but using the usual shadows $\Pi_{x}(y,r)$. Thus, $\Lambda_{r}^{F}(Go)\subseteq\Lambda_{c}(Go)$.
Let us restate the Definition 1.8 of conical points given in Introduction.
Definition 4.5.
A non-pinched point $\xi\in\mathcal{C}$ is called $(L,\mathscr{F})$-conical point for $L>0$ if for some $x\in\mathrm{Y}$, there exists a sequence of $X_{n}\in\mathscr{F}$ such that $d_{X_{n}}(x,\xi)\geq L$.
We clarify that these two definitions introduced so far are equivalent in the following quantitative sense.
Lemma 4.6.
Let $r,F$ be as above. There exist $L_{1},L_{2}>0$ depending on $r,F$ such that
  1. (1)
    An $(L_{1},\mathscr{F})$-conical point is $(r,F)$-conical;
  2. (2)
    An $(r,F)$-conical point is $(L_{2},\mathscr{F})$-conical.
Proof.
Let $\xi$ be an $L_{1}$-conical point in Definition 4.5 so there exists $X_{n}\in\mathscr{F}$ such that for each $n\geq 1$, we have
$d_{X_{n}}(x,\xi)\geq L_{1}$
Consider a convergent sequence of points $z_{m}\in\mathrm{Y}\to\xi$. By Lemma 3.22 we have $d_{X_{n}}(x,z_{m})\geq L_{1}-2C$ for any fixed $n\geq 1$ and for any $m\gg 0$. Choosing $L_{1}\gg\|Fo\|$ by Lemma 2.10, we have $[x,z_{m}]$ contains an $(r,f)$-barrier at $y_{n}\in\pi_{X_{n}}(x)$ for some $f\in F$. As $z_{m}\to\xi$, we have $\xi$ lies in the closure of $\Omega_{x}^{F}(y_{n},r)$, that is, the shadow $\Pi_{x}^{F}(y_{n},r)$. Thus, $\xi$ is $(r,F)$-conical.
Conversely, if $\xi\in\Pi_{x}^{F}(g_{n}o,r)$, choose a sequence of points $z_{m}\in\Omega_{x}^{F}(g_{n}o,r)$ tending to $\xi$ (as $m\to\infty$). There exists an $(r,F)$-barrier $y_{n}\in X_{n}=g_{n}\textrm{Ax}(f_{n})$ for some $f_{n}\in F$ so that
$\|[x,z_{m}]\cap N_{r}(X_{n})\|\geq\|Fo\|_{\min}$
Lemma 2.4 implies $d_{X_{n}}(x,z_{m})\geq L_{2}:=\|Fo\|_{\min}-4r$. Thus, $\xi$ is an $L_{2}$-conical point. The proof is complete. ∎
A point $\xi\in\partial{\mathrm{Y}}$ is called []delimited-[][$\cdot$][ ⋅ ]-visual (or simply visual): for any $x\in\mathrm{Y}$, there exists a geodesic ray starting at $x$ ending at $[\xi]$. We first prove that conical points are visual.
We remark that the next result uses crucially Assump B, which underlies the proof of Lemma 3.6.
Lemma 4.7.
Let $\xi\in\Lambda_{r}^{F}(Go)$ and $\hat{r}$ given by Lemma 2.9. For any $x\in\mathrm{Y}$, there exists a geodesic ray $\gamma$ from $x$, which accumulates into $[\xi]$. Moreover,
  1. (1)
    $\gamma$ contains infinitely many distinct $(\hat{r},F)$-barriers $y_{n}\in Go$, and $y_{n}\to[\xi]$.
  2. (2)
    There exists an $(L,B,\mathscr{F})$-admissible ray starting from $x$ and ending at $[\xi]$, where $L,B$ depending on $F,r$ are given by Lemma 2.16.
Proof.
By definition of $\xi\in\Lambda_{r}^{F}(Go)$, for each fixed $n\geq 1$, there exists a sequence of points $z_{n,m}\in\mathrm{Y}$ in the cone $\Omega_{x}^{F}(g_{n}o,r)$ for some $x_{0}\in Go$ and $g_{n}\in G$ such that $z_{n,m}$ tends to $\xi$ when $m\to\infty$. A geodesic segment $\alpha_{m}^{\prime}=[x_{0},z_{n,m}]$ contains the $(r,f_{n})$-barrier $y_{n}=g_{n}o$ for some $f_{n}\in F$: that is,
$d(g_{n}o,\alpha_{m}^{\prime}),\;d(g_{n}f_{n}o,\alpha_{m}^{\prime})\leq r.$
Given $x\in\mathrm{Y}$, consider a geodesic segment $\alpha_{m}=[x,z_{n,m}]$. We claim that $g_{n}o$ is (,^$(\hat{,}f_{n})$-barrier for $\alpha_{m}$, where $\hat{r}$ is given by Lemma 2.9. Indeed, we first note that $\alpha_{m}$ and $\alpha_{m}^{\prime}$ both intersect $N_{r}(g_{n}\textrm{Ax}(f_{n}))$. if $d(x,g_{n}\textrm{Ax}(f_{n}))\geq d(x_{0},x)$, we have $d_{g_{n}\textrm{Ax}(f_{n})}(x,x_{0})\leq C$.
By assumption, $\mathrm{Y}$ is a proper metric space. Up to passing to a subsequence, the sequence of $\alpha_{m}$ converges pointwise to a limiting geodesic ray denoted by $\gamma_{n}$, whence $g_{n}o$ is an $(\hat{r},f_{n})$-barrier for $\gamma_{n}$. Passing to a subsequence, a limiting geodesic ray $\gamma$ of $\gamma_{n}$ still possesses infinitely many distinct $(\hat{r},F)$-barriers $g_{n}o$.
In particular, $\|\gamma\cap N_{\hat{r}}(g_{n}\textrm{Ax}(f))\|\geq\|Fo\|_{\mathrm{min}}$. If $\|Fo\|_{\mathrm{min}}$ is sufficiently large, we obtain the convergence of $g_{n}o\to[\xi]$ from Lemma 3.6.
The general case for $x\in\mathrm{Y}$ could further derived from Lemma 2.9: we leave the proof to the interested reader.
By Lemma 2.16, the truncation of $\gamma$ relative to barriers is an $(L,B)$-admissible ray of infinite type, for some $L=L(\|Fo\|_{\min})$. The assertion (2) thus follows by Lemma 3.7. ∎
As an immediate corollary, we can fix the light source $x$ in (19) for convenience.
Corollary 4.8.
For any $r>0$ there exists $\hat{r}>0$ such that for any $x\in Go$, we have
$\Lambda_{r}^{F}(Go)\subseteq\limsup_{y\in Go}\Pi_{x}^{F}(y,\hat{r})$
4.2. Conical points using admissible paths
The second assertion in the “moreover” statement of Lemma 4.7 motivates the third formulation of conical points. It is based on a special class of $(L,B)$-admissible paths. Namely, an $(L,B)$-admissible ray $\gamma$ is said of infinite type if one of the following holds:
  1. (1)
    either its saturation $\mathscr{F}(\gamma)$ contains infinitely many elements, or
  2. (2)
    $\gamma$ is eventually contained in a finite neighborhood of a contracting subset $X\in\mathscr{F}(\gamma)$. In particular, $\mathscr{F}(\gamma)$ is finite.
In the first case, $\mathscr{F}(\gamma)$ is infinite, so it forms an escaping sequence of $C$-contracting subsets (as the action $G\curvearrowright\mathrm{Y}$ is proper). In the second case, if an admissible path $\gamma$ has a finite collection $\mathscr{F}(\gamma)$ of associated contracting subsets, then we can write $\gamma=p_{0}q_{1}p_{1}\cdots q_{n}p_{n}$, where $p_{n}$ is a non-trivial geodesic ray in a finite neighborhood of $X$. We refer the reader to Definition 2.12 of admissible paths.
Let the constant $B$ satisfy the truncation of a geodesic in Lemma 2.16.
Definition 4.9.
A non-pinched point $\xi\in\mathcal{C}$ is called an $(L,B,\mathscr{F})$-conical point if there exists an $(L,B)$-admissible ray of infinite type, which accumulates into $[\xi]$.
By definition, the set of $(L,B,\mathscr{F})$-conical points is $G$-invariant.
We continue to clarify the quantitative equivalence between different formulations of conical points.
Lemma 4.10.
Let $r>4C,F\subseteq G$ be as above. There exist $L_{1},L_{2}>0$ such that
  1. (1)
    the []delimited-[][$\cdot$][ ⋅ ]-locus of an $(L_{1},B,\mathscr{F})$-conical point $\xi\in\partial{\mathrm{Y}}$ contains an $(r,F)$-conical point,
  2. (2)
    the []delimited-[][$\cdot$][ ⋅ ]-locus of an $(r,F)$-conical point contains $(L_{2},B,\mathscr{F})$-conical point $\xi\in\partial{\mathrm{Y}}$.
Proof.
The statement (2) already follows from the second assertion of Lemma 4.7: if $\xi$ is an $(r,F)$-conical point, then there exists an $(L_{2},B,\mathscr{F})$-admissible ray $\gamma$, which accumulates in $[\xi]$.
Let $L_{1}=\|Fo\|+L_{0}+6r>0$, where $L_{0}$ is given by Lemma 2.10. Let us now assume that $\xi$ is an $(L_{1},B,\mathscr{F})$-conical point, and prove the statement (1).
Let $\gamma$ be an $(L_{1},B)$-admissible ray of infinite type ending in $[\xi]$. Take a sequence of geodesics $\alpha_{n}:=[o,x_{n}]$ for an unbounded sequence of points $x_{n}$ on $\gamma$. Any subpath of $\gamma$ is $(L_{1},B)$-admissible, so has the $r$-fellow travel property by Proposition 2.15:
$\|\alpha_{n}\cap N_{r}(X_{i})\|>L_{1}-2r$
for a sequence of $X_{i}\in\mathscr{F}(\gamma)$. An Cantor argument using Ascoli-Arzela lemma obtains a limiting geodesic ray $\alpha$ from $o$, satisfying $\|\alpha\cap N_{r}(X_{i})\|>L_{1}-2r$ as well.
From Lemma 3.6, we know that $\alpha$ accumulates into $[\xi]$. By Lemma 2.4, we have
$d_{X_{i}}(\alpha^{-},\alpha^{+})\geq\|\alpha\cap N_{r}(X_{i})\|-4r\geq L_{1}-6r\geq\|Fo\|+L_{0}.$
By Lemma 2.10, $\alpha$ contains infinitely many $(r,F)$-barriers, so by definition 4.3, any accumulation point of $\alpha$ in the boundary is an $(r,F)$-conical point. ∎
Furthermore, conical points are visual from each other.
Lemma 4.11.
For any $[\xi]\neq[\eta]$ in $\Lambda_{r}^{F}(Go)$, there exists a bi-infinite geodesic so that the two half rays, both with infinitely many $(\hat{r},F)$-barriers, accumulate into $[\xi],[\eta]$ respectively.
Proof.
By Lemma 4.10, let $\alpha,\beta$ be two $(L,B)$-admissible rays of infinite type ending at $[\xi],[\eta]$ respectively. Let $\mathscr{F}(\alpha)=\{X_{n}\}$ and $\mathscr{F}(\beta)=\{Y_{n}\}$ be the corresponding saturation of $\alpha$ and $\beta$.
According to the definition of infinite type, if $\mathscr{F}(\alpha),\mathscr{F}(\beta)$ are both finite, then $\alpha,\beta$ are eventually contained in a finite neighborhood of contracting quasi-geodesics, the last one in $\mathscr{F}(\alpha)$ and $\mathscr{F}(\beta)$ respectively. In this case, the proof is completed by Lemma 3.18.
Let us thus assume that one of $\mathscr{F}(\alpha)$ and $\mathscr{F}(\beta)$, say $\mathscr{F}(\alpha)$, is infinite. By Lemma 3.17, we have $\xi\notin[\Lambda X_{m}]$ for each fixed $X_{m}$. Now, let $x_{n}\in\alpha\cap N_{r}(X_{n})$ and $y_{n}\in\beta\cap N_{r}(Y_{n})$ be two unbounded sequences. Consider the sequence of geodesics $\gamma_{n}:=[x_{n},y_{n}]$. We claim that $d(o,\gamma_{n})$ are uniformly bounded for all $n\gg 0$. See Fig. 7.
[…]
Figure 7. Conical points are visual.
Indeed, suppose to the contrary that the sequence of $[x_{n},y_{n}]$ is escaping. For $\xi\notin[\Lambda X_{m}]$, Corollary 3.23 shows
$\forall n\gg m:\quad d_{X_{m}}(x_{n},y_{n})\leq C$
By Proposition 2.15, $\alpha$ has $r$-fellow travel property for $X_{m}$ and then by Lemma 2.4, we have
$d_{X_{m}}(o,x_{n})\geq\|[o,x_{n}]\cap N_{r}(X_{m})\|-4r\geq L-4r$
which yields by the triangle inequality $d_{X_{m}}$:
$d_{X_{m}}(o,y_{n})>L-C-4r.$
By Lemma 2.4, if $L>C-7r$, we have $\beta$ intersects $N_{r}(X_{m})$. As $X_{n}$ accumulates into $[\xi]$ by Lemma 3.6, we obtain that $\beta$ does so as well. This leads to a contradiction with $\alpha\neq\beta$. The claim follows.
As $\mathrm{Y}$ is a locally compact complete metric space, a subsequence of $\gamma_{n}$ converges locally uniformly to a bi-infinite geodesic $\gamma$, which intersects $N_{r}(X_{m})$ and $N_{r}(Y_{m})$ for all $m\gg 0$. Moreover, $X_{m}$ and $Y_{m}$ produce an infinite sequence of $(\hat{r},F)$-barriers on the two half rays $\gamma^{-},\gamma^{+}$ of $\gamma$, where $\hat{r}$ is given by Lemma 2.9. According to Lemma 3.6, $\gamma^{-},\gamma^{+}$ accumulate into $[\xi]\neq[\eta]$ respectively.
In summary, we proved a bi-infinite geodesic $\gamma$ with desired properties, connecting $[\xi]$ and $[\eta]$. The proof of lemma is complete. ∎
We next show that our notion of conical points satisfies the dynamical property of conical points in the theory of convergence groups.
Lemma 4.12.
Let $\xi\in\Lambda_{r}^{F}(Go)$ be a conical point. Then there exists a sequence of elements $g_{n}\in G$ and a pair of $[a]\neq[b]\in[\partial{\mathrm{Y}}]$ such that the following holds
$g_{n}(\xi,x)\to([a],[b]),$
for any point $x\in\mathrm{Y}$.
Remark 4.13.
This conclusion could hold for certain boundary points $x\in\partial{\mathrm{Y}}$, for instance if $\zeta$ is a conical point. We restrict to the points in $\mathrm{Y}$ to simplify the proof.
Proof.
By the proof of Lemma 4.7, there exists a geodesic $\gamma$ from $x$ to $\xi$, so that there exists a sequence $X_{n}\in\mathscr{F}$ such that $\|\gamma\cap N_{r}(X_{n})\|\geq\|Fo\|_{\min}$. Passing to subsequence, we assume that $X_{n}=g_{n}^{-1}\textrm{Ax}(f)$ for the same $f\in F$ and $g_{n}\in G$.
If $X:=X_{n}$ are eventually constant, then $\gamma$ is eventually contained in a finite neighborhood of $X$, so $[\xi]$ is the end point of a positive ray of $X$ (with appropriate arc-length parametrization). Thus, the diameter of $g_{n}\gamma\cap N_{r}(\textrm{Ax}(f))$ tends to $\infty$, that is: $\pi_{\textrm{Ax}(f)}(g_{n}x,g_{n}\xi)\to\infty$. According to our convention, $F$ consists of non-pinched contracting elements. Let $[a]\neq[b]$ be the attracting and repelling fixed points of non-pinched $f$. The conclusion in this case now follows from Lemma 3.28.
Passing to a subsequence, we next assume that $X_{n}$ are distinct. Let $\alpha_{n}$ be the geodesic segment of $g_{n}\gamma$ before entering $N_{r}(\textrm{Ax}(f))$, and $\beta_{n}$ the remaining half ray. By Lemma 2.9, two subsequences of $\alpha_{n}$ and $\beta_{n}$ converge respectively to two geodesic rays $\alpha,\beta$ with infinitely many $(\hat{r},F)$-barriers. By Lemma 3.6 using crucially Assump B, $\alpha,\beta$ accumulate to $[a],[b]$ respectively, which depend only on the sequence $\{X_{n}\}$. The proof is complete. ∎
4.3. Myrberg points
We are in a position to formulate a more intrinsic subclass of conical points, called Myrberg points, in a sense that Myrberg points do not refer to a contracting system as conical points do. It has several equivalent characterizations (cf. [Tuk94b]); we give the one in line with the convergence property of conical points in Lemma 4.12.
Recall that a point $\xi\in\partial{\mathrm{Y}}$ is called visual if there exists a geodesic ray starting from any point in $\mathrm{Y}$ and accumulating into $[\xi]$. In Lemma 4.11, we proved that conical points are visual.
Definition 4.14.
A visual point $\xi\in\mathcal{C}$ is called Myrberg point if for any pair of $[a]\neq[b]\subseteq[\Lambda(Go)]$ there exists a sequence of elements $g_{n}\in G$ such that the following
$g_{n}(\xi,x)\longrightarrow([a],[b])$
holds for any $x\in\mathrm{Y}$. Denote by $\Lambda_{m}{(Go)}$ the set of Myrberg points.
First of all, let us note that the fixed points of a contracting element are conical, but never Myrberg points.
Lemma 4.15.
Assume that a non-elementary $G$ contains a non-pinched contracting element. Then the locus of fixed points of any contracting element contain no Myrberg point.
Proof.
By Lemma 3.20, a non-elementary $G$ must contain at least two independent non-pinched contracting elements $k_{1},k_{2}$. Any contracting element $h$ must be independent with one, say $k$, of $\{k_{1},k_{2}\}$. Hence, by independence, any $G$-translated axis of $h$ has uniformly bounded projection to the axis of $k$. If $[h^{+}]$ contains a Myrberg point $\xi$, we have $g_{n}x\to[k^{-}]$ and $g_{n}\xi\to[k^{+}]$. By Corollary 3.28, it follows that $d_{\textrm{Ax}(k)}(g_{n}x,g_{n}\xi)\to\infty$. This is a contradiction. Thus, $[h^{+}]$ contains no Myrberg points. The proof is complete. ∎
$o$$g_{n}A\mathrm{x}(h)$$g_{n}o$$g_{n}h^{k}o$$\xi$$\leq r$$\leq r$
Figure 8. Myrberg Points: set $f=h^{k}\in E(h)$ in Lemma 4.16
The following result explains the canonical feature of a Myrberg point. Let $A$ be a set of non-pinched contracting elements so that its fixed points pair $([h^{+}],[h^{-}])$ of elements $h\in A$ is dense in the distinct pairs $\Lambda G\Join\Lambda G$ of limit points defined (2).
Lemma 4.16.
A point $\xi\in\mathcal{C}$ is a Myrberg point if and only if the following holds.
Let $f\in E(h)$ be a non-pinched contracting element for some $h\in A$. Any geodesic ray $\gamma$ ending at $[\xi]$ contains infinitely many distinct $(r,f)$-barriers. Here $r$ depends only on the contracting constant of $\textrm{Ax}(h)$ (but not on the choice of $f$).
Proof.
$\Longrightarrow$”: Let $\gamma$ be a geodesic ray starting at $x\in\mathrm{Y}$ and ending at a Myrberg point $[\xi]$. Fix a non-pinched contracting element $h$. Applying the definition of Myrberg points to $[a]:=[h^{-}],[b]:=[h^{+}]$, we have $g_{n}(\xi,x)\to([a],[b])$ for a sequence of $g_{n}\in G$. By Lemma 3.22 we have a well-defined projection map $\pi_{X}:\mathrm{Y}\cup\mathcal{C}\setminus[\Lambda X]\to X$ where $X=\textrm{Ax}(h)$.
Set $X_{n}:=g_{n}^{-1}X$. As $g_{n}(\xi,x)\to([a],[b])$, by Corollary 3.28, there exist $n_{0}>0$ such that for any $n\geq n_{0}$,
$d_{X_{n}}(x,\xi)\gg d(o,fo)$
By Lemma 2.10, $\gamma$ contains an $(r,f)$-barrier $g_{n}^{-1}X$. So the direction $\Longrightarrow$” follows.
[…]
$\|g_{n}\gamma\cap N_{r}(X)\|\geq d(o,f_{n}o)$
which tends to $\infty$ and thus yields $g_{n}(\xi,x)\to([h^{-}],[h^{+}])$ by Corollary 3.28. As $h$ is arbitrary, Lemma 3.15 shows that $\xi$ is a Myrberg point. The proof is complete. ∎
The following corollary summarizes the relation between conical points and Myrberg points.
Corollary 4.17.
We have
$\Lambda_{m}{(Go)}=\bigcap\Lambda_{r}^{F}(Go)$
where the intersection is taken over the sets $F\subseteq G$ of three mutually independent non-pinched contracting elements in $G$.
4.4. Gromov boundary of projection complex
Recall that $\mathscr{F}$ defined in (3) is the collection of $G$-translated axis of three independent non-pinched elements $F$ in $G$. We recall some facts in §2.6 on the projection complex and quasi-tree of spaces built from $\mathscr{F}$, which are used in this and next subsections.
Let $\mathcal{P}_{K}(\mathscr{F})$ be the projection complex built from $\mathscr{F}$, where $K\gg 0$ satisfies Lemma 2.21. We denote by $\partial\mathcal{P}_{K}(\mathscr{F})$ the Gromov boundary.
Set $\widetilde{X}:=\bigcup\{X\in\mathscr{F}\}$, which is the union of all quasi-lines in $\mathscr{F}$ as a subset in $\mathrm{Y}$. Let us recall the shadow map $\widetilde{X}\to\mathcal{P}_{K}(\mathscr{F})$ in (14), which is a coarsely Lipschitz map, collapses all points in $X\in\mathscr{F}$ to a vertex $\bar{x}$:
[…]
By Lemma 2.25, a standard path in $\mathcal{P}_{K}(\mathscr{F})$ is lifted via (14) to an admissible path in $\mathrm{Y}$, which is sent under this map to an unparametrized quasi-geodesic.
The goal of this subsection is examining how the map (14) induces a boundary map from the Gromov boundary $\partial\mathcal{P}_{K}(\mathscr{F})$ to the conical limit set.
Let us start with the definition of Gromov boundary.
Fix a basepoint $\bar{o}\in\mathcal{P}_{K}(\mathscr{F})$. We say that a (unbounded) sequence $\{\bar{u}_{n}\}$ is admissible if $\langle\bar{u}_{n},\bar{u}_{m}\rangle_{\bar{o}}\to\infty$ as $n,m\to\infty$, and two sequences $\{\bar{u}_{n}\},\{\bar{v}_{n}\}$ are equivalent if $\langle\bar{u}_{n},\bar{v}_{n}\rangle_{\bar{o}}\to\infty$. By definition, the Gromov boundary $\partial\mathcal{P}_{K}(\mathscr{F})$ consists of equivalent classes of admissible sequences $\{\bar{u}_{n}\}$.
Lemma 4.18.
For any $\xi\in\partial\mathcal{P}_{K}(\mathscr{F})$, there exists a quasi-geodesic ray $\gamma$ in $\mathcal{P}_{K}(\mathscr{F})$ starting at $\bar{o}$ and ending at $\xi$, so that any subpath of $\gamma$ is a standard path (cf. §2.6).
The path $\gamma$ in the statement shall be refereed to as a standard ray.
Proof.
Let $\bar{u}_{n}\to\xi\in\partial\mathcal{P}_{K}(\mathscr{F})$, and $\alpha_{n}$ be the standard path from $\bar{o}$ to $\bar{u}_{n}$, which is a uniform quasi-geodesic. The triangle of standard paths has the tripod-like property in Lemma 2.24. As $\langle\bar{u}_{n},\bar{u}_{m}\rangle_{\bar{o}}\to\infty$, one can extract a subsequence of $\alpha_{n}$ such that it converges to a quasi-geodesic ray $\alpha$ in a sense that restricting on any bounded subsets in $\mathcal{P}_{K}(\mathscr{F})$, $\alpha_{n}$ coincides with $\alpha$. Thus, any subpath of $\alpha$ are standard paths, and $\alpha$ converges to the boundary point $\xi$. The tripod-like property actually implies that $\alpha$ is unique, provided that $\bar{o}$ is fixed. ∎
The first main result is that the Gromov boundary $\partial\mathcal{P}_{K}(\mathscr{F})$ embeds into the set of $(r,F)$-conical points. We emphasize this result crucially uses Assump B instead of Assump B’.
Lemma 4.19.
For any $r,F$ satisfying Convention 4.1, there exists $K>0$ such that the following holds.
There is a continuous embedding of the Gromov boundary $\partial\mathcal{P}_{K}(\mathscr{F})$ of $\mathcal{P}_{K}(\mathscr{F})$ into $[\Lambda_{r}^{F}(Go)]$ of $(r,F)$-conical points in the quotient space $[\partial{\mathrm{Y}}]$.
Proof.
Given $\xi\in\partial\mathcal{P}_{K}(\mathscr{F})$, let $\alpha$ be a standard ray ending at $\xi$ given by Lemma 4.18. By Lemma 2.25, there exist $L=L(K),B=B(\mathscr{F})$ so that we can lift $\alpha$ to get an $(L,B)$-admissible quasi-geodesic ray $\tilde{\alpha}$. According to definition 4.9, $\tilde{\alpha}$ accumulates into the $[\tilde{\alpha}^{+}]$-class of an $(L,B,\mathscr{F})$-conical point in $\partial{\mathrm{Y}}$. Moreover, $L\to\infty$ as $K\to\infty$.
[…]
Figure 9. Injectivity of $\Phi$: standard rays $\alpha,\beta$ in $\mathcal{P}_{K}(\mathscr{F})$ (left); admissible path $\gamma$ in $\mathrm{Y}$ (right)
Defining the map $\Phi$. For given $F$, we can choose $K$ and thus $L$ large enough, so that an $(L,B,\mathscr{F})$-conical point is an $(r,F)$-conical point by Lemma 4.10. Therefore, $\tilde{\alpha}$ determines a unique $(r,F)$-conical point $[\gamma^{+}]\in[\partial{\mathrm{Y}}]$. Setting $\Phi(\xi):=[\tilde{\alpha}^{+}]$ defines the desired map
$\Phi:\partial\mathcal{P}_{K}(\mathscr{F})\to[\Lambda_{r}^{F}(Go)]$
The map $\Phi$ is injective. Indeed, if $\Phi(\xi)=\Phi(\eta)$, assume to the contrary that the corresponding standard rays $\alpha,\beta$ ending at $\xi\neq\eta$ coincide until a vertex $\bar{x}_{0}$.
By the tripod triangle property in Lemma 2.24, any $\bar{u}\in\alpha$ and $\bar{w},\bar{v}\in\beta$ far from the departure $\bar{x}_{0}$ is contained on a standard path, and if they appear in this order, then Lemma 2.21 implies $d_{W}(U,V)>K$, where $U,V,W$ are associated vertex spaces. See Fig. 9 for a schematic illustration.
Let $\tilde{\alpha}$ (resp. $\tilde{\beta}$) be the lifted path of $\alpha$ (resp. $\beta$) described in Lemma 2.25. From the construction, each vertex space $U$ associated to a vertex $\bar{u}$ on $\alpha$ (resp. $\beta$) is contained in the saturation of the $(L,B)$-admissible path $\tilde{\alpha}$ (resp. $\tilde{\beta}$).
As $\tilde{\alpha},\tilde{\beta}$ terminate at the same conical point $\Phi(\xi)=\Phi(\eta)$, there exists a geodesic ray $\gamma$ by Lemma 4.7 which fellow travels both $\tilde{\alpha}$ and $\tilde{\beta}$ by Proposition 2.15. In particular, if $W,U,V$ appear in this order on $\gamma$, then $d_{W}(U,V)\leq B$. As $\bar{u}\in\alpha,\bar{v}\in\beta$ can be chosen arbitrarily far from $\bar{w}$, such an order exists. Thus, for $K\gg B$, we obtain a contradiction with the above inequality $d_{W}(U,V)>K$, whence the injectivity of $\Phi$ follows.
The continuity of $\Phi$. Let $\xi_{n}\to\xi$ in the Gromov boundary. Our goal is to prove that $\Phi(\xi_{n})$ converges to $[\Phi(\xi)]$.
By Lemma 4.18, there exist a standard ray $\alpha_{n}$ ending at $\xi_{n}$, and its lift $(L,B)$-admissible path $\tilde{\alpha}_{n}$ in $\mathrm{Y}$ ending at $\Phi(\xi_{n})$. Moreover, $\alpha_{n}$ converges in the above sense to a standard ray $\alpha$ ending at $\xi$, and the lift $\tilde{\alpha}$ is an $(L,B)$-admissible path, ending at $[\Phi(\xi)]$.
If $\bar{x}_{n}$ is the last vertex of the overlap $\alpha_{n}\cap\alpha$, the corresponding vertex space $X_{n}\in\mathscr{F}$ appears in the saturation of the $(L,B)$-admissible path $\tilde{\alpha}$. As $\bar{x}_{n}\to\xi$, we have $d(\bar{o},\bar{x}_{n})\to\infty$. The map $X_{n}\to\bar{x}_{n}$ is Lipschitz, so we obtain $d(o,X_{n})\to\infty$. By Lemma 3.6, the escaping sequence of $X_{n}$ accumulates into $[\Phi(\xi)]$.
Let $v_{n}\in\mathrm{Y}$ be the exit point of $\tilde{\alpha}_{n}$ at $N_{r}(X_{n})$, where $r\geq 0$ is given by Proposition 2.15. Look at the sub-ray $[v_{n},\Phi(\xi_{n})]_{\tilde{\alpha}_{n}}$ of $\tilde{\alpha}_{n}$ starting from $v_{n}$. For any $z_{n}\in[v_{n},\Phi(\xi_{n})]_{\tilde{\alpha}_{n}}$ sufficiently far from $v_{n}$, we have $\pi_{X_{n}}(z_{n})$ is $(r+C)$-close to $v_{n}$ by Lemma 2.3. So the triangle inequality for $d_{X_{n}}$ gives
$d_{X_{n}}(o,z_{n})\geq d_{X_{n}}(o,v_{n})-r-C$
As $d_{X_{n}}(o,v_{n})\geq\|\gamma_{n}\cap N_{r}(X_{n})\|-4r\geq L-4r$ by Lemma 2.4, we have
$d_{X_{n}}(o,z_{n})\geq L-5r-C>10C$
implying $z_{n}\in\Omega_{o}(N_{C}(X_{n}),10C)$, where $L>5r+11C$ is assumed. As $X_{n}$ converges to $[\Phi(\xi)]$, Assump B implies that $z_{n}$ converges to $\Phi(\xi)$.
Fix a metric $\delta$ on the metrizable boundary $\partial{\mathrm{Y}}$, and choose $z_{n}$ so that $\delta(z_{n},\Phi(\xi_{n}))\leq 1/n$. Thus, $\Phi(\xi_{n})$ converges to $[\Phi(\xi)]$.
[…]
Figure 10. Continuity of $\Phi$: Standard rays in $\mathcal{P}_{K}(\mathscr{F})$ (left); Admissible paths in $\mathrm{Y}$ (right)
The continuity of $\Phi^{-1}$. Conversely, let $[p_{n}]\to[p]$ in the convergence boundary $[\partial{\mathrm{Y}}]$, where $[p_{n}],[p]\in[\Lambda_{r}^{F}(Go)]$. Let $\alpha_{n},\alpha$ be the corresponding standard rays ending at $\xi_{n}:=\Phi^{-1}([p_{n}])$ and $\xi:=\Phi^{-1}([p])$ by Lemma 4.18. We assume that $\alpha_{n},\alpha$ depart at the vertex $\bar{x}_{n}$. To prove $\xi_{n}\to\xi$, it suffices to show that $d(\bar{o},\bar{x}_{n})$ is unbounded. Let us assume to the contrary that $d(\bar{o},\bar{x}_{n})$ is bounded for infinitely many $\alpha_{n}$.
Let us fix any $\bar{u}_{n}\in\alpha_{n}$ and $\bar{w},\bar{v}_{n}\in\alpha$ far from $\bar{x}_{n}$. By Lemma 2.24, $\bar{u}_{n},\bar{w},\bar{v}_{n}$ are necessarily contained in a standard path (say, in this order), so $d_{W}(U_{n},V_{n})>K$. Recall that the saturation of $\tilde{\alpha}_{n}$ (resp. $\tilde{\alpha}$) contains the vertex spaces associated to the vertices on $\alpha_{n}$ (resp. $\alpha$).
By definition of $\Phi$, the admissible paths $\tilde{\alpha}_{n}$ and $\tilde{\alpha}$ terminate at $[p_{n}],[p]$ respectively, so by Lemma 4.7, $U_{n}$ (resp. $V_{n}$) accumulates into $[p_{n}]$ (resp. $[p]$). However, as $d_{W}(U_{n},V_{n})>K$, the contracting property implies that any geodesic from $u_{n}\in U_{n}$ to $v_{n}\in V_{n}$ intersects a fixed ball. Since $u_{n}\to[p],v_{n}\to[p]$, we obtain a contradiction with $p\in\mathcal{C}$. Hence, the continuity of $\Phi^{-1}$ follows, and the proof of the lemma is complete. ∎
4.5. Gromov boundary of quasi-tree of spaces
Starting from a family $\mathscr{F}$ of quasi-lines as in (3), the construction of the projection complex $\mathcal{P}_{K}(\mathscr{F})$ records no information from those lines in $\mathscr{F}$. The quasi-tree of spaces $\mathcal{C}(\mathscr{F})$ is a blown-up version of $\mathcal{P}_{K}(\mathscr{F})$, so that it contains isometrically embedded copies of quasi-lines in $\mathscr{F}$. It is exercise to verify that the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$ as a set is surjective to the union of $\partial\mathcal{P}_{K}(\mathscr{F})$ and the vertex set $\mathscr{F}$:
$\partial\mathcal{C}(\mathscr{F})\longrightarrow(\partial\mathcal{P}_{K}(\mathscr{F}))\cup\mathscr{F}$
though this map is not necessarily continuous with respect to the obvious topology on the target. Note that the non-injectivity only comes from collapsing the two endpoints of each quasi-line $X\in\mathscr{F}$ to be one corresponding vertex $X$ in $\mathcal{P}_{K}(\mathscr{F})$.
Corollary 4.20.
Under the assumption of Lemma 4.19, there is a continuous embedding $\Phi$ from the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$ into $[\Lambda_{r}^{F}(Go)]$.
Proof.
By the construction of $\mathcal{C}(\mathscr{F})$, there are two types of sequences $u_{n}\in\mathcal{C}(\mathscr{F})$ tending to Gromov boundary. Either the corresponding vertices $\bar{u}_{n}$ are unbounded in the original projection complex $\mathcal{P}_{K}(\mathscr{F})$ or $u_{n}$ are eventually contained in one vertex space, i.e.: an axis $U$ in $\mathscr{F}$. The former case has been dealt in Lemma 4.19. In the latter case, $U$ is a axis of non-pinched contracting element in $\mathrm{Y}$ with distinct endpoints $[U^{-}]\neq[U^{+}]$. Let $\Phi(\xi)$ be the corresponding $[U^{-}]$ or $[U^{+}]$ to which $u_{n}$ converges.
The continuity follows a similar argument as in Lemma 4.19, and is left to the interested reader. ∎
The map $\partial\mathcal{C}(\mathscr{F})\to[\Lambda_{r}^{F}(Go)]$ in Corollary 4.20 admits the following inverse in a qualitative way. Compare the order of quantifiers in the two statements.
Lemma 4.21.
For any $K\gg 0$ in Lemma 2.21, if $\|Fo\|_{\min}\gg K$, then the set $[\Lambda_{r}^{F}(Go)]$ of $(r,F)$-conical points admits a continuous embedding into the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$.
Proof.
Let $[\xi]$ be the []delimited-[][$\cdot$][ ⋅ ]-class of an $(r,F)$-conical point. By Lemma 4.10, there exists an $(L,B)$-admissible ray $\gamma$ ending at $[\xi]$, where $L$ is comparable with $\|Fo\|_{\min}$ and $B$ depends on $\mathscr{F}$. Let $\mathscr{F}(\gamma)$ be the collection of contracting subsets, which is a subset of $\mathscr{F}$, associated to the admissible path $\gamma$ (cf. Definition 2.12).
Defining the desired map $\Psi:[\Lambda_{r}^{F}(Go)]\to\partial\mathcal{C}(\mathscr{F})$. The discussion is divided into the following two cases.
If $\gamma$ is eventually contained in the $r$-neighborhood of a quasi-line $U\in\mathscr{F}$, then the corresponding half-ray of $U$ defines a Gromov boundary point of $\mathcal{C}(\mathscr{F})$. Here, $U$ is isometrically embedded convex into $\mathcal{C}(\mathscr{F})$.
We are thus left consider the case where the saturation $\mathscr{F}(\gamma)$ is infinite. That is, $\gamma$ passes through the $r$-neighborhood of infinitely many distinct axes in $\mathscr{F}(\gamma)$.
If $\|Fo\|_{\min}\gg K$, then $L$ is large enough so that for any $U,W,V\in\mathscr{F}(\gamma)$ in this order, we have $d_{W}(U,V)\geq K$ by Proposition 2.15. In different words, any subset of consecutive elements in $\mathscr{F}(\gamma)$ forms a standard path in $\mathcal{P}_{K}(\mathscr{F})$.
Applying a limiting argument as in Lemma 4.18 to the sequence of standard paths between two vertices in $\mathscr{F}(\gamma)$, we obtain a standard ray $\alpha$ in $\mathcal{P}_{K}(\mathscr{F})$ which contains $\mathscr{F}(\gamma)$ as a (possibly proper) subset of vertices.
Now, the standard ray $\alpha$ in $\mathcal{P}_{K}(\mathscr{F})$ gives naturally a standard ray $\beta$ in $\mathcal{C}(\mathscr{F})$ by [BBF15, Def. 4.3]: $\beta$ is obtained from $\alpha$ by replacing vertices $\bar{u}$ of $\alpha$ in $\mathcal{P}_{K}(\mathscr{F})$ with appropriate geodesic segments in the associated axes $U$ (cf. Theorem 2.27). As $\alpha$ is a quasi-geodesic by [BBFS19, Cor. 3.7], so $\beta$ is a quasi-geodesic in $\mathcal{C}(\mathscr{F})$. Thus, we see that $\beta$ determines a boundary point $p$ in $\partial\mathcal{C}(\mathscr{F})$. We thus obtain the map $\Psi:[\Lambda_{r}^{F}(Go)]\to\partial\mathcal{C}(\mathscr{F})$ by defining $\Psi([\xi])=p$.
The map $\Psi$ is injective. Indeed, assume to the contrary that $\Psi([\xi])=\Psi([\eta])=p$ for two distinct $[\xi]\neq[\eta]\in[\Lambda_{r}^{F}(Go)]$.
Let $\gamma,\gamma^{\prime}$ be two $(L,B)$-admissible rays ending at $[\xi]$ and $[\eta]$ respectively. If $\gamma$ or $\gamma^{\prime}$ are eventually contained in the finite neighborhood of $X\in\mathscr{F}(\gamma)$ or of $X^{\prime}\in\mathscr{F}(\gamma^{\prime})$, then it is easy to arrive a contradiction, as $X\neq X^{\prime}$ have disjoint limit sets in $\mathcal{C}(\mathscr{F})$. Hence, let us focus on the generic case that $\mathscr{F}(\gamma)$ and $\mathscr{F}(\gamma^{\prime})$ are both infinite.
Let $\alpha$ be the standard ray ending at $p$. According to the above construction of $\Psi$, as $\Psi([\xi])=\Psi([\eta])=p$, we have that $\mathscr{F}(\gamma)$ and $\mathscr{F}(\gamma^{\prime})$ are contained in the vertex set of $\alpha$. On the one hand, the coarsely Lipschitz map $\widetilde{X}\to\mathcal{P}_{K}(\mathscr{F})$ in (20) sends $\mathscr{F}(\gamma)\cup\mathscr{F}(\gamma^{\prime})$ into the set of vertices on $\alpha$. On the other hand, by Lemma 4.11, there exists a bi-infinite geodesic between $[\xi]$ and $[\eta]$, which intersects each quasi-line in $\mathscr{F}(\gamma)\cup\mathscr{F}(\gamma^{\prime})$. Hence, we would obtain a contradiction, and the injectivity of $\Phi$ follows.
The continuity of $\Psi$ (resp. $\Psi^{-1}$) follows by repeating the same arguments as in Lemma 4.19 for the continuity of $\Phi^{-1}$ (resp. $\Phi$). Hence, $\Psi$ is a topological embedding. The proof is complete. ∎
As a corollary, we have
Corollary 4.22.
The quotient $[\Lambda_{m}{(Go)}]$ for $\Lambda_{m}{(Go)}\subseteq\partial{\mathrm{Y}}$ is a Hausdorff, second countable, metrizable topological space.
If $\partial{\mathrm{Y}}$ is the horofunction boundary we prove in Lemma 5.7, via different means, that $[\Lambda_{m}{(Go)}]$ is a Hausdorff and second countable topological space.
Proof.
Via the embedding in Lemma 4.21, $[\Lambda_{m}{(Go)}]$ admits an topological embedding into the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$. Moreover, as fixed points of a contracting element cannot be Myrberg point (Lemma 4.15), the image of $[\Lambda_{m}{(Go)}]$ is disjoint with the endpoints of any quasi-line in $\mathscr{F}$.
Recall that a subset of a separable metric space is separable, and a metric space is separable if and only if it is second countable. It suffices to prove that the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$ is a complete separable space with respect to the visual metric. This actually holds for hyperbolic graphs with countably many vertices. Indeed, as $\mathcal{C}(\mathscr{F})$ has countably many vertices, the shadows $\Pi_{o}(v,r)$ at vertices $v\in\mathcal{C}(\mathscr{F})$ form a countable base for the Gromov boundary. To see the completeness, we make use of the standard rays in $\mathcal{C}(\mathscr{F})$ for simplicity; in general, one could appeal to the construction of quasi-geodesic rays ([BH99]).
Every boundary point $\xi$ is represented by a standard ray $\alpha$ by Lemma 4.18. If $\xi_{n}$ is a Cauchy sequence, then the almost tripod property of standard paths implies that the associated ray $\alpha_{n}$ converges to a standard ray. This shows the completeness of $\partial\mathcal{C}(\mathscr{F})$. ∎
5. Horofunction boundary as convergence boundary
In this section, we verify that the horofunction boundary with finite difference relation is a convergence boundary, i.e. satisfying Assump A,Assump B and Assump C.
Let us first recall the definition of limit sets with respect to the horofunction boundary. The limit set $\Lambda(Go)$ of a group $G$ acting isometrically on $\mathrm{Y}$ is defined to be the topological closure of a $G$-orbit $Go$ in $\partial_{\mathrm{H}}{\mathrm{Y}}$. We have $[\Lambda(Go)]=[\Lambda(Go^{\prime})]$ by Lemma 2.29 (note that $G$ is not necessarily discrete). We sometimes denote $[\Lambda(G)]=[\Lambda(Go)]$ for simplicity.
5.1. Verifying Assumptions
This subsection is devoted to the proof of Theorem 1.1. We first verify Assump A for the horofunction boundary.
Lemma 5.1.
Let $X\subseteq\mathrm{Y}$ be a $C$-contracting subset for $C\geq 0$. If $y_{n}\in\mathrm{Y}\to\eta\in\partial_{\mathrm{H}}{\mathrm{Y}}$ and $x_{n}\to\xi\in\partial_{\mathrm{H}}{\mathrm{Y}}$ for a sequence $x_{n}\in\pi_{X}(y_{n})$, then
$\|b_{\xi}-b_{\eta}\|_{\infty}\leq 4C$
Proof.
The proof boils down to show that the difference
$\lim\limits_{n\to\infty}[b_{x_{n}}(z)-b_{y_{n}}(z)]=\lim\limits_{n\to\infty}[d(z,{x_{n}})-d(o,x_{n})+d(o,y_{n})-d(z,y_{n})]$
is uniformly bounded over any $z\in\mathrm{Y}$. As $x_{n}\in\pi_{X}(y_{n})$ exits every compact set, the $C$-contracting property of $X$ thus implies for all $n\gg 0$:
$d(x_{n},[y_{n},o])\leq C,\quad d(x_{n},[y_{n},z])\leq C.$
Combining these estimates, we get for $n\gg 0$,
$\begin{array}[]{lll}|d(z,{x_{n}})-d(o,x_{n})+d(o,y_{n})-d(z,y_{n})|\leq 4C\end{array}$
completing the proof. ∎
The following corollary is implicit in the proof.
Corollary 5.2.
Let $\gamma$ be a $C$-contracting geodesic ray for some $C\geq 0$. For any $\xi,\eta\in[\gamma^{+}]$, we have
$\|b_{\xi}-b_{\eta}\|_{\infty}\leq 4C$
Proof.
Let $y_{n}\in\mathrm{Y}\to\zeta\in[\gamma^{+}]$. By Lemma 5.1, it suffices to prove that $\pi_{\gamma}(y_{n})$ is an escaping sequence. Indeed, if not, $\{\pi_{\gamma}(y_{n}):n\geq 1\}$ is contained in an $R$-ball at $\gamma^{-}$ for $R>0$. By $C$-contracting property, as $z\in\gamma$ tends to infinity, the geodesic $[y_{n},z]$ intersects $B(\gamma^{-},R+C)$. Estimating the value of $b_{y_{n}}(\cdot)$ at $z\in\gamma$ shows $b_{y_{n}}(z)\to+\infty$. However, the horofunction $b_{\gamma}$ associated to $\gamma$ takes negative values as $z\in\gamma$ tends to infinity. This contradiction completes the proof. ∎
We now associate the fixed points to a contracting element. Let $h$ be a contracting element on $\mathrm{Y}$, so that $\langle h\rangle o$ is a contracting quasi-geodesic by definition. The repelling and attracting fixed points $[h^{-}],[h^{+}]$ in $\partial_{\mathrm{H}}{\mathrm{Y}}$ do not depend on $o\in\mathrm{Y}$. Indeed, by Corollary 5.2, any two horofunctions in $[h^{+}]$ (or $[h^{-}]$) have a finite difference depending on the contracting constant of $\langle h\rangle o$.
We now prove Assump B and Assump B’ for the horofunction boundary.
Lemma 5.3.
Let $X_{n}\subseteq\mathrm{Y}$ be an escaping sequence of $C$-contracting subsets for some $C\geq 0$. Fix a basepoint $o\in\mathrm{Y}$. Let $x_{n}\in\pi_{X_{n}}(o)$ be a sequence of points such that $x_{n}\to\xi\in\partial_{\mathrm{H}}{\mathrm{Y}}$. Consider another sequence of points $y_{n}\in\mathrm{Y}\to\eta\in\partial_{\mathrm{H}}{\mathrm{Y}}$ so that either $y_{n}\in X_{n}$ or $\|N_{C}(X_{n})\cap[o,y_{n}]\|>6C$ for all $n\geq 1$. Then $\|b_{\xi}-b_{\eta}\|_{\infty}\leq 6C$.
Proof.
Given $z\in\mathrm{Y}$, we are going to give a uniform bound on $|b_{\xi}(z)-b_{\eta}(z)|$. To this end, choose a projection point $z_{n}\in\pi_{X_{n}}(z)$. As $d(o,X_{n})\to\infty$, the ball centered at $o$ of radius $d(o,z)$ misses $X_{n}$, so the $C$-contracting property by Lemma 2.2 implies
$d(z_{n},x_{n})\leq d_{X_{n}}(z,o)\leq C$
See Fig. 11. If $y_{n}\in X_{n}$, the contracting property gives $d(z_{n},[z,y_{n}])\leq 2C$ and then $d(x_{n},[z,y_{n}])\leq 3C$. If $\|N_{C}(X_{n})\cap[o,y_{n}]\|>6C$, Lemma 2.3 implies $d_{X_{n}}(o,y_{n})>2C$, yielding
$d_{X_{n}}(z,y_{n})\geq d_{X_{n}}(o,y_{n})-d_{X_{n}}(z,o)\geq C$
by triangle inequality of $d_{X_{n}}$.
[…]
Figure 11. Schematic configuration in verifying Assumption B
Using again the contracting property in Lemma 2.3,
$d(z_{n},[z,y_{n}])\leq 2C$
which implies also $d(x_{n},[z,y_{n}])\leq 3C$. In both cases, a straightforward computation gives that
$|b_{x_{n}}(z)-b_{y_{n}}(z)|\leq|d(z,x_{n})-d(z,y_{n})|\leq 2d(x_{n},[z,y_{n}])\leq 6C.$
Letting $x_{n}\to\xi$ and $y_{n}\to\eta$ completes the proof. ∎
If $\mathrm{Y}$ is CAT(0) or hyperbolic, the assumption “$\|N_{C}(X_{n})\cap[o,y_{n}]\|>6C$” could be relaxed as follows.
Lemma 5.4.
Under the assumption of Lemma 5.3, assume that $y_{n}\in\mathrm{Y}\to\eta\in\partial_{\mathrm{H}}{\mathrm{Y}}$ so that $X_{n}\cap[o,y_{n}]\neq\emptyset$ for each $n\geq 1$. If $\mathrm{Y}$ is CAT(0) or hyperbolic, then $\|b_{\xi}-b_{\eta}\|_{\infty}\leq 6C$.
Proof.
We resume the notation in the proof of Lemma 5.3.
As $x_{n}\in X_{n}$ are the shortest projection points of $o$, the $C$-contracting property of $X_{n}$ shows that $d(x_{n},[o,y_{n}])\leq 2C$. Consider the triangle $\Delta(o,z,y_{n})$. Note that $d(o,z)$ is fixed and $x_{n}$ is $2C$-close to the side $[z,y_{n}]$. If $\mathrm{Y}$ is CAT(0) or $\delta$-hyperbolic, it follows that $d(x_{n},[z,y_{n}])<2C+\delta$: in the former case, we use the Euclidean comparison triangle and in the latter, the $\delta$-thin triangle property.
The same proof as in Lemma 5.3 then proves the conclusion with $|b_{x_{n}}(z)-b_{y_{n}}(z)|\leq 2C+\delta$ over $z\in\mathrm{Y}$. ∎
We prove now that all boundary points in the horofunction boundary $\partial_{\mathrm{H}}{\mathrm{Y}}$ are non-pinched; that is, $\mathcal{C}=\partial_{\mathrm{H}}{\mathrm{Y}}$ in Assump C. As a corollary, every contracting elements are non-pinched in the horofunction boundary.
Lemma 5.5.
For any $\xi\in\partial_{\mathrm{H}}{\mathrm{Y}}$, there exist no sequences $x_{n},y_{n}\in\mathrm{Y}$ such that $x_{n},y_{n}\to[\xi]$ and $[x_{n},y_{n}]$ intersects a fixed ball for any $n\geq 1$.
Proof.
For concreteness, assume that $x_{n}\to\xi,y_{n}\to\eta$ such that $[\xi]=[\eta]$. The finite difference relation implies that there exists $K\geq 10$ such that $|B_{\xi}(z,w)-B_{\eta}(z,w)|\leq K$ for any $z,w\in\mathrm{Y}$. Arguing by contradiction, as $\mathrm{Y}$ is proper, assume that $[x_{n},y_{n}]$ converges locally uniformly to a bi-infinite geodesic $\alpha$. Choose $z,w\in\alpha$ with distance at least $2K$ such that $z,w\in N_{1}([x_{n},y_{n}])$. For concreteness, we may assume that $z$ is closer to $x_{n}$ than $w$. By direct computation,
$|B_{x_{n}}(z,w)+d(z,w)|\leq 2,\quad|B_{y_{n}}(z,w)-d(z,w)|\leq 2.$
For $x_{n},y_{n}\to\xi$, we have $B_{x_{n}}(z,w),B_{y_{n}}(z,w)\to B_{\xi}(z,w)$ as $n\to\infty$. With $|B_{\xi}(z,w)-B_{\eta}(z,w)|\leq K$, this contradicts the assumption of $d(z,w)>2K$. ∎
5.2. Horofunctions associated to conical points
We are not presumably working with the horofunction boundary in this subsection. Still, the goal is to define a horofunction for conical points in any nontrivial convergence boundary $\overline{\mathrm{Y}}=\mathrm{Y}\cup\partial{\mathrm{Y}}$.
Lemma 5.6.
Let $\xi\in\Lambda_{r}^{F}(Go)$. Consider two sequences $x_{n}\in\mathrm{Y}\to\zeta\in[\xi]$, $z_{n}\in\mathrm{Y}\to\eta\in[\xi]$ in $\partial{\mathrm{Y}}$, and simultaneously, $x_{n}\to b_{\zeta}$, $z_{n}\to b_{\eta}$ in the horofunction boundary $\partial_{\mathrm{H}}{\mathrm{Y}}$. Then
$\|b_{\zeta}-b_{\eta}\|_{\infty}\leq 20C.$
Moreover, the locus of an $(r,F)$-conical point $\xi$ consists of $(\hat{r},F)$-conical points.
As a consequence, we can associate to any $\xi\in\Lambda_{r}^{F}(Go)$ a Busemann quasi-cocycle as follows
$B_{\xi}(x,y)=\limsup_{y\to\xi}\big{(}d(x,z_{n})-d(y,z_{n})\big{)}$
where the convergence $z\to\xi$ takes place in $\overline{\mathrm{Y}}=\mathrm{Y}\cup\partial{\mathrm{Y}}$.
Proof.
By Lemma 4.7, let $\gamma$ be a geodesic ray with infinitely many distinct $(\hat{r},F)$-barriers $y_{n}:=g_{n}o\in X_{n}\in\mathscr{F}$ and ending at $[\xi]$. For given $X_{n}$ and $m\gg n$, we have
(21) $d_{X_{n}}(o,y_{m})\geq\|N_{r}(X_{n})\cap\gamma\|-4r\geq\|Fo\|_{\min}-4r>10C,$
If $X:=X_{n}$ are eventually the same, we see that $\gamma$ is (eventually) contracting (after removing a finite initial path). In this case, the conclusion follows by Corollary 5.2.
We now assume that $\{X_{n}\}$ are escaping. Take any sequence $z_{m}\to\eta\in[\gamma^{+}]$, where $z_{m}\in\mathrm{Y}$ may be not on $\gamma$. As $y_{n}\to[\xi]=[\eta]$, Lemma 3.26 implies that $d_{X_{n}}(y_{m},z_{m})\leq 6C$ for all $m\gg 0$. By triangle inequality with (21), we obtain $\pi_{X_{n}}(z_{m},o)\geq 4C$. So $[o,z_{m}]\cap N_{C}(X_{n})\neq\emptyset$ by Lemma 2.2. In other words, $z_{m}\in\Omega_{o}(N_{C}(X_{n}))$ for every $m\gg n$. The same argument applied for $x_{n}\to\zeta$ shows that $x_{m}\in\Omega_{o}(N_{C}(X_{n}))$. By Lemma 5.3, we thus see $\|b_{\zeta}-b_{\eta}\|_{\infty}\leq 20C$.
It remains to see that $\eta$ is an $(\hat{r},F)$-conical point, where $\hat{r}$ is given by Lemma 2.9. As $d_{X_{n}}(y_{m},z_{m})\leq 6C$, $[o,z_{m}]$ contains an $(\hat{r},f)$-barrier $y_{n}=g_{n}o$ by Lemma 2.9. Thus, $z_{m}\in\Omega_{o}^{F}(g_{n}o,\hat{r})$, so $\eta$ is an $(\hat{r},F)$-conical point. ∎
We close this section with the following property of Myrberg set in the horofunction boundary.
Lemma 5.7.
The Myrberg set $\Lambda_{m}{(Go)}\subseteq\partial{\mathrm{Y}}$ is saturated in any nontrivial convergence boundary $\partial{\mathrm{Y}}$, and the finite difference relation []delimited-[][$\cdot$][ ⋅ ] of the horofunction boundary restricted on $\Lambda_{m}{(Go)}\subseteq\partial_{\mathrm{H}}{\mathrm{Y}}$ is $K$-finite for some uniform constant $K>0$.
Proof.
As Myrberg points are conical, the finite difference relation on $\Lambda_{m}{(Go)}$ is $K$-finite for some uniform $K$ by Lemma 5.6. By Lemma 4.16, a point $\xi\in\Lambda_{m}{(Go)}$ is Myrberg if and only if for any non-pinched contracting element $h$ and for any $L\gg 10C$, there exists $g\in G$ such that
$d_{X}(o,\xi)\geq L$
where $X=g\textrm{Ax}(h)$ is $C$-contracting for some $C>0$. It remains to show that $\eta\in[\xi]$ is Myrberg. By Lemma 3.26, we have $d_{X}(\xi,\eta)\leq 6C$, so by the triangle inequality for $d_{X}$,
$d_{X}(o,\eta)\geq L-6C.$
As $h$ and $L$ is arbitrary, it is proved that $\eta$ is Myrberg as well. ∎
As a corollary of Lemma 2.30 and Lemma 4.22, we have
Corollary 5.8.
The quotient $[\Lambda_{m}{(Go)}]$ for $\Lambda_{m}{(Go)}\subseteq\partial_{\mathrm{H}}{\mathrm{Y}}$ is a Hausdorff, second countable, metrizable topological space.
6. Conformal density on the convergence boundary
This section develops a general theory of quasi-conformal density on a convergence compactification $\overline{\mathrm{Y}}=\mathrm{Y}\cup\partial{\mathrm{Y}}$, under an additional Assump D introduced below.
6.1. Quasiconformal density
Buseman (quasi-)cocyles are indispensable to formulate (quasi-)conformal density on $\partial{\mathrm{Y}}$. As shown in Lemma 5.6, conical points in any convergence boundary can be associated with Buseman quasi-cocycles. Let $\mathcal{C}$ be the $G$-invariant subset of non-pinched points in Assump C. Our last assumption requires that Buseman quasi-cocycles extends to the set $\mathcal{C}$.
Given $z\in\mathrm{Y}$, let $B_{z}(x,y):=d(x,z)-d(y,z)$ for $x,y\in\mathrm{Y}$.
Assump D.
There exists a family of Buseman quasi-cocycles based at points in $\mathcal{C}$
$\big{\{}B_{\xi}:\quad\mathrm{Y}\times\mathrm{Y}\to\mathbb{R}\big{\}}_{\xi\in\mathcal{C}}$
so that for any $x,y\in\mathrm{Y}$, we have
$\limsup_{z\to\xi}|B_{\xi}(x,y)-B_{z}(x,y)|\leq\epsilon,$
where $\epsilon\geq 0$ is a universal constant not depending on $x,y$ and $\xi$. If $\epsilon=0$, we say that Buseman cocyles extend continuously to $\mathcal{C}$.
Let $\mathcal{M}^{+}(\overline{\mathrm{Y}})$ be the set of finite positive Borel measures on $\overline{\mathrm{Y}}:=\partial{\mathrm{Y}}\cup\mathrm{Y}$, on which $G$ acts by push-forward:
$g_{\star}\mu(A)=\mu(g^{-1}A)$
for any Borel set $A$.
Definition 6.1.
Let $\omega\in[0,\infty[$. Under Assump D, a map $\mu:\mathrm{Y}\to\mathcal{M}^{+}(\overline{\mathrm{Y}})$
$x\longmapsto\mu_{x}$
is a $\omega$-dimensional $G$-quasi-equivariant quasi-conformal density if for any $g,h\in G$ and any $x,y\in\mathrm{Y}$, we have
(22) $\forall\xi\in\partial{\mathrm{Y}}:\quad\frac{dg_{\star}\mu_{x}}{d\mu_{x}}(\xi)\in[\frac{1}{\lambda},\lambda],$
(23) $\forall\mu_{y}\textrm{ a.e. }\xi\in\mathcal{C}:\quad\frac{1}{\lambda}e^{-\omega B_{\xi}(x,y)}\leq\frac{d\mu_{x}}{d\mu_{y}}(\xi)\leq\lambda e^{-\omega B_{\xi}(x,y)}$
for a universal constant $\lambda\geq 1$. We normalize $\mu_{o}$ to be a probability measure: its mass $\|\mu_{o}\|=\mu_{o}(\overline{\mathrm{Y}})=1$.
If $\lambda=1$ for (22), the map $\mu:\mathrm{Y}\to\mathcal{M}^{+}(\overline{\mathrm{Y}})$ is $G$-equivariant. If both $\lambda=1$, we call $\mu$ a conformal density. Note that $\mu_{gx}=g_{\star}\mu_{x}$; equivalently, $\mu_{gx}(gA)=\mu_{x}(A)$.
Remark 6.2.
We clarify by examples the advantage of the conformality (23) restricted on a smaller subset $\mathcal{C}$ of $\partial{\mathrm{Y}}$.
  1. (1)
    First of all, it seems implausible to define Buseman cocyles at pinched points, e.g. if the pinched point is the two endpoints of a bi-infinite geodesic (i.e.: a horocycle). For example, the horocycles exist in the Cayley graph of relatively hyperbolic groups, so that in [Yan22], we can only define Buseman (quasi-)cocyles at conical points in Floyd boundary.
  2. (2)
    Secondly, if the action $G\curvearrowright\mathrm{Y}$ is of divergent type, one may prove via other means that the quasi-conformal measures are fully supported on the set $\mathcal{C}$ (in view of Theorem 1.14).
    As finitely generated groups are of divergent type for the action on the Cayley graph, we then prove in [Yan22] the atomless of the Patterson–Sullivan measures on Bowditch boundary via an argument in [DOP00]. As there are only countably many non-conical points, the conformality is recovered on almost every boundary points.
    As another example, the conformal density constructed from Thurston measures in [ABEM12] (not using Patterson’s construction) is supported on uniquely ergodic points in Thurston boundary (this follows by Masur-Veech’s result [Mas82a, Vee82] on Teichmüller geodesic flow). It turned out that the horofunction could be defined on those points (this is nontrivial as Thurston boundary is not the horofunction boundary of Teichmüller metric), on which the conformality suffices in application.
For simplicity, we write $\mu_{g}=\mu_{go}$ if the basepoint $o\in\mathrm{Y}$ does not matter in context. In particular, $\mu_{1}$ denotes the measure $\mu_{o}$ where $1$ is the identity of $G$.
Patterson–Sullivan measures
Let $(\mathrm{Y},d)$ be a proper geodesic space. Choose a basepoint $o\in\mathrm{Y}$. The Poincaré series for the action of $G\curvearrowright\mathrm{Y}$
$\mathcal{P}_{G}(s,x,y)=\sum\limits_{g\in G}e^{-sd(x,gy)},\;s\geq 0$
diverges at $0\leq s<\omega_{G}$ and converges at $s>\omega_{G}$. The action of $G$ on $\mathrm{Y}$ is called divergent if $\mathcal{P}_{G}(s,x,y)$ diverges at the critical exponent $\omega_{G}$. Otherwise, $G$ is called convergent.
We start to construct a family of measures $\{\mu_{x}^{s,y}\}_{x\in\mathrm{Y}}$ supported on $Gy$ for any given $s>\omega_{G}$ and $x,y\in\mathrm{Y}$. Assume that $\mathcal{P}_{G}(s,x,y)$ is divergent at $s=\omega_{G}$. Set
(24) $\mu_{x}^{s,y}=\frac{1}{\mathcal{P}_{G}(s,o,y)}\sum\limits_{g\in G}e^{-sd(x,gy)}\cdot{\mbox{Dirac}}{(gy)},$
where $s>\omega_{G}$. Note that $\mu^{s,y}_{o}$ is a probability measure supported on $Gy$. If $\mathcal{P}_{G}(s,x,y)$ is convergent at $s=\omega_{G}$, the Poincaré series in (24) needs to be replaced by a modified series as in [Pat76].
Fix $y\in\mathrm{Y}$. Choose $s_{i}\to\omega_{G}$ such that $\mu_{x}^{s_{i},y}$ are convergent in $\mathcal{M}(\Lambda(Gy))$. The Patterson–Sullivan measures $\mu_{x}^{y}=\lim\mu_{v}^{s_{i},y}$ are the limit measures, which may depend on the sequence $s_{i}$ and $y$. Note that $\mu_{o}^{y}(\Lambda(Gy))=1$. In what follows, we usually write $\mu_{x}=\mu_{x}^{o}$ for $x\in\mathrm{Y}$.
Lemma 6.3.
Let $G$ act properly on a proper geodesic space $(\mathrm{Y},d)$ with a contracting element, and compactified by a nontrivial convergence boundary $\partial{\mathrm{Y}}$. Then under Assump D, for any fixed $y\in\mathrm{Y}$, the family of Patterson–Sullivan measures $\{\mu_{x}^{y}\}_{x\in\mathrm{Y}}$ on $\partial{\mathrm{Y}}$ is a $\omega_{G}$-dimensional $G$-equivariant quasi-conformal density supported on the limit set $\Lambda(Gy)$.
Moreover, if $\partial{\mathrm{Y}}$ is the horofunction boundary, then it is a conformal density.
Proof.
As the orbit $Gy$ is discrete in $\mathrm{Y}$, the measure $\mu_{x}^{y}$ is supported on the closure of $Gy$. The proof for quasi-conformal density follows the same argument in [Coo93]. If $\partial{\mathrm{Y}}$ is the horofunction boundary, the Buseman cocycle extends continuously to $\partial_{\mathrm{H}}{\mathrm{Y}}$ so the constant $\epsilon=0$ in Assump D, and we obtain a conformal density: $\lambda=1$. See [BM96] for more details. ∎
In the sequel, we write PS-measures as shorthand for Patterson–Sullivan measures.
6.2. Shadow Lemma
We assume the set $\mathcal{C}$ in Assump C is measurably significant: $\mu_{1}(\mathcal{C})>0$. The goal of this subsection is to prove the shadow lemma.
The partial shadows $\Pi_{o}^{F}(go,r)$ and cones $\Omega_{o}^{F}(go,r)$ given in Definition 4.2 depend on the choice of a contracting system $\mathscr{F}$ as in (3). Without index $F$, $\Pi_{o}(go,r)$ denotes the usual shadow.
Lemma 6.4 (Shadow Lemma).
Let $\{\mu_{g}\}_{g\in G}$ be a $\omega$-dimensional $G$-quasi-equivariant quasi-conformal density on $\partial{\mathrm{Y}}$ for some $\omega>0$. Assume that $\mu_{1}(\mathcal{C})>0$. Then there exists $r_{0}>0$ such that
$\begin{array}[]{rl}e^{-\omega\cdot d(o,go)}\;\prec_{\lambda}\;\mu_{1}(\Pi_{o}^{F}(go,r)\cap\mathcal{C})\;\leq\;\mu_{1}(\Pi_{o}(go,r)\cap\mathcal{C})\;\prec_{\lambda,\epsilon,r}\;e^{-\omega\cdot d(o,go)}\\ \end{array}$
for any $g\in G$ and $r\geq r_{0}$, where $\epsilon,\lambda$ are given respectively in Assump D and Definition 6.1.
Proof.
We start proving the lower bound which is the key part of the proof.
1. Lower Bound. Fix a Borel subset $U$ in $\mathcal{C}$ such that $\mu_{1}(U)>0$ (e.g.: let $U=\mathcal{C}$). For any given $\xi\in U$, there exists a sequence of points $z_{n}\in\mathrm{Y}$ such that $z_{n}\to\xi$. By Lemma 2.17, there exist $f_{n}\in F$ and $r>0$ such that $g$ is an $(r,f_{n})$-barrier for any geodesic $[o,gfz_{n}]$. Since $F$ consists of three elements, we assume that $f:=f_{n}$ are all equal, up to taking a subsequence of $z_{n}$. This implies that $gfz_{n}\in\Omega_{o}^{F}(go,r)$ tends to $gf\xi$, so by definition,
(25) $gf\xi\in\Pi_{o}^{F}(go,r)\cap\mathcal{C}.$
We refer to Fig. 12 for an illustration.
Consequently, we decompose the set $U$ into a disjoint union of three sets $U_{1},U_{2},U_{3}$: for each $U_{i}$, there exists $f_{i}\in F$ such that $gf_{i}U_{i}\subseteq\Pi_{o}^{F}(go,r)$.
$go$$gfo$$o$$\Pi_{o}^{F}(go,r)\ni\ gf\xi$$\leq r$$\leq r$$z_{n}\rightarrow\xi\in U$$gfz_{n}$$g\textrm{Ax}(f)$
Figure 12. Mass transport of $U$ into shadows
Since $B_{\xi}(\cdot,\cdot)$ is Lipschitz, we obtain a constant $\theta=\theta(\|Fo\|)>0$ from (23) such that for each $1\leq i\leq 3$, we have
$\mu_{1}(f_{i}U_{i})\geq\theta\cdot\mu_{f_{i}}(f_{i}U_{i})\geq\theta\cdot\lambda^{-1}\cdot\mu_{1}(U_{i})$
where the universal constant $\lambda$ comes from the almost $G$-equivariance (22). Thus,
$\sum_{1\leq i\leq 3}\mu_{1}(f_{i}U_{i})\geq\theta\lambda^{-1}\cdot\mu_{1}(U)/3$
so $f_{i}U_{i}\subseteq g^{-1}\Pi_{o}^{F}(go,r)$ as above shows
$\mu_{1}(g^{-1}\Pi_{o}^{F}(go,r_{0})\cap\mathcal{C})\geq{\theta}\lambda^{-1}\cdot\mu_{1}(U)/3.$
Using again (22) and (23), we have
$\begin{array}[]{rl}\mu_{1}(\Pi_{o}^{F}(go,r_{0})\cap\mathcal{C})&\geq\lambda^{-1}\cdot\mu_{g^{-1}}(g^{-1}\Pi_{o}^{F}(go,r_{0})\cap\mathcal{C})\\ \\ &\geq\lambda^{-2}\cdot e^{-\omega\cdot d(o,go)}\cdot\mu_{1}(g^{-1}\Pi_{o}^{F}(go,r_{0}))\end{array}$
where the last line uses $B_{\xi}(g^{-1}o,o)\leq d(o,go)$. Setting $M:=\theta\cdot\lambda^{-3}\cdot\mu_{1}(U)/3$, we get the lower bound:
$\mu_{1}(\Pi_{o}^{F}(go,r_{0})\cap\mathcal{C})\geq M\cdot e^{-\omega\cdot d(o,go)}.$
We remark here that by the same proof, the subset $\mathcal{C}$ could be replaced with any $G$-invariant subset of positive $\mu_{1}$-measure. We restrict to the $\mathcal{C}$ satisfying Assump D only for the upper bound.
2. Upper Bound. Fix $r\geq r_{0}$. Given $\xi\in g^{-1}\Pi_{o}(go,r)\cap\mathcal{C}$, there is a sequence of $z_{n}\in\mathrm{Y}$ tending to $\xi$ such that $\gamma_{n}\cap B(o,r)\neq\emptyset$ for $\gamma_{n}:=[g^{-1}o,z_{n}]$. By Assump D, we obtain
$|B_{\xi}(g^{-1}o,o)-d(g^{-1}o,o))|\leq 2r+\epsilon.$
As $\mu_{1}(g^{-1}\Pi_{o}(go,r))\leq\|\mu_{1}\|=1$, the upper bound is given as follows:
[…]
The proof of lemma is complete. ∎
Remark 6.5 (Alternative proof).
We sketch another proof for the lower bound which was carried out in Teichmüller space [TY21]. In the proof of Lemma 3.27, as $\langle f\rangle\cdot\partial K=\partial{\mathrm{Y}}\setminus[f^{\pm}]$, the set $\partial K\cap\mathcal{C}$ has positive $\mu_{1}$-measure. By Lemma 3.29, there exists an open set $U$ of $[f^{+}]$ of positive $\mu_{1}$-measure so that $U\subseteq\Pi_{g^{-1}o}^{F}(o,r)$, and then the remaining proof proceeds as above.
Standing Assumption
From now on, we assume further that the constant $r>0$ in Convention 4.1 satisfies the Shadow Lemma 6.4.
As all boundary points in the horofunction boundary are non-pinched by Lemma 5.5, we obtain a full version of shadow lemma.
Lemma 6.6.
Let $\{\mu_{g}\}_{g\in G}$ be a $\omega$-dimensional $G$-equivariant conformal density for some $\omega>0$ on the horofunction boundary $\partial_{\mathrm{H}}{\mathrm{Y}}$. Then there exists $r_{0}>0$ such that
$\begin{array}[]{rl}\exp(-\omega\cdot d(o,go))\prec\mu_{1}(\Pi_{o}^{F}(go,r))\leq\mu_{1}(\Pi_{o}(go,r))\prec_{r}\exp(-\omega\cdot d(o,go))\end{array}$
for any $go\in Go$ and $r\geq r_{0}$.
Corollary 6.7.
Under the asusmption as Lemma 6.6, $\mu_{1}$ has no atoms at conical points in $\partial_{\mathrm{H}}{\mathrm{Y}}$.
Proof.
Recall that any conical point $\xi\in\Lambda_{r}^{F}(Go)$ is contained in the $(r,F)$-shadows of an unbounded sequence of orbital points in $Go$. By Lemma 3.29, the $(r,F)$-shadows are sandwiched between open sets, so by increasing $r$, we can assume that $\xi$ is contained in a sequence of open sets whose $\mu_{1}$-measure tends to 00. This shows that $\xi$ is not atom. ∎
By the existence of $\omega_{G}$-dimensional conformal density in Lemma 6.3, we obtain a upper bound on the growth of balls first proved in [Yan19], by other methods without using PS-measures.
Proposition 6.8.
Let $\{\mu_{g}\}_{g\in G}$ be a $\omega$-dimensional $G$-quasi-equivariant, quasi-conformal density on $\partial{\mathrm{Y}}$ for $\omega>0$. Assume that $\mu_{1}$ charges positive measure on $\mathcal{C}$. Then for any $n\geq 0$,
$|N(o,n)|\prec\exp(\omega n).$
In particular, $\omega\geq\omega_{G}$ and $|N(o,n)|\prec\exp(\omega_{G}n).$
To prove this result, we need the following technical result, which shall be also used in later on. Recall that $[\Pi_{o}^{F}(v,r)]$ denotes the union of []delimited-[][$\cdot$][ ⋅ ]-classes of $\xi\in\Pi_{o}^{F}(v,r)$.
Lemma 6.9.
There exist $\hat{r},R>0$ with the following property. Assume that $\|Fo\|_{\min}\gg 0$. Let $v,w\in Go$ such that the intersection
$\xi\in[\Pi_{o}^{F}(v,r)]\cap[\Pi_{o}^{F}(w,r)]\cap\mathcal{C}$
is non-empty. If $d(o,v)\leq d(o,w)$, then $d(v,[o,w])\leq R$ and $\Pi_{o}^{F}(w,r)\subseteq\Pi_{o}^{F}(v,\hat{r})$.
Proof.
Write $v=go,w=g^{\prime}o$. By definition, there exist $x_{m}\in\Omega_{o}^{F}(go,r),y_{m}\in\Omega_{o}^{F}(g^{\prime}o,r)$ such that $x_{m}$ and $y_{m}$ converge to $[\xi]$. Let $X=g\textrm{Ax}(f)$ and $Y=g^{\prime}\textrm{Ax}(f^{\prime})$ be the corresponding $(r,F)$-barriers satisfying
(26) $go,gfo\in N_{r}([o,x_{m}]),\quad g^{\prime}o,g^{\prime}f^{\prime}o\in N_{r}([o,y_{m}])$
Case 1. Assume that $X=Y$. For $go,g^{\prime}o\in X=Y$, $[o,go]$ and $[o,g^{\prime}o]$ intersect the $C$-ball at $\pi_{X}(o)$ by Lemma 2.2. As $d(o,go)\leq d(o,g^{\prime}o)$ is assumed, the Morse quasi-geodesic $X$ implies the existence of a constant $R=R(C)>0$ independent of $go,g^{\prime}o$ such that $d(go,[o,g^{\prime}o])\leq R$.
$o$$go$$g^{\prime}o$$x_{m}\in\Omega_{o}^{F}(go,r)$$\xi\in\mathcal{C}$$y_{m}\in\Omega_{o}^{F}(g^{\prime}o,r)$$gfo$$g^{\prime}f^{\prime}o$$X$$Y$
Figure 13. Case (2) in the proof of Lemma 6.9
Case 2. Assume now that $X\neq Y\in\mathscr{F}$. As $\mathscr{F}$ has bounded intersection, we have $[\Lambda X]\cap[\Lambda Y]=\emptyset$. For concreteness, assume that $\xi\notin[\Lambda X]$. By Corollary 3.23, we have $d_{X}(x_{m},y_{m})\leq C$ for all $m\gg 0$. By triangle inequality for $d_{X}$, we have $d_{X}(o,y_{m})\geq d_{X}(o,x_{m})-C$, so Lemma 2.3 shows
$\|N_{C}(X)\cap[o,y_{m}]\|\geq d_{X}(o,x_{m})-5C\geq L-4r-5C>0.$
By Lemma 2.2, the entry points of $[o,x_{m}]$ and of $[o,y_{m}]$ into $N_{C}(X)$ have a distance at most $2C$. As $\|N_{C}(X)\cap N_{C}(Y)\|\leq B$ for some $B=B(C)$, we see that $[o,y_{m}]$ exits $N_{C}(X)$ within at most $B$ distance to the entry point in $N_{C}(Y)$. By (26), we have $d(go,[o,x_{m}]),d(g^{\prime}o,[o,y_{m}])\leq r$. Thus, there exists a constant $R=R(C)>0$ such that $d(go,[o,g^{\prime}o])\leq R$.
Apply Lemma 2.9 to the geodesic $\alpha:=[o,x_{m}]$ and $\beta:=[o,y_{m}]$, where $\alpha$ contains $(r,f)$-barrier $go\in X$. Recalling $d_{X}(x_{m},y_{m})\leq C$ as above, we have $go$ is an $(\hat{r},f)$-barrier for $[o,y_{m}]$ with $m\gg 0$. This implies by definition that $\Pi_{o}^{F}(w,r)\subseteq\Pi_{o}^{F}(v,\hat{r})$. Hence, the proof is complete. ∎
Consider the annulus of radius $n$ centered at $o\in\mathrm{Y}$ with width $\Delta\geq 1$:
(27) $A(o,n,\Delta)=\{v\in Go:|d(o,v)-n|\leq\Delta\}$
Lemma 6.10.
For given $\Delta>0$, there exists $N=N(\Delta)$ such that for every $n\geq 1$, any $\xi\in\mathcal{C}$ is contained in at most $N$ shadows $\Pi_{o}^{F}(v,r)$ where $v\in A(o,n,\Delta)$.
Proof.
Indeed, let $v,w\in A(o,n,\Delta)$ such that $\xi\in\Pi_{o}^{F}(v,r)\cap\Pi_{o}^{F}(w,r)$. Lemma 6.9 gives an constant $R>0$ so that $d(v,[o,w])\leq R$. Taking on account of the relation
$d(o,v),\;d(o,w)\in[n-\Delta,n+\Delta],$
there exists $D=D(\Delta,R)$ such that $d(v,w)\leq D$. Therefore it suffices to set $N=|N(o,D)|$ to complete the proof. ∎
All the ingredients are prepared for the proof of Proposition 6.8.
Proof of Proposition 6.8.
By Lemma 6.10, that every point $\xi\in\mathcal{C}$ are contained in a uniformly finite number of shadows. Hence,
$\sum_{v\in A(o,n,\Delta)}\mu_{1}(\Pi_{o}^{F}(v,r)\cap\mathcal{C})\prec\mu_{1}(\mathcal{C})$
which by the shadow lemma gives $\sharp A(o,n,\Delta)\leq e^{n\omega}$, and thus the desired estimates on $|N(o,n)|$. ∎
6.3. Shadow Principle
Let $N(G)=\{g\in\textrm{Isom}(\mathrm{Y}):gGg^{-1}=G\}$ be the normalizer of $G$ in $\textrm{Isom}(\mathrm{Y})$. Following Roblin, the next result is called Shadow Principle, which was first proved in CAT(-1) spaces [Rob02]. The shadow lemma is a special case where $x,y\in Go$ and $\|\mu_{x}\|=\|\mu_{y}\|$.
Lemma 6.11 (Shadow Principle).
Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be a $\omega$-dimensional $G$-quasi-equivariant quasi-conformal density supported on $\partial{\mathrm{Y}}$ with $\mu_{x}(\mathcal{C})>0$ for some $\omega>0$. Let $Z=\Gamma o$ for $\Gamma=N(G)$. Then there exists $r_{0}>0$ such that
$\begin{array}[]{rl}\mu_{y}(\mathcal{C})e^{-\omega d(x,y)}\;\prec_{\lambda}\;\mu_{x}(\Pi_{x}^{F}(y,r)\cap\mathcal{C})\;\leq\mu_{x}(\Pi_{x}(y,r)\cap\mathcal{C})\;\prec_{\lambda,\epsilon,r}\;\mu_{y}(\mathcal{C})e^{-\omega d(x,y)}\end{array}$
for any $x,y\in Z$ and $r\geq r_{0}$, where $\epsilon,\lambda$ are given respectively in Assump D and Definition 6.1.
Proof.
Write explicitly $x=g_{1}o$ and $y=g_{2}o$ for $g_{1},g_{2}\in\Gamma$. Set $U=\mathcal{C}$. Repeating the proof for the lower bound of Lemma 6.4 with $g:=g_{1}^{-1}g_{2}$, we decompose the set $U$ into a disjoint union of three sets $U_{1},U_{2},U_{3}$: for each $U_{i}$, there exists $f_{i}\in F$ such that
(28) $g_{2}f_{i}U_{i}\subseteq\Pi_{g_{1}o}^{F}(g_{2}o,r)\cap\mathcal{C}.$
By quasi-conformality (23) for $d(o,f_{i}o)\leq L,$ there exists $\theta=\theta(L)>0$ such that for each $1\leq i\leq 3$,
$\mu_{g_{2}o}(g_{2}f_{i}U_{i})\geq\theta\cdot\mu_{g_{2}f_{i}o}(g_{2}f_{i}U_{i}).$
By applying the $G$-quasi-equivariance (22) to the above right-hand with the element $g_{2}f_{i}^{-1}g_{2}^{-1}\in G$, we obtain
(29) $\mu_{g_{2}o}(g_{2}f_{i}U_{i})\geq\lambda^{-1}\theta\cdot\mu_{g_{2}o}(g_{2}U_{i}).$
($g_{2}f\in\Gamma\setminus G$ cannot be cancelled directly for $\mu_{x}$ is not known $\Gamma$-equivariant). Summing (29) over $1\leq i\leq 3$,
$\sum_{1\leq i\leq 3}\mu_{g_{2}o}(g_{2}f_{i}U_{i})\geq\lambda^{-1}\theta\cdot\mu_{g_{2}o}(g_{2}U).$
As $U=\mathcal{C}$ is $\Gamma$-invariant and so $g_{2}U=U$, we obtain from (28):
$\mu_{g_{2}o}(\Pi_{r}^{F}(g_{1}o,g_{2}o))\geq\lambda^{-1}{\theta}\cdot\mu_{g_{2}o}(g_{2}U)/3\geq\lambda^{-1}\theta\cdot\mu_{g_{2}o}(U)/3.$
By quasi-conformality (23) again,
$\begin{array}[]{rl}\mu_{g_{1}o}(\Pi_{g_{1}o}^{F}(g_{2}o,r)\cap\mathcal{C})&\geq\lambda^{-2}\cdot e^{-\omega\cdot d(g_{1}o,g_{2}o)}\cdot\mu_{g_{2}o}(\Pi_{g_{1}o}^{F}(g_{2}o,r))\\ \\ &\geq M\cdot\mu_{g_{2}o}(U)\cdot e^{-\omega\cdot d(g_{1}o,g_{2}o)}.\end{array}$
where $M:=\lambda^{-3}\cdot\theta/3$. The upper bound proceeds exactly as in Lemma 6.4 so we omit it. ∎
7. Hopf–Tsuji–Sullivan Dichotomy
This section is devoted to the proof of Theorem 1.14, the so-called Hopf–Tsuji–Sullivan dichotomy, for the conformal density for groups with contracting elements.
Let $G\curvearrowright\mathrm{Y}$ as in (1), equipped with a non-elementary convergence boundary $\partial{\mathrm{Y}}$. Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be any $\omega$-dimensional $G$-quasi-equivariant quasi-conformal density charging positive measure on the set $\mathcal{C}$ of non-pinched points, where $\mathcal{C}$ satisfies Assump C and Assump D. By Lemma 6.8, it is necessary that $\omega\geq\omega_{G}$.
The proof is based on the construction of a family of visual spheres described in the sequel. A similar strategy was used by Tukia [Tuk94b] for real hyperbolic spaces whose arguments are hard to be implemented in general metric spaces. We are eventually turned to the technique of projection complex.
7.1. Outline of the proof
Before giving the details, let us outline the key feature we want for visual spheres, and how it is used in the equivalence between divergence of Poincaré series and positivity of conical points, the core of Theorem 1.14.
The general principle to prove the equivalence is well-known in the fields, which relies on a certain variation of the Borel-Cantelli lemma: if
$\sum_{i=1}^{\infty}\mu_{o}(A_{n})<\infty$
then setting $\Lambda=\cap_{n\geq 1}\cup_{i\geq n}A_{i}$, we have $\mu_{o}(\Lambda)=0$.
In our specific setup, $A_{n}$ is usually the union of shadows $\Pi_{o}(v,r)$ over a sphere $v\in S_{n}$ (which is actually the visual sphere to be precised in next subsection), then $\Lambda$ coincides with the set of conical points, which is by definition contained in infinitely many shadows. Recall that
$\mathrm{e}^{-\omega_{G}d(o,v)}\asymp\mu_{o}(\Pi_{o}(v,r))$
from the Shadow Lemma 6.4. As a consequence, this implies the easy direction: if the conical point has positive measure, the Poincaré series $\sum_{g\in G}\mathrm{e}^{-\omega_{G}d(o,go)}$ is divergent.
The converse direction is the difficult part, boiling down to secure a converse to Borel-Cantelli lemma. The basic idea which is standard in the community is to establish a certain independence between the events $A_{n}$ such as
(30) $\mu_{o}(A_{n+m})\leq c\mu_{o}(A_{n})\mu_{o}(A_{m})$
for a constant $c$ independent of $n$. Therefore, it becomes crucial how to choose the “sphere” $S_{n}$ with this desired property, which usually hinges on the hyperbolicity.
In our setup, let us describe the scenario of constructing $S_{n}$ as follows. Standing on a base point $o$, the first visual sphere $S_{1}$ consists of the set of vertices in $Go$ visible from $o$. In general, $S_{1}$ would be an infinite set, as $Go$ is typically not a subset in any sense of quasi-convexity. This is actually the advantage of the construction, rather than a drawback, to make the above inequality (30) plausible.
Inductively, the next visual sphere $S_{n}$ would be the set of vertices, which are visually blocked by the previous one $S_{n-1}$.
This construction could be worked out with explicit hyperbolicity, which was done by Tukia [Tuk94b] for real hyperbolic spaces. However, if the construction is similarly performed in a general metric space, one might run into cumbersome technicalities due to the lack of hyperbolicity of the ambient space.
As fore-mentioned, we shall implement the construction of visual spheres using the genuine sphere $\bar{S}_{n}$ in the projection complex in its combinatorial metric. To be precisely, the visual sphere, to be denoted by $T_{n}$ below, are the set of orbits in $Go$ carried by the sphere $\bar{S}_{n}$. With hyperbolicity of $\mathcal{P}_{K}(\mathscr{F})$, this allows us to use our version of Shadow Lemma to prove (30) in Lemma 7.5 in the end.
Let us now precise the construction of the visual spheres, which is the main novelty in this proof.
7.2. Family of visual spheres
Let $F\subseteq G$ be a set of three independent non-pinched contracting elements and =$\mathscr{F}=\{g\textrm{Ax}(f):g\in G,f\in F\}$ be the family of their $G$-translated $C$-contracting axes.
By the construction, the quasi-tree of spaces $\mathcal{C}(\mathscr{F})$ contains totally geodesically embedded copies of every $U\in\mathscr{F}$, and the collection $U\in\mathscr{F}$ is the vertex set of the projection complex $\mathcal{P}_{K}(\mathscr{F})$. We are therefore making the following.
Convention 7.1.
In the sequel, $U\in\mathscr{F}$ denotes either the subset $U$ in $\mathrm{Y}$ or $U$ in $\mathcal{C}(\mathscr{F})$, so the points $u\in U$ are understood according to the context. The corresponding vertex in $\mathcal{P}_{K}(\mathscr{F})$ is denoted by $\bar{u}$.
We fix a preferred axis $O=\textrm{Ax}(f)\in\mathscr{F}$, and the corresponding base vertex $\bar{o}$ in $\mathcal{P}_{K}(\mathscr{F})$. The basepoint $o\in\mathrm{Y}$ is contained in the axis $O=\textrm{Ax}(f)$.
Visual sphere construction
For an integer $n\geq 0$, consider the sphere of radius $n$ centered at the preferred base vertex $\bar{o}$ in the projection complex:
$\bar{S}_{n}:=\{\bar{v}\in\mathcal{P}_{K}(\mathscr{F}):d(\bar{o},\bar{v})=n\}$
The $n$-th visual sphere of orbital points in $Go$ is defined as
$T_{n}:=\bigcup_{\bar{v}\in\bar{S}_{n}}V$
where $V$ is of the form $g\textrm{Ax}(f)=gE(f)o$. That is, $T_{n}$ is the union of all $V\in\mathscr{F}$ corresponding to $\bar{v}\in\bar{S}_{n}$. If $go\in U$, we say that $U$ (resp. $\bar{u}$) are an associated (non-unique) vertex space (resp. vertex).
We now introduce some auxiliary sets. For any $V\in\mathscr{F}$ and $L,\Delta>0$, denote
(31) $\begin{array}[]{ll}V_{L}&:=\quad\{v\in V:d_{V}(o,v)>L\}\\ T_{n}^{\prime}(L)&:=\quad\bigcup_{\bar{v}\in\bar{S}_{n}}V_{L}\end{array}$
(recall that, according to Convention 7.1, $V$ is the axis in $\mathrm{Y}$ corresponding to the vertex $\bar{v}$.) By definition, $T_{n}^{\prime}(L)$ forms an $L$-net in $T_{n}$: any point in $T_{n}$ has a distance less than $L$ to $T_{n}^{\prime}$.
[…]
Figure 14. Visual spheres in projection complex (left) and there lifts with $L$-truncation in the original space (right)
We are actually going to focus on a much smaller subset. Let us fix $\Delta:=\min\{d(o,go):g\in E(f),f\in F\}$, so for any $L>0$, the annulus set
$\begin{array}[]{rl}V(L,\Delta)&:=\;\{v\in V:|d_{V}(o,v)-L|\leq\Delta\}\end{array}$
is always non-empty, and has at most $|A(o,L,\Delta)|$ elements. We refer to Fig. 14 for illustrating these sets.
In contrast with the subset $T_{n}^{\prime}(L)$, the union of $V(L,\Delta)$ over $\bar{v}\in\bar{S}_{n}$ is definitely not a net of $T_{n}$, because $V(L,\Delta)$ truncates the part outside the ball $B(o,L+\Delta)$. Thanks to the linear growth of $V$, $V(L,\Delta)$ still contributes a major proportion of the Poincaré series associated to $V$. This fact stated below shall be useful in the proof of Lemma 7.4.
Lemma 7.2.
For any $\omega>0$ and $L>0$, there exists $\theta=\theta(L,\omega)>0$ such that for any $V\in\mathscr{F}$, we have
$\sum_{v\in V}e^{-\omega d(o,v)}\leq\theta\sum_{v\in V(L,\Delta)}e^{-\omega d(o,v)}.$
Proof.
Write $V=gE(f)\cdot o$ for some $g\in G$. Let $u\in V$ such that $d(o,u)=d(o,V)$, so $|d(u,v)-L|\leq\Delta$ for $v\in V(L,\Delta)$. The $C$-contracting property further implies for any $v\in V$:
$|d(o,v)-d(o,u)-d(u,v)|\leq 4C.$
Thus, by the triangle inequality,
$\begin{array}[]{rl}\displaystyle\sum_{v\in V}e^{-\omega d(o,v)}&\prec_{C,\Delta}e^{-\omega d(o,u)}\displaystyle\sum_{n\geq 1}|A(o,n,\Delta)\cap\textrm{Ax}(f)|e^{-\omega n}.\end{array}$
As the maximal elementary group $E(f)$ is virtually $\mathbb{Z}$ and $\textrm{Ax}(f)=E(f)o$ is a quasi-geodesic, we see $|A(o,n,\Delta)\cap\textrm{Ax}(f)|$ is a linear function of $n$. For any given $L>0$, we ignore the contribution from $V_{L+\Delta}$ in the series on the right-hand side, so that
$\begin{array}[]{rl}\displaystyle\sum_{v\in V}e^{-\omega d(o,v)}\leq\theta\cdot\left(\sum\limits_{v\in V(L,\Delta)}e^{-\omega d(o,v)}\right)\end{array}$
where the implied constant $\theta$ depends only on $L,C,\Delta$ and $E(f)$. ∎
The following lemma is analogous to Lemma 6.10, but the proof is much more involved. Recall the definition of $T_{n}^{\prime}(L)$ is given in (31).
Lemma 7.3.
There exists $L\gg 0$ such that for any $\xi\in\mathcal{C}$ and $n>0$, the set of orbital points
$T_{n}(\xi)\quad:=\quad\{v\in T_{n}^{\prime}(L):\;\xi\in\Pi_{o}(v,r)\}$
is contained in at most one axis $V$ in $\mathscr{F}$. That is, $T_{n}(\xi)=V_{L}$.
Proof.
Let $go,ho\in T_{n}^{\prime}(L)$ so that $\xi\in\Pi_{o}(go,r)\cap\Pi_{o}(ho,r)$. By definition of $T_{n}^{\prime}(L)$, there are $U,V\in\mathscr{F}$ so that $go\in U$ and $ho\in V$ and the associated vertices $\bar{u},\bar{v}$ in the projection complex lie on the $n$-sphere $\bar{S}_{n}$. Moreover, we have
(32) $d_{U}(o,go)\geq L,\quad d_{V}(o,ho)\geq L.$
Arguing by way of contradiction, assume that $U\neq V$, which thus a bounded intersection. We may assume further that $d(o,U)\leq d(o,V)$ for concreteness.
According to definition of $\xi\in\Pi_{o}(go,r)\cap\Pi_{o}(ho,r)$, there exist two sequences of points in the corresponding cones
$\{x_{m}\}\subseteq\Omega_{o}(go,r),\quad\{y_{m}\}\subseteq\Omega_{o}(ho,r)$
which tend to $\xi$ as $m\to\infty$. We refer to Fig. 15 for the configuration of these points.
By definition, $d_{U}(o,go)\geq L$ for $go\in U_{L}$, which yields $\|N_{r}(U)\cap[o,go]\|\geq L-4r$ by Lemma 2.4. The contracting property of $U$ implies that the entry point of $[o,x_{m}]$ at $N_{r}(U)$ is $2C$-close to $\pi_{U}(o)$ by Lemma 2.3(3). As $go$ lies within a distance $r$ to $[o,x_{m}]$, we obtain
$\|N_{r}(U)\cap[o,x_{m}]\|\geq\|N_{r}(U)\cap[o,go]\|-2C-r\geq L-5r-2C.$
Thus, setting $M=9r+2C$, Lemma 2.4 shows
$d_{U}(o,x_{m})\geq\|N_{r}(U)\cap[o,x_{m}]\|-4r\geq L-M.$
Similarly, the same reasoning shows $d_{V}(o,y_{m})\geq L-M.$
As $U\neq V$ is assumed, we have $[\Lambda U]\cap[\Lambda V]=\emptyset$. Indeed, recall that $\Lambda U$ means the limit set of the axis $U\in\mathscr{F}$, which is here the fixed points of a contracting element. By Lemma 3.20, distinct non-pinched contracting elements have disjoint fixed points, so $[\Lambda U]\cap[\Lambda V]=\emptyset$ follows.
For concreteness, assume that $\xi\notin[\Lambda U]$; the case $\xi\notin[\Lambda V]$ is similar.
[…]
Figure 15. If $d(o,U)\gg d(o,V)$, we shall see that $U$ appears before $V$ in $\mathcal{P}_{K}(\mathscr{F})$: $\bar{u}$ lies on $[\bar{o},\bar{v}]$. This contradicts the assumption of $\bar{u},\bar{v}\in\bar{S}_{n}$.
By Assump C, we have $d(o,[x_{m},y_{m}])\to\infty$ for $\xi\in\mathcal{C}$. By Lemma 3.23, as $\xi\notin\Lambda U$, we have $d_{U}(x_{m},y_{m})\leq C$ for $m\gg 0$. The triangle property for $d_{U}$ then gives
$d_{U}(o,y_{m})\geq d_{U}(o,x_{m})-d_{U}(x_{m},y_{m})\geq L-C-M$
If we choose $L\gg M+5C$, then $[o,y_{m}]\cap N_{C}(U)\neq\emptyset$ by Lemma 2.3. Let $z$ be the exit point $[o,y_{m}]$ at $N_{C}(U)$, so $d(z,\pi_{U}(y_{m}))\leq 2C$ by Lemma 2.2.
As $U\neq V$ has bounded intersection and $d(o,U)\leq d(o,V)$, $[o,y_{m}]$ exits $N_{C}(U)$ before entering $N_{C}(V)$ with a possible overlap bounded by a constant $B$. As $d(ho,[o,y_{m}])\leq r$ and $ho\in V$, there exists a constant still denoted by $M=M(r)>0$ such that
$d_{U}(o,ho)\geq d_{U}(o,y_{m})-M.$
Choose $L>\hat{K}+2M+C$, where $\hat{K}$ is given by Lemma 2.26. Thus, $d_{U}(o,ho)\geq\hat{K}$, so that $\bar{u}$ appears in any geodesic $[\bar{o},\bar{v}]$. This is a contradiction, as $d(\bar{o},\bar{u})=d(\bar{o},\bar{v})=n$.
In conclusion, the set $T_{n}(\xi)$ is contained in one axis $U=V$. In particular, $T_{n}(\xi)\subseteq V_{L}$ follows. The result is proved. ∎
Thanks to Lemma 7.3, any boundary point is contained in a uniform number of shadows over $T_{n}^{\prime}(L)$, so we deduce the following main corollary from it.
Lemma 7.4.
There exists a uniform constant $c>0$ independent of $n\geq 1$ such that
(33) $\forall n\geq 1:\quad\sum\limits_{go\in T_{n}}e^{-\omega d(o,go)}\leq c.$
Proof.
Fix a constant $L>0$ satisfying Lemma 7.3, so if $T_{n}(\xi)$ for $\xi\in\mathcal{C}$ is non-empty, then it is contained in one $V\in\mathscr{F}$. Precisely, $T_{n}(\xi)\subseteq V_{L}$. By Lemma 6.9, there exists $N=N(L)$ such that any point $\xi\in\mathcal{C}$ are shadowed by at most $N$ elements in $V(L,\Delta)$:
$\sum\limits_{\bar{v}\in\bar{S}_{n}}\left(\sum\limits_{u\in V(L,\Delta)}\mu_{1}(\Pi_{o}^{F}(u,r))\right)\leq N\cdot\mu_{o}(\mathcal{C})\leq N$
By Lemma 7.2, the majority of the Poincaré series is concentrated over $V(L,\Delta)$. We thus obtain
$\sum\limits_{v\in T_{n}}e^{-\omega d(o,v)}\leq\theta\sum\limits_{\bar{v}\in\bar{S}_{n}}\left(\sum\limits_{u\in V(L,\Delta)}e^{-\omega d(o,u)}\right)$
where $V$ is the axes associated to $\bar{v}$.
The conclusion then follows from the shadow lemma 6.4 that $\mu_{1}(\Pi_{o}^{F}(u,r))\asymp\mathrm{e}^{-\omega d(o,u)}$. ∎
Recall the definition of partial Poincaré series over the $n$-th visual sphere $T_{n}$:
$\mathcal{P}_{T_{n}}(s,o,o)=\sum_{v\in T_{n}}e^{-\omega d(o,v)}$
The following crucial lemma shall be used to write the Poincaré series as geometric series in Step 1 below in the next subsection.
Lemma 7.5.
For any integer $m,n\geq 1$ and for any $s>\omega_{G}$, we have
(34) $\mathcal{P}_{T_{n}}(s,o,o)\asymp\mathcal{P}_{T_{n+1}}(s,o,o)$
(35) $\mathcal{P}_{T_{n+m}}(s,o,o)\prec\mathcal{P}_{T_{n}}(s,o,o)\cdot\mathcal{P}_{T_{m}}(s,o,o)$
where the implied constants do not depend on $m,n$ and $s$.
Proof.
(1). We start by proving the $\prec$” direction of (34).
Let $\hat{K}=\hat{K}(K)\geq B$ be given by Lemma 2.26, where $B$ is the bounded projection constant of $\mathscr{F}$. Consider the $(2\hat{K})$-net $T_{n}^{\prime}:=T_{n}^{\prime}(2\hat{K})$ in $T_{n}$, so we have
$\mathcal{P}_{T_{n}}(s,o,o)\asymp_{\hat{K}}\mathcal{P}_{T_{n}^{\prime}}(s,o,o).$
The goal of the proof is to prove for any $go\in T_{n}^{\prime}$, there exists $f^{\prime}\in F$ such that $gf^{\prime}o\in T_{n+1}$. Recall that if $V$ is an axis given by $g\textrm{Ax}(f)$, we say that $f$ is the type of the axis $V$ or the vertex $\bar{v}$.
Indeed, according to definition of $go\in T_{n}^{\prime}(2\hat{K})$, there exist an associated vertex $\bar{u}\in\bar{S}_{n}$ of type $f\in F$ and the vertex space $U=g\textrm{Ax}(f)$ such that $d_{U}(o,go)>2\hat{K}.$ By Lemma 2.17 there exists $f^{\prime}\neq f\in F$ such that $gf^{\prime}$ labels a quasi-geodesic. The axis $V:=g\textrm{Ax}(f^{\prime})$ is adjacent to $U$ in $\mathcal{P}_{K}(\mathscr{F})$, so $d(\bar{u},\bar{v})=1$ and thus $|d(\bar{v},\bar{o})-n|\leq 1$. We claim that $\bar{v}\in\bar{S}_{n+1}$. The point is to exclude the possibility of $d(\bar{v},\bar{o})\leq n$.
Since $U=g\textrm{Ax}(f)$ and $V=g\textrm{Ax}(f^{\prime})$ have $B$-bounded intersection for some $B>0$, we obtain that $[go,gf^{\prime}o]$ has $B$-bounded projection to $U$. This implies
$d_{U}(O,V)\geq d(o,go)-B\geq 2\hat{K}-B\geq\hat{K}$
so by Lemma 2.26, $\bar{u}$ is contained in any geodesic from $\bar{o}$ to $\bar{v}$ in $\mathcal{P}_{K}(\mathscr{F})$.
If $d(\bar{o},\bar{v})\leq n$ was true, we would obtain a contradiction: $d(\bar{u},\bar{o})<n$. Hence, $d(\bar{v},\bar{o})\leq n$ is impossible: that is, $d(\bar{o},\bar{v})=n+1$ and $\bar{v}\in\bar{S}_{n+1}$.
In summary, we proved that for any $go\in T_{n}^{\prime}$ there exists $f^{\prime}\in F$ such that $gf^{\prime}o\in T_{n+1}$ and $d(o,gf^{\prime}o,o)\leq d(o,go)+\|Fo\|$. Moreover, as $Go$ is a discrete set, only a uniform finite number of $go\in T_{n}^{\prime}$ shares the common $gf^{\prime}o\in T_{n+1}$ in a ball of radius $\|Fo\|$. Unveiling the definition of $\mathcal{P}_{T_{n+1}}(s,o,o)$ with the triangle inequality, we see that
$\mathcal{P}_{T_{n+1}}(s,o,o)\geq c\cdot\mathcal{P}_{T_{n}^{\prime}}(s,o,o)$
some constant $c$ depending on $\|Fo\|$. Recall that $\mathcal{P}_{T_{n}^{\prime}}(s,o,o)\asymp_{\hat{K}}\mathcal{P}_{T_{n}}(s,o,o)$. The direction $\prec$” of the inequlity (34) follows as desired.
The other direction would follow from the inequality (35), which is the next goal.
(2). For any $go\in T_{n+m}$, let $\bar{u}\in\bar{S}_{n+m}$ be an vertex for which the associated axis $U$ contains $go$. The vertex $\bar{u}$ is non-unique, as any point of $Go$ is contained in three axis from $\mathscr{F}$.
Consider the standard path $\alpha$ from $\bar{o}$ to $\bar{u}$. By Lemma 2.25, there exists an $(L,B)$-admissible path $\tilde{\alpha}$ between $o$ and $go$ so that the saturation $\mathscr{F}(\tilde{\alpha})$ consists of the axes associated to the vertices on $\alpha$. In particular, we have $[o,go]$ intersects the $r$-neighborhood of every $V\in\mathscr{F}(\tilde{\alpha})$.
Let $\bar{v}\in\bar{S}_{n}\cap\alpha$ so that $d(\bar{v},\bar{o})=n$. To be explicit, write the corresponding axis $V=h\textrm{Ax}(f^{\prime})$ for some $h\in G$ and $f^{\prime}\in F$. By definition, any element $ho\in V$ in the axis $V$ is contained in the $n$-th visual sphere $T_{n}$. Let us choose the one $ho$ so that $d(ho,[o,go])\leq r$ (as $[o,go]$ intersects the $r$-neighborhood of $V$). Thus, denoting $a=h^{-1}g$,
$d(o,go)\geq d(o,ho)+d(o,ao)-2r$
As the standard path $\alpha$ is a $(2,1)$-quasi-geodesic ([BBFS19, Corollary 3.7]), it follows that $|d(\bar{v},\bar{u})-m|\leq N_{0}$ for a universal integer $N_{0}$. Indeed, according to the proof of Theorem 2.23, we can choose $N_{0}=2$, since it is proved that any geodesic passes with a distance at most 2 to any vertex on the standard path.
It is possible that $f^{\prime}\neq f$; that is, $\bar{v}$ might not be in the orbit $G\bar{o}$. Let $\bar{w}$ be the vertex adjacent to $\bar{v}$ so that the associated axis $W$ is of form $h\textrm{Ax}(f)$. As $|d(\bar{w},\bar{u})-m|\leq N_{0}+1$ and $h^{-1}\bar{w}=\bar{o}$, we thus have $h^{-1}go$ lies in the following $(N_{0}+1)$-thicken annulus:
$A_{m}:=\bigcup_{|i-m|\leq N_{0}+1}T_{i}$
That is, we can write $g=h\cdot a$ where $ho\in T_{n}$ and $ao\in A_{m}$. We now see that
$\mathcal{P}_{T_{n+m}}(s,o,o)$ $=\sum_{go\in T_{n+m}}e^{-\omega d(o,go)}$
$\prec\mathcal{P}_{T_{n}}(s,o,o)\cdot\mathcal{P}_{A_{m}}(s,o,o),$
By Lemma 7.4, $\mathcal{P}_{A_{1}}(s,o,o)\leq c_{1}:=2(N_{0}+1)c$ for $m=1$, so we get
$\mathcal{P}_{T_{n+1}}(s,o,o)\leq c_{1}\mathcal{P}_{T_{n}}(s,o,o)$
This completes the proof of the other direction in (34), as promised above. This further implies $\mathcal{P}_{A_{m}}(s,o,o)\leq c_{2}\mathcal{P}_{T_{m}}(s,o,o)$ for some $c_{2}$ depending on $c_{1}$ and $N_{0}$. Hence, the general case in (35) follows. ∎
7.3. Proof of Hopf–Tsuji–Sullivan Dichotomy
Recall that the union $\bigcup_{i\geq 1}T_{i}$ covers the $G$-orbit $Go$ with multiplicity at most 3. The shadow lemma 6.4 then implies
(36) $\mathcal{P}_{G}(\omega,o,o)\quad\asymp\quad\sum\limits_{i\geq 1}\Bigg{(}\sum\limits_{g\in T_{i}}e^{-\omega d(o,go)}\Bigg{)}\quad\asymp_{r}\quad\sum\limits_{i\geq 1}\Bigg{(}\sum\limits_{g\in T_{i}}\mu_{1}\big{(}\Pi_{o}^{F}(go,r)\big{)}\Bigg{)}$
We subdivide the proof of Theorem 1.14 in the following steps.
Before the proof, we remark that different definitions 4.5, 4.3 and 4.9 of conical points are quantitatively equivalent in Lemmas 4.6 and 4.10: that is, each set is contained into the other with appropriate parameters. Thus, we shall insist on the definition 4.3 of conical points via shadows, as the proof uses extensively the shadow lemma. This does not effect the generality of Theorem 1.14, as the positive / null measure charged on conical points shall be established to be equivalent with divergence / convergence of Poincaré series.
First of all, the direction “(2) $\Longrightarrow$ (4)” follows as $\Lambda_{r}^{F}(Go)\subseteq\Lambda_{m}{(Go)}$, and “(3) $\Longrightarrow$ (2)” is trivial. We now give the proof of the other directions.
(1) $\Longrightarrow$ (4)
If the set $\Lambda_{r}^{F}(Go)$ is $\mu_{1}$-null, then we prove that $G$ is of convergence type. Denote
$\Lambda_{n}:=\bigcup\big{\{}\Pi_{o}^{F}(go,r):g\in\cup_{k\geq n}T_{k}\big{\}}.$
We have $\lim_{n\to\infty}\Lambda_{n}\subseteq\Lambda_{r}^{F}(Go)$ by (19). By hypothesis, $\lim_{n\to\infty}\mu_{1}(\Lambda_{n})=0$. Using Lemma 6.4, there exist $n_{0}>0$ and $10<q<$ 1 such that $\theta q\leq 1$ and
$\sum\limits_{g\in T_{n_{0}}}e^{-\omega d(o,go)}<q$
where $\theta>0$ is the implied constant given by Lemma 7.5. For each $k\geq 0$ and $1\leq i\leq n_{0}$, we obtain from (34, 35) that
$\begin{array}[]{ccc}\sum\limits_{k\geq 1}\left(\sum\limits_{v\in T_{kn_{0}}}e^{-\omega d(o,v)}\right)&<\sum\limits_{k\geq 1}\left(\theta\sum\limits_{v\in T_{n_{0}}}e^{-\omega d(o,v)}\right)^{k}&\leq\sum\limits_{k\geq 1}(\theta q)^{k}<\infty\end{array}$
and
$\sum\limits_{v\in T_{kn_{0}+1}}e^{-\omega d(o,v)}<C\left(\sum\limits_{v\in T_{kn_{0}}}e^{-\omega d(o,v)}\right)$
By (36), we obtain that $\mathcal{P}_{G}(\omega,o,o)$ is convergent.
Recall the usual conical points $\Lambda_{c}(G)$ are defined in (6) using the usual shadow $\Pi_{x}(y,r)$ without index $F$:
$\Lambda_{c}(Go):=\bigcup_{r\geq 0}\bigcup_{x\in Go}\left(\limsup_{y\in Go}\Pi_{x}(y,r)\right).$
It is clear that $\Lambda_{r}^{F}(Go)$ is a subset of $\Lambda_{c}(Go)$, so $\mu_{1}(\Lambda_{c}(G))=0$ implies $\mathcal{P}_{G}(\omega,o,o)<\infty$.
(4) $\Longrightarrow$ (1)
If $G$ is of convergent type, then the set $\Lambda_{r}^{F}(Go)$ is $\mu_{1}$-null.
Indeed, assume that $\mu_{1}(\Lambda_{r}^{F}(Go))>0$ and we are going to prove that $\mathcal{P}_{G}(\omega,o,o)$ is convergent. By Corollary 4.8, we could fix the light source at $o$ and increase $r$, so that the following family of sets
$\Lambda_{n}:=\bigcup\limits_{k\geq n}\Bigg{(}\bigcup\limits_{v\in A(o,k,\Delta)}\Pi_{o}^{F}(v,r)\Bigg{)}$
tends to $\Lambda_{r}^{F}(Go)$, where $A(o,n,\Delta)$ is defined in (27). Thus, $\mu_{1}(\Lambda_{n})\geq\frac{1}{2}\mu_{1}(\Lambda_{r}^{F}(Go))$ for $n\gg 0$. By Lemma 6.4, we obtain a uniform lower bound of the partial sum of Poincaré series $\mathcal{P}_{G}(\omega,o,o)$ for any $n\gg 0$
$\bigcup\limits_{k\geq n}\Bigg{(}\bigcup\limits_{v\in A(o,k,\Delta)}e^{-\omega n}\Bigg{)}\succ\mu_{1}(\Lambda_{n})$
contradicting to the convergence of $\mathcal{P}_{G}(\omega,o,o)$.
To prove that $\Lambda_{c}(G)$ is $\mu_{1}$-null, we re-define $\Lambda_{n}$ as in the exact form as above where $\Pi_{o}^{F}(v,r)$ is replaced by $\Pi_{o}(v,r)$. As the shadow Lemma 6.4 works exactly for $\Pi_{o}(v,r)$, the remaining argument proves in the same way that $\Lambda_{c}(Go)$ is $\mu_{1}$-null.
(4) $\Longrightarrow$ (5)
If the set $A:=\Lambda_{r}^{F}(Go)$ or $A:=\Lambda_{c}(G)$ is $\mu_{1}$-positive then it is $\mu_{1}$-full in $\mathcal{C}$: $\mu_{1}(A)=\mu_{1}(\mathcal{C})$.
Suppose to the contrary that the complement $\mathcal{C}\setminus A$ is $\mu_{1}$-positive. The restriction of $\{\mu_{x}\}_{x\in\mathrm{Y}}$ on $\mathcal{C}\setminus A$ gives a conformal density $\{\nu_{x}\}_{x\in\mathrm{Y}}$ of the same dimension $\omega$, so that $\nu_{1}(\Lambda_{r}^{F}(Go))=0$. By “(1) $\Longrightarrow$ (4)”, we have that $G$ is of convergent type: this in turn contradicts the “(4) $\Longrightarrow$ (1)”.
(5) $\Longrightarrow$ (3)
Applying “(1) $\Longrightarrow$ (4)”, we have that for every $\mathscr{F}$ in (3), the set $\Lambda_{r}^{F}(Go)$ is $\mu_{1}$-full. By Corollary 4.17, their countable intersection gives the Myrberg set.
The proof of Theorem 1.14 is complete.
8. Conformal and harmonic measures on the reduced Myrberg set
This section investigates the uniqueness of conformal density with dimension at the critical exponent for divergence group actions. Our goal here is to prove Theorem 1.16 from Introduction, stated below as Lemma 8.5, that conformal measures shall be unique and ergodic, when pushed forward to the reduced horofunction boundary.
Suppose that $G\curvearrowright\mathrm{Y}$ as in (1) is of divergent type, where $\mathrm{Y}$ is compactified by a convergence boundary $\partial{\mathrm{Y}}$. Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be a $\omega$-dimensional $G$-quasi-equivariant quasi-conformal density on $\partial{\mathrm{Y}}$ with $\lambda\geq 1$ given in Definition 6.1.
According to the Hopf–Tsuji–Sullivan dichotomy (Theorem 1.14), as $G\curvearrowright\mathrm{Y}$ is of divergent type, the conformal measures $\mu_{x}$ are fully charged on the set of conical points or even the smaller subset of Myrberg points.
By Lemma 5.7 and 2.30, the finite difference relation restricted on $\Lambda_{m}{(Go)}$ induces a closed surjective map from $\Lambda_{m}{(Go)}$ to $[\Lambda_{m}{(Go)}]$ with compact fibers. This implies that $[\Lambda_{m}{(Go)}]$ is metrizable and second countable.
8.1. Push forward conformal density to the quotient
According to Lemma 4.21, there exists a topological embedding
$[\Lambda_{m}{(Go)}]\longrightarrow\partial\mathcal{C}(\mathscr{F})$
so we could endow $[\Lambda_{m}{(Go)}]$ with the visual metric on the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$. Push forward $\{\mu_{x}\}_{x\in\mathrm{Y}}$ to $[\Lambda_{m}{(Go)}]$ via the quotient map
$\Lambda_{m}{(Go)}\longrightarrow[\Lambda_{m}{(Go)}]$
We obtain a $\omega_{G}$-dimensional $G$-quasi-equivariant quasi-conformal density $\{[\mu_{x}]\}_{x\in\mathrm{Y}}$ on $[\Lambda_{m}{(Go)}]$, where $[\mu_{x}]$ denotes the image measure of $\mu_{x}$. The constant $\lambda^{\prime}$ now depends on $\lambda$ and $C$, as any two horofunctions in a []delimited-[][$\cdot$][ ⋅ ]-class of Myrberg points has $(20C)$-difference by Lemma 5.3. Here the contracting constant $C$ could taken to be the infimum over all $\mathscr{F}$ in (3).
The proof consists in releasing conformal measures as a Hausdorff-Caratheodory measure on $[\Lambda_{m}{(Go)}]$. To that end, we first state a variant of Vitali covering lemma in [Yan22, Appendix A].
Let $B\in\mathbb{B}$ be a family of subsets in a topological space $X$ with extensions $B\subseteq\tilde{B}$ of each $B$. Then $\mathbb{B}$ is called hierarchically locally finite if there exist a positive integer $N>0$ and a disjoint decomposition $\mathbb{B}=\bigsqcup\mathbb{B}_{n}$ for $n\geq 0$ where $\mathbb{B}_{n}$ is possibly empty for certain $n$ such that the following two properties are true:
  1. (1)
    Given $B\in\mathbb{B}_{n}$ for $n\geq 0$, the set $\mathcal{N}(B)$ of $A\in\mathbb{B}_{n-1}\cup\mathbb{B}_{n}\cup\mathbb{B}_{n+1}$ with $A\cap B\neq\emptyset$ has cardinality at most $N,$ where $\mathbb{B}_{-1}:=\emptyset$.
  2. (2)
    If $A\in\mathbb{B}_{n}$ intersects $B\in\mathbb{B}_{m}$ for $n\geq m-2$, then $A$ is contained in the extension $\tilde{B}$ of $B$.
By definition, any subfamily of $\mathbb{B}$ is hierarchically locally finite.
Lemma 8.1.
[Yan22, Lemma A.1] Let $\mathbb{B}$ be a hierarchically locally finite family of subsets in a space $X$. Then there exists a sub-family $\mathbb{B}^{\prime}\subseteq\mathbb{B}$ of pairwise disjoint subsets such that
(37) $\bigcup\limits_{B\in\mathbb{B}}B\subseteq\bigcup_{A\in\mathbb{B}^{\prime}}\left(\bigcup_{B\in\mathcal{N}(A)}\tilde{B}\right).$
8.2. Hierarchically locally finite covers
In order to use Lemma 8.1, we define the family $\mathbb{B}$ of sets $B(v)$ and their extension $\tilde{B}(v)$ as follows
$B(v):=[\Pi_{o}^{F}(v,r)],\quad\tilde{B}(v):=[\Pi_{o}^{F}(v,\hat{r})]$
for every $v\in Go$. Here $r,\tilde{r}$ are given by Convention 4.1.
By abuse of language, we define the diameter of it: $\|B(v)\|:=e^{-d(o,v)}$, even though there may exist no metric on $\partial{\mathrm{Y}}$ inducing it. A subfamily of $\mathbb{B}^{\prime}\subseteq\mathbb{B}$ is called an $\epsilon$-covering, if $\|B\|\leq\epsilon$ for every $B\in\mathbb{B}^{\prime}$.
Lemma 8.2.
For any $v\in Go$, we have
(38) $\mu_{1}(B(v))\asymp\|B(v)\|^{\omega}.$
Proof.
The lower bound on $\mu_{1}(B(v))$ follows directly from the shadow lemma 6.6. For the upper bound, $B(v)$ contains the locus of generic points in $\Pi_{o}^{F}(v,r)$. Since $G$ is of divergent type, the measure $\mu_{1}$ is supported on $\Lambda_{r}^{F}(Go)$, so by Lemma 5.6, any two horofunctions in the same locus of points in $\Pi_{o}^{F}(v,r)$ have uniform difference. Thus, the proof of Lemma 6.4 for the upper bound works to give an uniform coefficient before $e^{-d(o,v)}.$
We define a Hausdorff-Carathéodory outer measure $\mathcal{H}^{\omega}$ on $[\Lambda_{m}{(Go)}]$ relative to $\mathbb{B}$ and the gauge function $r^{\omega}$. Namely, for any $\epsilon>0$ and $A\subset[\Lambda_{m}{(Go)}]$,
(39) $\mathcal{H}_{\epsilon}^{\omega}(A)=\inf\Bigg{\{}\sum_{B\in\mathbb{B}^{\prime}}\|B\|^{\omega}:A\subseteq\bigcup_{B\in\mathbb{B}^{\prime}}B,\;\|B\|\leq\epsilon\Bigg{\}}$
where the infimum is taken over $\epsilon$-cover $\mathbb{B}^{\prime}$ over $A$. Such $\epsilon$-covers exist for any $\epsilon$, since Myrberg points and conical points are contained in infinitely many shadows. Then
$\mathcal{H}^{\omega}(A):=\sup_{\epsilon>0}\mathcal{H}_{\epsilon}^{\omega}(A).$
We now prove that $\mathcal{H}^{\omega}$ is a metric outer measure. Thus, Borel subsets are $\mathcal{H}^{\omega}$-measurable.
Lemma 8.3.
If $A,B$ are disjoint closed subsets in $[\Lambda_{m}{(Go)}]$, then
$\mathcal{H}^{\omega}(A\cup B)=\mathcal{H}^{\omega}(A)+\mathcal{H}^{\omega}(B)$
Proof.
By Lemma 4.21, $[\Lambda_{m}{(Go)}]$ is topologically embedded into the Gromov boundary $\partial\mathcal{C}(\mathscr{F})$. Let $A,B$ be disjoint closed subsets in $\Lambda_{m}{(Go)}$. If $\partial\mathcal{C}(\mathscr{F})$ is equipped with visual metric, it is readily checked that the corresponding $\epsilon$-covers for $A$ and $B$ are disjoint for small enough $\epsilon$. The details is left to the interested reader. Consequently, the conclusion follows from the definition of Hausdorff measures. ∎
Fix a large $L>0$, and subdivide $\mathbb{B}$ into a sequence of (possibly empty) sub-families $\mathbb{B}_{n}$ for $n\geq 0$:
$\mathbb{B}_{n}:=\{B(v)\in\mathbb{B}:nL\leq d(o,v)<(n+1)L\}.$
Lemma 8.4.
There exists a uniform constant $L>0$ such that $\mathbb{B}$ defined as above is a hierarchically locally finite family over $[\Lambda_{m}{(Go)}]$.
Proof.
We verify the two conditions in order. Let $B(v)=[\Pi_{o}^{F}(v,r)]\in\mathbb{B}_{n}$. Applying Lemma 6.10 to the annulus $A(o,n,2L)$, there are at most $N=N(L)$ sets
$B(w)\in\mathbb{B}_{n-1}\cup\mathbb{B}_{n}\cup\mathbb{B}_{n+1}$
which intersect $B(v)$.
Let $B(w)\in\mathbb{B}_{m}$ for $m<n$ intersecting $B(v)$, so $d(v,w)\geq d(o,v)-d(o,w)\geq L$. We have $B(v)\subseteq\tilde{B}(w)=[\Pi_{o}^{F}(w,\hat{r})]$ by Lemma 6.9. Both conditions are thus verified and the lemma is proved. ∎
8.3. Uniqueness of conformal measures
Assuming the action is of divergent type, we are now ready to prove the main result in this Section Theorem 1.16 restated below, saying that the push-forward of PS measures becomes unique on the reduced Myrberg set.
Lemma 8.5.
Let $\{[\mu_{x}]\}_{x\in\mathrm{Y}}$ be a $\omega_{G}$-dimensional $G$-quasi-equivariant quasi-conformal density on the quotient $[\Lambda_{m}{(Go)}]$. Then we have
$\mathcal{H}^{\omega}(A)\asymp[\mu_{1}](A)$
for any Borel subset $A\subseteq[\Lambda_{m}{(Go)}]$, where the implied constants depend on the constant $\lambda$ in Definition 6.1.
In particular, $[\mu_{x}]$ is unique in the following sense: if $[\mu],[\mu^{\prime}]$ are two such quasi-conformal densities, then the Radon-Nikodym derivative $d[\mu_{o}]/d[\mu_{o}^{\prime}]$ is bounded from above and below by a constant depending only on $\lambda$.
Proof.
Take an $\epsilon$-covering $\mathbb{B}^{\prime}\subseteq\mathbb{B}$ of $A$, so that $[\mu_{1}(A)]\leq\sum_{B\in\mathbb{B}^{\prime}}[\mu_{1}](B)$. Letting $\epsilon\to 0$, we obtain from (38) that $[\mu_{1}](A)\prec\mathcal{H}^{\sigma}(A)$.
To establish the other inequality, for any $0<\epsilon<\epsilon_{0}$, let $\mathbb{B}_{1}\subseteq\mathbb{B}$ be an $\epsilon$-covering of $A$. By Lemma 8.1, there exist a disjoint sub-family $\mathbb{B}_{2}$ of $\mathbb{B}_{1}$ and an integer $N>1$ such that
$\mathcal{H}^{\omega}(A)\prec\sum_{B\in\mathbb{B}_{2}}N\cdot\|{B}\|^{\omega}\prec[\mu_{1}](U),$
Letting $\epsilon\to 0$ yields $\mathcal{H}^{\omega}(A)\prec[\mu_{1}](A)$. ∎
As in the usual Hausdorff measure, we define the Hausdorff dimension $\textrm{HDim}(A)$ of a subset $A\subseteq[\Lambda_{m}{(Go)}]$:
(40) $\textrm{HDim}(A):=\sup\{\omega\geq 0:\mathcal{H}^{\omega}(A)=\infty\}=\inf\{\omega\geq 0:\mathcal{H}^{\omega}(A)=0\}$
where $\sup\emptyset:=\infty$ and $\inf\emptyset:=0$ by convention. The following corollary is immediate.
Corollary 8.6.
The Hausdorff dimension of $[\Lambda_{m}{(Go)}]$ equals $\omega_{G}$.
We also obtain the ergodicity from the uniqueness.
Lemma 8.7.
Under the assumption of Lemma 8.5, let $\{[\mu_{x}]\}_{x\in\mathrm{Y}}$ be a $\omega_{G}$-dimensional $G$-quasi-equivariant quasi-conformal density on $[\Lambda_{m}{(Go)}]$. Then $\{[\mu_{x}]\}_{x\in\mathrm{Y}}$ is ergodic with respect to the action of $G$ on $[\Lambda_{m}{(Go)}]$.
Proof.
Let $A$ be a $G$-invariant Borel subset in $[\Lambda_{m}{(Go)}]$ such that $[\mu_{x}](A)>0$. Restricting $[\mu_{x}]$ on $A$ gives rise to a $\omega_{G}$-dimensional conformal density $\nu_{x}:=[\mu_{x}]|_{A}$ on $[\Lambda_{m}{(Go)}]$. By Lemma 8.5, for any subset $X\subseteq[\Lambda_{m}{(Go)}]$, we have $\nu_{x}(X)\prec\mathcal{H}^{\omega}(A\cap X)$. Thus, $[\mu_{x}]([\Lambda_{m}{(Go)}]\setminus A)=0$. ∎
8.4. Myrberg set as Poisson boundary
In this subsection, we explain that the reduced Myrberg set could serve as the Poisson boundary for random walks on groups with contracting elements. This is the content of Theorem 1.19.
Let $G\curvearrowright\mathrm{Y}$ be a proper action with contracting elements as in (1) with a convergence boundary $\partial{\mathrm{Y}}$. Let $\mu$ be a probability measure on $G$ so that its support generates $G$ as a semigroup. This defines a $\mu$-random walk on the countable state set $G$ as follows.
Let $\{s_{n}\in G:n\neq 1\}$ be the sequence of identical independent distributed (i.i.d.) random variables with law $\mu$ defined on $(G^{\mathbb{N}},\mu^{\mathbb{N}})$. Formally, we define $s_{n}$ to the $n$-th projection of $G^{\mathbb{N}}$ to $G$. The $\mu$-random walk $\omega_{n}=\prod_{i=1}^{n}s_{i}$ (starting at identity) is then product of i.i.d. random variables, recording the random position of the walker with i.i.d. increments $s_{i}$. The law of $\omega_{n}$ is given by the $n$-th convolution $\mu^{\star n}$.
We look at the trajectory of $\omega_{n}$ on the space $\mathrm{Y}$, the orbital points $\{\omega_{n}\cdot o\}$, and assume that the $\mu$-random walk has finite logarithmic moment with respect to the metric on $\mathrm{Y}$:
$\sum_{g\in G}\mu(g)\cdot|\log d(o,go)|<\infty.$
Let $F$ be a set of three independent non-pinched contracting elements in $G$. Let $\mathscr{F}$ as in (3) be the associated $G$-translated axes of elements in $F$. By [BBFS19, Theorem 5.10], the action of $G$ on the projection complex $\mathcal{P}_{K}(\mathscr{F})$ is acylindrical (see [GY22, Thm 6.3] for a detailed proof).
By the construction of projection complex, the orbital map
$go\in\mathrm{Y}\longmapsto g\bar{o}\in\mathcal{P}_{K}(\mathscr{F})$
is clearly a Lipschitz map. In particular, if a $\mu$-random walk on $\mathrm{Y}$ has finite logarithmic moment, then the induced random walk on $\mathcal{P}_{K}(\mathscr{F})$ has the finite logarithmic moment as well. [MT18, Theorem 1.5] then says that almost every trajectory on $\mathcal{P}_{K}(\mathscr{F})$ tends to a point in the Gromov boundary of $\mathcal{P}_{K}(\mathscr{F})$.
If $H(\mu)<\infty$, the Gromov boundary of $\mathcal{P}_{K}(\mathscr{F})$ endowed with hitting measure $\mu$ is the Poisson boundary for the $\mu$-random walk on $G$. Via Lemma 4.19, we obtain that almost every trajectory on $\mathrm{Y}$ tends to the locus of a conical point in $\partial_{\mathrm{H}}{\mathrm{Y}}$.
We vary $F$ over all such sets as above in $G$. By Corollary 4.17, the countable intersection of the embeddings of Gromov boundaries $\partial\mathcal{P}_{K}(\mathscr{F})$ gives the Myrberg set $\Lambda_{m}{(Go)}$. Hence, we conclude that almost every trajectory on the original space $\mathrm{Y}$ converges to the locus of a Myrberg point.
By Lemma 4.22, $[\Lambda_{m}{(Go)}]$ is a second countable metrizable space. As the hitting measure $\nu$ is supported on $\Lambda_{m}{(Go)}$, the partition []delimited-[][$\cdot$][ ⋅ ] thus forms a measurable partition in sense of Rokhlin: there are a countably many of measurable sets $B_{n}$ in $\partial{\mathrm{Y}}$, which are union of []delimited-[][$\cdot$][ ⋅ ]-classes, separating almost every two []delimited-[][$\cdot$][ ⋅ ]-classes ([Cou16, Def. 15.1]).
By Rokhlin’s theory, the measurable quotient space $([\Lambda_{m}{(Go)}],[\nu])$ by the above discussion is isomorphic to the Poisson boundary with hitting measure. The proof of Theorem 1.19 is therefore complete.
Remark 8.8.
Using the recent work [CFFT22] 111appearing after the first version of this paper, the finite logarithmic moment condition could be removed, still under the finite entropy. Indeed, there the authors proved that, without moment condition, the Gromov boundary is identified with Poisson boundary for the random walk for the acylindrical action on a hyperbolic space. The above argument works without changes modulo this ingredient.
9. Cogrowth of normal subgroups in groups of divergent type
From this section, we begin to present the applications of results established in previous sections. The first application is Theorem 1.17 from Introduction, which is a general result concerning the co-growth tightness of normal subgroups.
Theorem 9.1.
Assume that $G\curvearrowright\mathrm{Y}$ is a non-elementary group of divergent type with a contracting element. Then for any non-elementary normal subgroup $H$ of $G$, we have
$\omega_{H}>\frac{\omega_{G}}{2}.$
Moreover, if $H$ is of divergent type, then $\omega_{H}=\omega_{G}$.
The strict inequality “$\omega_{H}>\frac{\omega_{G}}{2}$” is sometimes referred to as cogrowth tightness in literature.
The remainder of this section is devoted to the proof, which are subdivided into the following two cases.
9.1. Normal subgroups are of divergent type
Assume that $H\curvearrowright\mathrm{Y}$ is of divergent type. We are going to prove that $\omega_{H}=\omega_{G}$. The proof presented here is inspired by the one in [MYJ20].
Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be a PS measure on the horofunction boundary $\partial_{\mathrm{H}}{\mathrm{Y}}$, which is an accumulation point of $\{\mu_{x,s}^{o}\}$ in (24) supported on $Ho$ for a sequence $s\searrow\omega_{H}$. This is a $\omega_{H}$-dimensional $H$-equivariant conformal density by Lemma 6.3. That is, $\lambda=1$ and $\epsilon=0$ in Definition 6.1 and Assump D.
Keep in mind that all the Patterson construction below uses the action $H\curvearrowright\mathrm{Y}$ instead of $G\curvearrowright\mathrm{Y}$, so gives (quasi)conformal density for $H$ (rather than for $G$).
The goal of proof is providing a constant $M>0$ such that
(41) $\forall x\in Go:\quad\|\mu_{x}\|\leq M$
Once this is proved, the shadow principal 6.11 implies that
$\mu_{1}([\Pi_{o}^{F}(go,r)])\asymp e^{-\omega_{H}d(o,go)}$
for any $go\in Go$. The same argument as in Proposition 6.8 proves $\omega_{H}\geq\omega_{G}$. Thus the remainder of the proof is to prove (41). For this purpose, for given $go$, we consider another PS-measure $\{\mu_{x}^{go}\}_{x\in\mathrm{Y}}$ on $\partial_{\mathrm{H}}{\mathrm{Y}}$ as an accumulation point of $\{\mu_{x,s}^{go}\}$ in (24) supported on $Hgo$ for a sequence $s\searrow\omega_{H}$.
As $H\curvearrowright\mathrm{Y}$ is of divergent type, the Patterson’s construction gives
$\forall x\in\mathrm{Y}:\quad\|\mu_{x}^{go}\|=\lim_{s\to\omega+}\frac{\mathcal{P}_{H}(s,x,go)}{\mathcal{P}_{H}(s,o,go)},\quad\|\mu_{x}^{o}\|=\lim_{s\to\omega+}\frac{\mathcal{P}_{H}(s,x,o)}{\mathcal{P}_{H}(s,o,o)},$
From the definition of Poincaré series (5), we have
$\forall s>0:\quad\frac{\mathcal{P}_{H}(s,go,go)}{\mathcal{P}_{H}(s,o,go)}=\Bigg{[}\frac{\mathcal{P}_{H}(s,go,o)}{\mathcal{P}_{H}(s,o,o)}\Bigg{]}^{-1}$
For given $go\in Go$, let us take the same sequence of $s$ (depending on $go$) so that $\{\mu_{x,s}^{o}\}_{x\in\mathrm{Y}}$ and $\{\mu_{x,s}^{go}\}_{x\in\mathrm{Y}}$ converge to $\{\mu_{x}^{o}\}_{x\in\mathrm{Y}}$ and $\{\mu_{x}^{go}\}_{x\in\mathrm{Y}}$ respectively, so
(42) $\|\mu_{go}^{go}\|=\|\mu_{go}^{o}\|^{-1}$
Push forward the limiting measures $\{\mu_{x}^{o}\}_{x\in\mathrm{Y}}$ and $\{\mu_{x}^{go}\}_{x\in\mathrm{Y}}$ to the quasi-conformal densities on the quotient $[\partial_{m}H]$, which we keep by the same notation. The constant $\lambda$ in Definition 6.1 is universal for any $\{\mu_{x}^{go}\}_{x\in\mathrm{Y}}$ with $go\in Go$, as the difference of two horofunctions in the same locus of a Myrberg point is universal by Lemma 5.3. Thus, Lemma 8.5 gives a constant $M=M(\lambda)$ independent of $go$ such that for any $go\in Go$:
$M^{-1}\|\mu_{go}^{go}\|\leq\|\mu_{go}^{o}\|\leq M\|\mu_{go}^{go}\|$
so the relation (42) implies $\|\mu_{go}^{o}\|^{2}\leq M$.
It still remains to note that $\{\mu_{x}\}$ and $\{\mu_{x}^{o}\}$ may be different limit points of $\{\mu_{x,s}^{o}\}_{x\in\mathrm{Y}}$, where $\{\mu_{x}\}$ is the PS measure fixed at the beginning of the proof. However, applying Lemma 8.5 again gives $\|\mu_{x}\|\asymp_{\lambda}\|\mu_{x}^{o}\|$ for any $x\in\mathrm{Y}$ and thus for $x=go$. As desired in (41), we obtain a upper bound on $\|\mu_{go}\|$ depending only on $\lambda,M$. The proof of the Case 1 is completed.
9.2. Normal subgroups are of convergent type
Assume that $H\curvearrowright\mathrm{Y}$ is of convergent type. It suffices to prove
$\omega(H)>\frac{\omega(G)}{2}.$
Indeed, if denote $A_{n}=A(o,n,\Delta)$ defined in (27), we have
$\mathcal{P}_{G}(\omega,o,o)\asymp_{\Delta}\sum_{n\geq 0}|A_{n}|e^{-n\omega}.$
A simple exercise shows that for any $\Gamma\subseteq G$, we have
(43) $\limsup_{n\to\infty}\frac{\log\sharp\left(A(o,n,\Delta)\cap\Gamma\right)}{n}=\omega_{\Gamma}$
Let $r,F>0$ given by Lemma 2.17 and $R=\|Fo\|+4r+4\Delta$. Let $B_{n}$ be a maximal $R$-separated subset of $A_{n}$:
  1. (1)
    For any $v\neq w\in B_{n}$ we have $d(v,w)>R$
  2. (2)
    For any $w\in A_{n}$ there exists $w\in B_{n}$ such that $d(v,w)\leq R.$
The idea of the proof is to produce at least $|B_{n}|$ elements in $H\cap A(o,{2n},R)$:
(44) $|B_{n}|\leq|H\cap A(o,{2n},R)|$
Indeed, this follows from [Yan19, Lemma 2.19]. As it is short, let us give the proof at the convenience of reader.
Apply Lemma 2.17 to the pair $(g,g^{-1})$ for each $go\in B_{n}$. There exists $f\in F$ such that the admissible path labeled by $gfg^{-1}$ $r$-fellow travels a geodesic $\gamma=[o,gfg^{-1}o]$ so $d(go,\gamma)\leq r$ and
$|d(o,gfg^{-1}o)-2d(o,go)|\leq\|Fo\|+4r.$
If $gfg^{-1}o=g^{\prime}f^{\prime}g^{\prime-1}o$ for two $go\neq g^{\prime}o\in B_{n}$, then $d(go,\gamma),d(g^{\prime}o,\gamma)\leq r$. As $go,g^{\prime}o\in A_{n}$ lies in the same annulus of width $\Delta$, we see that $d(go,g^{\prime}o)\leq 2r+2\Delta$. This contradicts to the item (1). Hence, (44) is proved and by (43),
$\omega_{H}\geq\frac{\omega_{G}}{2}.$
By the item (2), $|A_{n}|\asymp_{R}|B_{n}|$. If $\omega_{H}=\frac{\omega_{G}}{2}$, then
$\mathcal{P}_{G}(\omega_{G},o,o)\asymp_{R}\sum_{n\geq 1}|B_{n}|e^{-n\omega_{G}}\leq\sum\limits_{n\geq 1}|A(o,2n,R)\cap H|e^{-\omega_{H}n}<\infty$
where the last inequality follows from the convergence of $\mathcal{P}_{H}(\omega_{H},o,o)$ at $\omega_{H}$. This contradicts the divergent action of $G\curvearrowright\mathrm{Y}$. Hence, the proof of the second case is completed.
10. Relatively hyperbolic groups: Case Study (I)
This section is to illustrate some aspects of the convergence boundary through Floyd and Bowditch boundaries of relatively hyperbolic groups. The latter two boundaries were serving as a main motivation to the introduction of Assump A, Assump B and Assump C of a convergence boundary.
Most materials presented here are not really new to the experts in the field, though the discussion concerning the Myrberg sets in this setup seems not be found in literature. In this sense, a new result that we shall establish, Theorem 10.21, is that the Myrberg set in Floyd boundary is homeomorphic to the one in Bowditch boundary. Consequently, the Myrberg set is invariant among the peripheral structures endowed on a fixed relatively hyperbolic group. The same conclusion holds for Cannon–Thurston map (Theorem 10.23).
10.1. Convergence group actions
We start by introducing the notion of convergence group actions, which sets the stage for further discussion. We refer the reader to [Bow99, Tuk94a] for more details and relevant discussion.
A group action of $G$ by homeomorphisms on a compact metrizable space $M$ is called convergence group action if the diagonal action of $G$ on the triple space $\Theta^{3}(M)=\{(x,y,z)\in M^{3}:x\neq y\neq z\neq x\}$ of $M$ is properly discontinuous: for any compact set $K$ in $\Theta^{3}(M)$, the following set
$\sharp\{g\in G:gK\cap K\neq\emptyset\}<\infty$
The prototype example of a convergence action is the boundary action on the Gromov boundary induced from a proper action on the hyperbolic space, and the action on the Floyd boundary introduced in the next subsection.
By the general theory of convergence groups, the elements in $G$ are subdivided into the classes of elliptic, parabolic and hyperbolic elements. The latter two are infinite order elements having exactly one and two fixed points in $M$ accordingly. We denote by $h^{-},h^{+}$ the repelling and attracting fixed points of a hyperbolic element $h$.
Lemma 10.1.
Let $h\in G$ be a hyperbolic element with respect to $G\curvearrowright M$. Then
  1. (1)
    $\langle h\rangle$ acts properly and compactly outside $M\setminus h^{\pm}$. In particular,
    • it admits North–South dynamics as in Lemma 3.27 (with []delimited-[][$\cdot$][ ⋅ ] being singleton).
    • the set stabilizer of $h^{\pm}$ is the maximal elementary subgroup containing $h$.
  2. (2)
    The fixed points of two hyperbolic elements are either disjoint or equal.
  3. (3)
    If $h,k$ are two hyperbolic elements with disjoint fixed points, then $g_{n}=h^{n}k^{n}$ for $n\gg 0$ is hyperbolic, so that the fixed points $g_{n}^{+}$ and $g_{n}^{-}$ tend to $h^{+}$ and $k^{-}$ as $n\to\infty$.
Proof.
These are proved by Tukia [Tuk94a]: Theorem 2G, Theorem 2I and Lemma 2J. ∎
The limit set $\Lambda H$ of a subgroup $H<G$ is the set of accumulation points of all $H$-orbits in $M$. If the limit set $\Lambda H$ contains more 3 points, then it is the unique minimal $H$-invariant closed subset. The limit set contains the essential dynamical information, on which $H$ restricts as a convergence group action.
We hereafter assume that $\sharp\Lambda G\geq 3$ and $G$ admits a minimal convergence action on $M$ by restricting on its limit set. That is, $M=\Lambda G$ consists of limit points.
We introduce several classes of limit points, which shall be used later on.
Definition 10.2.
Let $\Lambda G\Join\Lambda G$ denote the set of unordered pairs of distinct points in $\Lambda G$.
  1. (1)
    A point $\xi\in\Lambda G$ is called conical if there exists a sequence of elements $g_{n}\in G$ ($n\geq 1$) such that the closure of $\{g_{n}(\xi,\eta):n\geq 1\}$ in $\Lambda G\Join\Lambda G$ is disjoint from the diagonal $\Delta(\Lambda G)=\{(x,x):x\in\Lambda G\}$ for any $\eta\in\Lambda G\setminus\xi$.
  2. (2)
    A point $\xi\in\Lambda G$ is called Myrberg point if for any $\eta\in\Lambda G\setminus\xi$, the set of $G$-orbits of $(\eta,\xi)$ is dense in the set of distinct pairs $\Lambda G\Join\Lambda G$.
  3. (3)
    A point $\xi\in\Lambda G$ is called bounded parabolic if the stabilizer $G_{\xi}$ of $\xi$ in $G$ is infinite, and acts properly and co-compactly on $\Lambda G\setminus\xi$. The subgroup $G_{\xi}$ is called maximal parabolic subgroup.
  4. (4)
    A convergence group action of $G$ on $M$ is called geometrically finite if every limit point $\xi\in\Lambda G$ is either a conical point or a bounded parabolic point.
Remark 10.3.
  1. (1)
    We demand the dense property in the space of distinct pairs only for symmetric reasons: if it is dense in $\Lambda G\Join\Lambda G$, it is so in the product $\Lambda G\times\Lambda G$.
  2. (2)
    According to definition, a Myrberg point must be a conical point, but the converse is false in general. By the South-North dynamics, the fixed points $h^{\pm}$ of a hyperbolic element $h\in G$ are conical. However, $h^{\pm}$ are not Myrberg if the two fixed points $\{h^{-},h^{+}\}$ of a hyperbolic element $h\in G$ forms a null sequence under the action of $G$: only finitely many pairs have diameter bigger than any fixed constant. For example, this property holds for the convergence action on Gromov boundary induced from a proper action on a hyperbolic space. See Lemma 4.15 for the similar results concerning about fixed points of contracting elements.
In practice, it suffices to approach the fixed points of all hyperbolic elements in the above definition of a Myrberg point.
Lemma 10.4.
A point $\xi\in M$ is a Myrberg point if and only if for any hyperbolic element $h\in G$ and for any $\eta\neq\xi$, there exists $g_{n}\in G$ so that $g_{n}(\eta,\xi)\to(h^{-},h^{+})$.
Proof.
As the limit set is $G$-minimal, the set of fixed points of hyperbolic element is dense in $\Lambda G$. Thus by Lemma 10.1.(3), the set of pairs of fixed points of all hyperbolic elements is dense in $\Lambda G\times\Lambda G$: if $h,k$ are hyperbolic element, then $g_{n}=h^{n}k^{n}$ for $n\gg 0$ is a hyperbolic element so that $g_{n}^{+}$ and $g_{n}^{-}$ tends to $h^{+}$ and $k^{-}$ respectively. ∎
In the next lemma, we could restrict to a smaller class of hyperbolic elements.
Lemma 10.5.
Suppose that $\Phi:G\curvearrowright M\to G\curvearrowright N$ is a $G$-equivariant map between two minimal convergence group actions. Then the following hold,
  1. (1)
    The preimage of a conical point $\xi\in N$ consists of one conical point.
  2. (2)
    Let $A$ be the set of all hyperbolic elements with respect to the action $G\curvearrowright N$. Then the pairs of fixed points of $h\in A$ are dense in $M\Join M$.
Proof.
We only prove the assertion (2) here; the assertion (1) is due to [Ger12, Prop. 7.5.2] and [MiOY13, Lemma 2.3]. By the classification of elements, a hyperbolic element in $G\curvearrowright N$ is also hyperbolic with respect to $G\curvearrowright M$, for an infinite order element cannot have three fixed points. Let $\Lambda\subseteq M$ be the union of fixed points of hyperbolic elements in $A$ with respect to the convergence action $G\curvearrowright M$.
Let $U,V$ be two open neighborhoods of $\xi$ and $\eta$ respectively in $M$. Note that $M$ is the unique minimal $G$-invariant set, so $\Lambda$ is dense in $M$. Let us choose $h,k\in A$ so that $h^{+}\in U$ and $k^{-}\in V$ are fixed points in $M$. By Lemma 10.1.(3), the element $g_{n}:=h^{n}k^{n}$ for $n\gg 0$ is hyperbolic with respect to $G\curvearrowright N$ (not being considered w.r.t. $G\curvearrowright M$), so $g_{n}$ is contained in $A$. At the same time, their two fixed points of $g_{n}$ in $M$ tend to $\xi$ and $\eta$ as $n\to\infty$. This establishes the assertion (2) and the proof is complete. ∎
10.2. Floyd boundary as convergence boundary
We introduce a compactification of any locally finite graph due to W. Floyd [Flo80]. The Cayley graph of a finitely generated group shall be our main focus. We follows closely the exposition in [Ger12], [GP13] and [Kar03].
Let $G$ be a group with a finite generating set $S$. Assume that $1\notin S$ and $S=S^{-1}$. Let $\mathscr{G}(G,S)$ denote the Cayley graph of $G$ with respect to $S$, equipped with the word metric $d$. We define a Floyd metric on $\mathscr{G}(G,S)$ by rescaling the word metric as follows.
Fix a Floyd scaling function $f:\mathbb{N}_{\geq 0}\to\mathbb{R}_{>0}$, satisfying for some $10<\lambda<$ 1, so that
  • $\sum_{n\geq 0}f(n)<\infty$
  • $\lambda f(n)\leq f(n+1)\leq f(n)$
The Floyd length $\ell_{o,f}(e)$ of an edge $e$ in $\mathscr{G}(G,S)$ is $\lambda^{n}$, where $n=d(o,e)$. The Floyd length $\ell_{o,f}(\gamma)$ of a path $\gamma$ is the sum of Floyd lengths of its edges. This induces a length metric $\rho_{o,f}$ on $\mathscr{G}(G,S)$, which is the infimum of Floyd lengths of all possible paths between two points.
Let $\overline{G}_{\mathrm{F}}$ be the Cauchy completion of $G$ with respect to $\rho_{o,f}$. The complement $\partial_{\mathrm{F}}{G}$ of $\mathscr{G}(G,S)$ in $\overline{G}_{\mathrm{F}}$ is called Floyd boundary of $G$. The $\partial_{\mathrm{F}}{G}$ is called non-trivial if $\sharp\partial_{\mathrm{F}}{G}>2$. The non-triviality of Floyd boundary does not depend on the choice of generating sets [Yan22, Lemma 7.1]. Most of groups do have trivial Floyd boundary: we refer to [KN04, Lev20] for relevant discussion. Currently, the most general class of groups known with non-trivial Floyd boundary are relatively hyperbolic groups, to be introduced in the next subsection.
In what follows, we shall omit the Floyd function $f$ in the notations $\rho_{o,f},\ell_{o,f}$ and write $\rho_{o},\ell_{o}$ for simplicity.
By construction, we have the following equivariant property
(45) $\rho_{o}(x,y)=\rho_{go}(gx,gy)$
(46) $\lambda^{d(o,o^{\prime})}\leq\frac{\rho_{o}(x,y)}{\rho_{o^{\prime}}(x,y)}\leq\lambda^{-d(o,o^{\prime})}$
for any two points $o,o^{\prime}\in G$. So for different basepoints, the corresponding Floyd compactifications are bi-Lipschitz. As a consequence of (45 and 46), one derives that the left-multiplication on $G$ extends to the boundary as bi-Lipschitz homeomorphism. Note that the topology may depend on the choice of the rescaling function and the generating set.
As fore-mentioned, the action on the Floyd boundary provides an important source of convergence group actions. If $\sharp\partial_{\mathrm{F}}{G}\geq 3$, Karlsson proved in [Kar03] that $\Gamma$ acts by homeomorphism on $\partial_{\mathrm{F}}{G}$ as a convergence group action. Moreover, the cardinality of $\partial_{\mathrm{F}}{G}$ is either 0, 1, 2 or uncountably infinite. By [Kar03, Proposition 7], $\partial_{\mathrm{F}}{G}=2$ exactly when the group $\Gamma$ is virtually infinite cyclic. These are based on the following fact in [Kar03], which is the main tool in understanding the Floyd geometry.
Lemma 10.6 (Visibility lemma).
There is a function $\varphi:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ such that for any $v\in G$ and any geodesic $\gamma$ in $\mathscr{G}(G,S)$, we have if $\ell_{v}(\gamma)\geq\kappa,$ then $d(v,\gamma)\leq\varphi(\kappa)$.
From this lemma, we could deduce that Floyd boundary is visual: any quasi-geodesic ray converges to a boundary point, and any two points $x,y\in G\cup\partial_{\mathrm{F}}{G}$ can be connected by a two-sided or one-sided infinite geodesic. See [GP13, Prop. 2.4] for a proof.
Let $f\in G$ be a hyperbolic element with two distinct fixed points $f^{-},f^{+}\in\partial_{\mathrm{F}}{G}$, so that the action on $\partial_{\mathrm{F}}{G}\setminus\{f^{\pm}\}$ admits South-North dynamics (cf. Lemma 10.1). If $E(f)$ denotes the set stabilizer of the two fixed points $f^{\pm}$, then it is the unique maximal elementary subgroup containing $f$. Moreover, by [Yan14a, Lemma 7.2] and [GP16, Prop 8.2.4], the quasi-axis $\textrm{Ax}(f)=E(f)$ is contracting in the sense of Definition 2.1. Compare these results with the ones in §3 using the sole contracting property.
The following easy consequence shall be used later on.
Lemma 10.7.
If $\gamma_{n}=[x_{n},y_{n}]$ is a sequence of geodesics so that $\|\textrm{Ax}(f)\cap N_{r}(\gamma_{n})\|$ tends to $\infty$ for a large fixed $r>0$, then $x_{n}\to f^{-}$ and $y_{n}\to f^{+}$ (up to swapping $f^{-}$ and $f^{+}$).
Proof.
Let $u_{n}\in\textrm{Ax}(f)$ be a shortest projection point of $x_{n}$ to $\textrm{Ax}(f)$. The $C$-contracting property of $\textrm{Ax}(f)$ implies that any geodesic $[x_{n},z]$ for given $z\in\textrm{Ax}(f)$ intersects the $C$-neighborhood of $u_{n}$. Let $r>C$. By assumption, $\|\textrm{Ax}(f)\cap N_{r}(\gamma_{n})\|\to\infty$, so we see that $d(1,x_{n})\to\infty$. Letting $z\to f^{-}$, we obtain that $d(1,[z,x_{n}])\to\infty$, so $x_{n}\to f^{-}$ follows from Lemma 10.6. The proof for $y_{n}\to f^{+}$ is completely analogous and left to the interested reader. ∎
Here is another easy corollary relating the Floyd geometry to the existence of a barrier along a geodesic in Definition 2.8.
Lemma 10.8.
There exists $\kappa>0$ depending on $\textrm{Ax}(f)$ with the following property for any $L\gg 0$. Assume that a geodesic $\gamma$ contains a segment $\alpha$ in $N_{r}(\textrm{Ax}(f))$ with length at least $L$. Let $x$ be the middle point of $\alpha$. Then $\rho_{x}(\gamma^{-},\gamma^{+})>\kappa$.
To conclude this subsection, we prove that Floyd boundary is a convergence boundary with maximal partition in §3. We first start with the following fact due to A. Karlsson.
Lemma 10.9.
There exists a continuous $G$-equivariant map from $\partial_{\mathrm{H}}G$ to $\partial_{\mathrm{F}}G$.
Proof.
We show that if $x_{n}\to\xi$ in the horofunction compactification, then $x_{n}$ converges to a point $\eta$ in the Floyd compactification. Arguing by contradiction, let us assume (after dropping and re-indexing) the sequences $x_{2n}\to\eta$ and $x_{2n+1}\to\zeta$ for $\eta\neq\zeta$ in the horofunction compactification. By [GP13, Prop. 2.4], there exists a geodesic between any two distinct $\xi$ and $\eta$. This contradicts to the Assump C of the horofunction boundary. Therefore, we obtain a well-defined map from $\partial_{\mathrm{H}}G$ to $\partial_{\mathrm{F}}G$. ∎
We now verify the following.
Lemma 10.10.
The nontrivial Floyd boundary of the Cayley graph of a finitely generated group is a convergence boundary with maximal partition, where conical points are non-pinched.
Proof.
Any geodesic ray converges to a limit point in $\partial_{\mathrm{F}}{G}$ by [GP13, Prop. 2.4]. By Lemma 10.9, the Assump B for horofunction boundary implies Assump B for Floyd boundary. These verify the Assump A and Assump B. It remains to show that a conical point is non-pinched. This is a consequence of the fact that there exists no horocycle at a conical point by [GP13, Lemma 3.6]. Recall that a horocycle at a boundary point $\xi$ means a bi-infinite geodesic with two half-rays tending to $\xi$.
Indeed, let $x_{n},y_{n}\in G$ tend to a conical point $\xi$. If $[x_{n},y_{n}]$ for all $n\geq 1$ intersects a fixed ball around the identity $1$, then we could extract a subsequence of $[x_{n},y_{n}]$ still denoted by $[x_{n},y_{n}]$ so that $[x_{n},y_{n}]$ converges locally uniformly to a bi-infinite geodesic $\gamma=[\gamma^{-},\gamma_{+}]$. We claim that $\gamma$ would be a horocycle at $\xi$; that is, $\gamma^{-}=\gamma_{+}=\xi$. This gives the desired contradiction to [GP13, Lemma 3.6]. To this end, consider the common overlap $[u_{n},v_{n}]=[x_{n},y_{n}]\cap\gamma$. The length of $[u_{n},v_{n}]$ tends to $\infty$, so $\gamma\setminus[u_{n},v_{n}]$ leaves every compact set, so $\rho_{1}(u_{n},\gamma^{-})\to 0$ and $\rho(v_{n},\gamma^{+})\to 0$ by Lemma 10.6. Hence, it suffices to show that $\rho_{1}(u_{n},x_{n})\to 0$ and $\rho_{1}(v_{n},y_{n})\to 0$. This again follows from Lemma 10.6, as $[u_{n},x_{n}]$ and $[v_{n},y_{n}]$ are escaping. Therefore, we conclude that $\gamma^{-}=\gamma_{+}=\xi$ and the proof of the lemma is complete. ∎
Recall that Floyd boundary surjects onto the end boundary for any finitely generated group, where the map is injective on the set of thin ends (cf. [GGPY21, Prop. 7.18]). We refer to [GGPY21] for the relevant discussion. We record the following corollary.
Corollary 10.11.
The end boundary of (the Cayley graph of) a finitely generated group is a convergence boundary, where the thin ends are non-pinched.
10.3. Homeomorphic Myrberg sets for relatively hyperbolic groups
There are various equivalent ways in literature ([Far98, Bow12, Osi06b, DS05, Ger09] etc) to introduce the relative hyperbolicity for a finitely generated group $G$ (even for countable groups which however is beyond the scope of this paper). We adopt here the point of view from the boundary theory. Recall Definition 10.2 for introduction of several classes of limit points.
Definition 10.12.
Let $\mathbb{P}$ be a finite collection of subgroups in $G$. A pair $(G,\mathcal{P})$ is relatively hyperbolic if $G$ admits a geometrically finite group action on a compact metrizable space $M$ such that $\{gPg^{-1}:P\in\mathbb{P},g\in G\}$ are exactly the collection of all maximal parabolic subgroups.
Using the relative Cayley graph, one can construct the limit set $\Lambda G$ of the action $G\curvearrowright M$ with the Gromov boundary of this graph [Bow12]. We will often call Bowditch boundary, denoted by $\partial_{\mathrm{B}}G$, the limit set $\Lambda G$ of a geometrically finite action. Bowditch proved that if $G$ is finitely generated then $\partial_{\mathrm{B}}G$ up to an equivariant homeomorphism depends only on the pair $(G,\mathcal{P})$ [Bow99].
Remark 10.13.
To be distinguished with the limit set on Bowditch boundary, we shall use $\partial H$ to denote the limit set of $H$ in the Floyd boundary $\partial_{\mathrm{F}}{G}$, which are the accumulation points of $H$ in $\partial_{\mathrm{F}}{G}$. Equivalently, it is the topological closure of the subset $H$ at the boundary in the compactification $\overline{G}_{\mathrm{F}}$.
Assume that $G$ is non-elementary: each maximal parabolic subgroup is of infinite index; equivalently, the action on the Bowditch boundary has at least three limit points.
The following result was first proved by Floyd [Flo80] for geometrically finite Kleinian group $\Gamma$, and then generalized to any relatively hyperbolic group by Gerasimov [Ger12].
Theorem 10.14.
[Ger12, Map Theorem] There exists a Floyd scaling function $f$ so that the identity on $G$ extends to a continuous and equivariant surjection $\phi$ from the Floyd compactification to the Bowditch compactification of $G$.
In [Ger12], Gerasimov proved the theorem for the scaling function $f(n)=\lambda^{n}$ for $\lambda\in(0,1)$, but any scaling function $f(n)>\lambda^{n}$ such as $f(n)=n^{p}$ for $p>1$ would also work. Indeed, the Floyd boundary with respect to $f(n)=n^{p}$ covers the one constructed using $f(n)=\lambda^{n}$.
10.4. Conical and Myrberg points
Let $(G,\mathcal{P})$ be a relatively hyperbolic pair. We say that an element $g\in G$ is hyperbolic if it is infinite order and is not conjugated into any $P$. Equivalently, it is a hyperbolic element on the Bowditch boundary in the sense of convergence boundary. Two hyperbolic elements are called independent if no power of them are conjugate.
The notion of a hyperbolic element plays a key role here in that it is contracting with respect to any word metric on $G$. The results obtained from the sole contracting property in previous subsections hold of course in this specific setup. In particular, a hyperbolic element $f$ is always contained in the unique maximal elementary subgroup $E(f)$ (in Lemma 2.6), which is equivalently the stabilizer in $G$ of the set of two fixed points $\{f^{-},f^{+}\}$ of $f$. In what follows, we however shall take the point of view from Floyd metric, which would be more powerful and convenient, to study the conical and Myrberg points.
Let $F$ be a finite set of independent hyperbolic elements. From a different perspective, Osin [Osi06a] proved that $G$ is hyperbolic relative to the union $\mathbb{P}\cup\mathbb{E}$, where $\mathbb{E}=\{E(f):f\in F\}$. Let $\mathcal{P}=\{gP:g\in G,P\in\mathbb{P}\}$. Note that the previously-defined collection $\mathscr{F}$ of $G$-translated axis of $f\in F$ are exactly the one of all left cosets of $E(f)$ for $f\in F$. It is well-known that $\mathcal{P}\cup\mathscr{F}$ forms a contracting system with bounded intersection (cf. §2.2).
Let us first recall the following characterization of conical points.
Lemma 10.15.
[Yan19, Lemma 2.14] A point $\xi$ in $\partial_{\mathrm{B}}G$ is conical if and only if for any geodesic ray $\gamma$ ending at $\xi$, there exists infinitely many elements $g_{n}\in G$ so that $d(g_{n},\gamma)\leq r$, where $r$ is a uniform constant independent of $\xi$ and $\gamma$.
Moreover, the direction $\Longrightarrow$” holds for a conical point $\xi$ in $\partial_{\mathrm{F}}{G}$ with the constant $r$ depending on $\xi$.
According to the definition 4.3, a point $\xi$ is $(F,r)$-conical for some large $r>0$, if it is contained in the partial shadows $\Pi_{1}^{F}(g_{n},r)$ for infinitely many $g_{n}\in G$. That is, there exists a geodesic ray $\gamma=[1,\xi]$ so that $d(g_{n},\gamma),d(g_{n}f,\gamma)\leq r$.
Lemma 10.16.
Let $\xi$ be an $(F,r)$-conical point in $\partial_{\mathrm{F}}{G}$. Then for $\|F\|\gg 0$, $\xi$ is a conical point in the sense of convergence boundary (cf. Definition 4.3).
Proof.
Let $\gamma$ be a geodesic ray ending at an $(F,r)$-conical point $\xi\in\partial_{\mathrm{F}}{G}$. We first show that the endpoint $\eta$ of $\gamma$ in $\partial_{\mathrm{B}}G$ is a conical point by using Lemma 10.15.
Indeed, according to definition, $\gamma$ contains infinitely many $(r,f)$-barriers $g_{n}$ for $f\in F$. That is, $\gamma\cap g_{n}N_{r}(\textrm{Ax}(f))$ has diameter at least $d(o,fo)$. By Lemma 10.8, if $\|F\|=\min\{d(o,fo):f\in F\}\gg 0$, $\gamma$ contains infinitely many points $x_{n}$ so that $\rho_{x_{n}}(\gamma^{-},\gamma^{+})\geq\kappa$, where $\kappa$ is a positive constant independent of $x_{n}$. According to Lemma 10.6, it is easy to show that any other geodesic $\beta$ ending at $\eta$ has to pass through a fixed neighborhood of all but finitely many $x_{n}$. Hence, $\eta$ is a conical point by Lemma 10.15, so the preimage $\xi$ is a conical point in $\partial_{\mathrm{F}}{G}$ by Lemma 10.5. The proof is complete. ∎
According to Definition 10.2, let us recast the notion of Myrberg points in the Floyd boundary. Let $\partial_{\mathrm{F}}{G}\Join\partial_{\mathrm{F}}{G}$ denote the set of unordered pairs of distinct points in $\partial_{\mathrm{F}}{G}$.
Definition 10.17.
A point $\xi\in\partial_{\mathrm{F}}{G}$ is called a Myrberg point if for any $x\in G$, the $G$-orbit of $(x,\xi)$ is dense in $\partial_{\mathrm{F}}{G}\Join\partial_{\mathrm{F}}{G}$. To be precise, given $\eta\neq\zeta\in\partial_{\mathrm{F}}{G}$, there exists a sequence of $g_{n}\in G$ so that $g_{n}x\to\eta$ and $g_{n}\xi\to\zeta$.
Remark 10.18.
According to definition of Floyd length, if $\sup_{n\geq 1}d(x_{n},y_{n})<\infty$ and $d(1,x_{n})\to\infty$, we see that $\rho_{1}(x_{n},y_{n})\to 0$. Hence, if $g_{n}x\to\xi$, then $g_{n}y\to\xi$ for any $y\in G$. Thus, we could assume $x=1$ in the definition for convenience. This also implies that the point $x$ can be assumed to any boundary point different from $\xi$.
Using the geometry of the Cayley graphs, we describe the Myrberg point in the Floyd boundary. Denote by $\partial_{m}G$ the set of Myrberg points in $\partial_{\mathrm{F}}{G}$.
Lemma 10.19.
Let $A$ be a set of hyperbolic elements in $G$, which are closed under conjugation, so that their fixed point pairs are dense in $\partial_{\mathrm{F}}{G}\Join\partial_{\mathrm{F}}{G}$. Then $\xi\in\partial_{\mathrm{F}}{G}$ is a Myrberg point if and only if the following property holds:
  • for any $f\in A$ and for any integer $n>0$, there exist infinitely many distinct $g\in G$ so that
    $d(g,[x,\xi]),d(gf^{n},[x,\xi])\leq r$
    where $r$ depends on the contracting constant of $\textrm{Ax}(f)$.
The same characterization holds for Myrberg points in Bowditch boundary, with $\partial_{\mathrm{F}}{G}$ replaced by $\partial_{\mathrm{B}}G$ in the above statement.
Proof.
First of all, by Corollary 4.17, it suffices to show that $\xi\in\partial_{\mathrm{F}}{G}$ satisfying the following property is a Myrberg point:
  • Given any $f\in A$ and any integer $n>0$, a geodesic ray $\gamma=[x,\xi]$ starting at $x\in G$ contains infinitely many $(r,f^{n})$-barrier, where $r>C$ is a large constant depending on the contraction constant of $\textrm{Ax}(f)$.
To this end, consider a pair of distinct points $(\zeta,\eta)$ in $\partial_{\mathrm{F}}{G}$. We shall find a sequence of $g_{n}\in G$ so that $[x,\xi]\to(\zeta,\eta)$. By assumption, let $h_{n}$ be a sequence of hyperbolic elements in $A$ so that $(h_{n}^{-},h_{n}^{+})$ tends to $(\zeta,\eta)$. According to the above property, for each fixed $f:=h_{n}$, we have that $\gamma$ contains infinitely many $(r,f^{n})$-barriers, where $r$ depends on the contracting constant of the axis of $f$. Let us choose one $(r,f^{n})$-barrier $g_{n}$ among them, so that the diamter of $g_{n}\gamma\cap N_{r}(\textrm{Ax}(f))$ tends to $\infty$. By Lemma 10.7, we obtain
$\rho_{o}(g_{n}x,f^{-}),\;\rho_{o}(g_{n}\gamma^{+},f^{+})\leq 1/n.$
Note that this holds for each $f:=h_{n}$, hence the assumption that $(h_{n}^{-},h_{n}^{+})$ tends to $(\zeta,\eta)$ shows that $g_{n}(x,\xi)$ tends to $(\zeta,\xi)$. This concludes the proof of the lemma. ∎
The same proof also gives the following description of Myrberg points in the limit set of discrete groups on hyperbolic spaces.
Lemma 10.20.
Assume that $G$ act properly on a hyperbolic space $\mathrm{Y}$. Let $A$ be a set of hyperbolic elements in $G$ closed under conjugation so that their fixed point pairs are dense in $\Lambda G\Join\Lambda G$. Then $\xi\in\Lambda G$ is a Myrberg point if and only if the following property holds:
  • for any $f\in A$ and for any integer $n>0$, there exist infinitely many distinct $g\in G$ so that
    $d(g,[x,\xi]),d(gf^{n},[x,\xi])\leq r$
    where $r$ depends on the contracting constant of $\textrm{Ax}(f)$.
Here $[x,\xi]$ denotes a geodesic in the hyperbolic space $\mathrm{Y}$.
10.5. Homeomorphisms between Myrberg sets
Let $\Phi:\partial_{\mathrm{F}}{G}\to\Lambda G$ be the map from the Floyd boundary to Bowditch boundary given by Theorem 10.14. By the continuity, a Myrberg point in $\partial_{\mathrm{F}}{G}$ is of course sent to a Myrberg point in $\Lambda G$. Let $\Phi_{m}:\partial_{m}G\to\Lambda G$ denotes the map of $\Phi$ restricted onto the Myrberg set $\partial_{m}G$.
Theorem 10.21.
The map $\Phi_{m}:\partial_{m}G\to\Lambda_{m}G$ is a homeomorphism onto the set of Myrberg points in $\Lambda G$.
Proof.
Let $\xi\in\Lambda G$ be a Myrberg point. It is a conical point, so the preimage under $\Phi$ is a singleton consisting of a conical point in $\partial_{\mathrm{F}}{G}$ by Lemma 10.5. Hence, it follows that $\Phi_{m}$ must be injective. We are now going to prove the surjectivity of the map.
Let $A$ be the set of hyperbolic elements relative to the convergence action of $G$ on the Bowditch boundary. The pairs of fixed points in $\Lambda G$ of the elements in $A$ is dense in $\Lambda G\Join\Lambda G$, and moreover, by Lemma 10.5, their preimages under $\Phi$ (i.e. fixed points in $\partial_{\mathrm{F}}{G}$) are dense in $\partial_{\mathrm{F}}{G}\Join\partial_{\mathrm{F}}{G}$.
As $\xi\in\Lambda_{m}G$ is Myrberg point, for any fixed $n$ and hyperbolic element $f$ in $A$, the geodesic ray $[1,\xi)$ contains infinitely many segments labeled by $f^{n}$ in its $r$-neighborhood for some $r$ depending on $\textrm{Ax}(f)$. This property involves the geometry of the Cayley graph, without using the boundary, so by Lemma 10.19, the preimage $\Phi^{-1}(\xi)$ is also a Myrberg point in the Floyd boundary. Therefore, we proved that $\Phi_{m}$ is a continuous and bijective map to $\Lambda_{m}G$.
At last, we prove that the inverse of $\Phi$ is continuous on $\Lambda_{m}G$. To this end, let $\xi_{n}\to\xi\in\Lambda_{m}G$. Let $\alpha=[1,\xi]$ and $\alpha_{n}=[1,\xi_{n}]$. If $\xi_{n}\to\xi$, then by Lemma 10.20, $\alpha_{n}$ eventually contains infinitely copies of $g_{n}f^{m}$ in its $r$-neighborhood. This proves $\xi_{n}\to\xi\in\partial_{m}G$. The proof is complete. ∎
We now record the main conclusion from the above result, saying that Myrberg set is invariant independent of the peripheral structure of $G$.
Corollary 10.22.
Let $(G,\mathbb{P})$ and $(G,\mathbb{P}^{\prime})$ be two relative hyperbolic pairs for the same group $G$. Then there exists a homeomorphism between their Myrberg sets in $\partial_{\mathrm{B}}(G,\mathbb{P})$ and $\partial_{\mathrm{B}}(G,\mathbb{P}^{\prime})$, which are homeomorphic to the Myrberg set in the Floyd boundary $\partial_{\mathrm{F}}(G)$.
To conclude this subsection, let us prove that the same conclusion holds for the Cannon–Thurston map.
Assume that $G$ is a hyperbolic group, which acts properly on a hyperbolic space $\mathrm{Y}$. Assume that the orbit map
$g\in G\longmapsto go\in\mathrm{Y}$
extends continuously (and uniquely) to a map from the Gromov boundary $\partial G$ of $G$ to the limit set of $G$. Such a boundary map $\Phi$ is referred to as Cannon–Thurston map in literature ([CT07, Mj14]).
Theorem 10.23.
The map $\Phi:\partial G\to\Lambda G$ restricts to a homeomorphism from $\partial_{m}G$ to the Myrberg set in $\Lambda G$.
Proof.
The map $\Phi$ is a $G$-equivariant map from $\partial G\to\Lambda G$. By Lemma 10.5, the preimage of any conical point in $\Lambda G$ is just singleton of one (conical) point in $\partial G$.
Let $A$ be the set of hyperbolic elements in $G$ acting on the hyperbolic space $\mathrm{Y}$. By Lemma 10.5, $A$ satisfies the assumption of Lemma 10.20. The proof follows closely the one for Lemma 10.21. We leave it to the interested reader. ∎
11. CAT(0) groups with rank-1 elements: Case Study (II)
The groups acting properly on CAT(0) spaces with rank-1 elements are another important class of groups motivating the study of the paper. Applications are presented with an emphasis to the Myrberg set and the contracting boundary for CAT(0) spaces.
We assume the reader is familiar with the basics on CAT(0) geometry and refer the reader to [BH99] for the reference.
11.1. Myrberg set in the visual boundary
We first prove that the visual boundary of a proper CAT(0) space is a convergence boundary. The proper assumption is only used to guarantee the compactification.
Lemma 11.1.
Let $\mathrm{Y}$ be a proper CAT(0) space. Then the visual boundary $\partial_{\mathrm{Vis}}\mathrm{Y}$ is a convergence boundary with respect to the maximal partition, so that all boundary points are non-pinched.
Proof.
It is well-known that the visual boundary is homeomorphic to horofunction boundary ([BH99]). By Theorem 1.1, it suffices to prove that this finite difference partition is trivial: every point is minimal.
Indeed, arguing by contradiction, let $\xi\neq\eta\in\partial_{\mathrm{Vis}}\mathrm{Y}$ be represented respectively by two geodesic rays $\alpha,\beta$ from $o\in\mathrm{Y}$, so that $\|b_{\xi}-b_{\eta}\|_{\infty}\leq K$ for some $K>0$. On one hand, for any $x\in\alpha$, we have $b_{\xi}(x)=d(o,x)$ so $|b_{\eta}(x)-d(o,x)|\leq K$. Taking a sequence $y_{n}\in\beta$ so that $b_{y_{n}}(x)=d(x,y_{n})-d(o,y_{n})$ tends to $b_{\xi}(x)$, we have
$|d(o,x)+d(o,y_{n})-d(x,y_{n})|\leq K+1$
for all $n\gg 0$. On the other hand, the Alexandrov angle $\theta:=\angle(\alpha,\beta)$ is positive for $\alpha\neq\beta$ (cf. [BH99]), so the Euclidean comparison triangle for $\Delta(o,x,y_{n})$ has the Euclidean angle at $o$ bigger than $\theta>0$. Thus the Euclidean geometry tells us that $|d(o,x)+d(o,y_{n})-d(x,y_{n})|$ must tend to $\infty$ as $d(o,y_{n})\to\infty$. This is a contradiction. The lemma is thus proved. ∎
The following result is crucial in the further discussion, roughly saying that the set of geodesics with positive flat strip is an open set.
Lemma 11.2.
[Bal95, Lemma III.3.1] Let $\gamma:\mathbb{R}\to\mathrm{Y}$ be a geodesic in a proper CAT(0) space $\mathrm{Y}$ without bounding a flat strip of width $R>0$. Then there exist neighborhoods $U$ of $\gamma^{+}$ and $V$ of $\gamma^{-}$ in $\overline{\mathrm{Y}}$ such that any $\xi\in U$ and any $\eta\in V$, there exists a geodesic from $\xi$ to $\eta$ within $R$-distance to $\gamma(0)$ and any such geodesic does not bound a $(2R)$-flat strip.
We now give the main result of this subsection, recovering [Lin18, Lemma 7.1] via completely different methods (also compare with [Ric17, Theorem 2]). Such results are crucial in applying Hopf’s argument in proving ergodicity of geodesic flows (see [Lin18, Section 7] for details).
Lemma 11.3.
Assume that $G\curvearrowright\mathrm{Y}$ in (1) contains a rank-1 element with zero width axis. Then
  1. (1)
    any bi-infinite geodesic $\gamma$ with one endpoint at a Myrberg point has zero width.
  2. (2)
    any two geodesic rays $\beta,\gamma$ ending at a Myrberg point are asymptotic: there exists $a\in\mathbb{R}$ such that
    $d(\beta(t),\gamma(t+a))\to 0,\quad t\to\infty.$
Proof.
Let $h$ be a rank-1 element with a geodesic axis $\alpha$ with zero width: for any $R>0$ there is no flat strip of width $R$ with $\alpha$ as one boundary component. Fix $o\in\alpha$.
(1). We are going to prove that $\gamma$ has zero width. Suppose to the contrary that $\gamma$ bounds an flat strip of width $R$ for some $R>0$. By Corollary 4.17, $\gamma$ contains infinitely many $(r,f_{n})$-barriers $g_{n}o$, where $f_{n}\in\langle h\rangle$ is any sequence of elements with unbounded length. Thus,
(47) $\|N_{r}(g_{n}\alpha)\cap\gamma\|\to\infty$
Thus, there exists a segment $p_{n}$ of $g_{n}\alpha$ such that $p_{n}\subseteq N_{r}(\gamma)$ and $\ell(p_{n})\to+\infty$.
For this $R/2$, let $U,V$ be their open neighborhoods of $\alpha^{-},\alpha^{+}$ satisfying Lemma 11.2. As $\ell(p_{n})\to+\infty$, the cone topology on the visual boundary of $\mathrm{Y}$ implies that $g_{n}^{-1}\gamma^{-}\in U,g_{n}^{-1}\gamma^{+}\in V$ for $n\gg 0$. By Lemma 11.2, the geodesic $g_{n}^{-1}\gamma$ does not bound an flat strip of width $R$. This contradicts the hypothesis, so we proved that $\gamma$ has zero width.
(2). As the CAT(0) distance has convexity property, it suffices to find two unbounded sequences of points $x_{n}\in\beta,y_{n}\in\gamma$ such that $d(x_{n},y_{n})\to 0$ as $n\to\infty$. By the same argument as above, we have (47) holds for $\gamma$ and $\beta$, so for any $R>0$, Lemma 11.2 implies that $\gamma,\beta$ are $R/2$-close to $g_{n}\alpha(0)$ for any $n\gg 0$. Thus we obtain $x_{n}\in\beta,y_{n}\in\gamma$ so that $d(x_{n},y_{n})\leq R$. Letting $R\to 0$ completes the proof. ∎
11.2. Roller boundary of CAT(0) cube complexes
Let $\mathrm{Y}$ be a proper CAT(0) cube complex, that is, a proper CAT(0) space built out of unit Euclidean $n$-cubes $[0,1]^{n}$ for $n\geq 1$ by side isometric gluing. The maximal dimension of $n$-cubes is called the dimension of $\mathrm{Y}$. In the sequel, we only consider finite dimensional CAT(0) cube complexes.
The collection $\mathcal{H}$ of hyperplanes or called walls as connected components of mid-cubes (i.e. copies of $[0,1]^{n-1}\times\{1/2\}$ in $n$-cubes) endows $\mathrm{Y}$ with a wall structure: every wall separates the space into two connected components called half-spaces. We say that an edge is duel to a wall if the edge is parallel to the mid-cube in the wall. An orientation of walls picks up exactly one half-space for each wall, and is called consistent if any two such chosen half-spaces intersect non-trivially.
For every vertex $x$ of $\mathrm{Y}^{0}$, the set $U_{x}$ of half-spaces containing $x$ forms a consistent orientation denoted by $\mathbf{x}\in 2^{\mathcal{H}}$. We equip the space $2^{\mathcal{H}}$ of all orientations with the compact product topology, where all consistent orientations form a closed subset. The Roller compactification $\overline{\mathrm{Y}}$ of $\mathrm{Y}$ is thus the closure of the vertex set $\mathrm{Y}^{0}$ (identified with just described consistent orientations) in the space of all consistent orientations on walls, and Roller boundary is $\partial_{\mathrm{R}}\mathrm{Y}:=\overline{\mathrm{Y}}\setminus\mathrm{Y}^{0}$.
The CAT(0) metric induces an $\ell^{2}$-metric on the 1-skeleton $\mathrm{Y}^{1}$. In this subsection, we are mainly interested in the $\ell^{1}$-metric on $\mathrm{Y}^{1}$:
$\forall x,y\in\mathrm{Y}^{0}:\;d(x,y):=\frac{1}{2}|U_{x}\Delta U_{y}|$
where $\Delta$ denotes the symmetric difference of sets. If $\mathrm{Y}$ is finite dimensional, the $\ell^{1}$-metric and $\ell^{2}$-metric are quasi-isometric. The $\ell^{1}$-metric is exactly the combinatorial metric on $\mathrm{Y}^{1}$ and usually non-unique geodesic metric. However, the set of walls duel to the edges on any geodesic is more canonical and does not depend on the geodesic, which is exactly the set of walls separating $x$ and $y$. By abuse of language, we also say that the wall separating $x$ and $y$ is duel to a geodesic $[x,y]$.
Recall that an action of $G$ on the CAT(0) cubical complex $\mathrm{Y}$ is called essential if no $G$-orbit is contained in a finite neighborhood of some half-space. If $G$ has no fixed points in the visual boundary of $\mathrm{Y}$ or has only finitely many orbits of walls, then the action could be made essential by passing to its (convex) essential core (see [CS11, Prop 3.5]).
Assume that $\mathrm{Y}$ is irreducible: it cannot be written as a nontrivial metric product of two CAT(0) cube complexes. If $G$ acts essentially on $\mathrm{Y}$ without fixed points in the visual boundary, this is equivalent to one of the following
  1. (1)
    existence of strongly separated of two half spaces: no wall transverses both of them ([CS11, Prop. 5.1]).
  2. (2)
    existence of a contracting isometry on $\mathrm{Y}$ with respect to CAT(0) metric ([CS11, Theorem 6.3]).
Lemma 11.4.
[GY22, Lemma 8.3] Let $G\curvearrowright\mathrm{Y}$ be an essential proper action on an irreducible CAT(0) cube complex. Then $G$ contains a contracting element with respect to the action on 1-skeleton of $\mathrm{Y}$ with the $\ell^{1}$-metric. Moreover,
  • A contracting element in CAT(0) metric is contracting in the $\ell^{1}$-metric, and vice versa.
  • A contracting element in $\ell^{1}$-metric preserves an $\ell^{1}$-geodesic by translation.
Following [Fer18], a boundary point (i.e. a consistent orientation of walls) is called regular if the orientation of walls contains an infinite sequence of a descending chain of pairwise strongly separated half-spaces. We equip $\partial_{\mathrm{R}}\mathrm{Y}$ with the finite symmetric difference partition []delimited-[][$\cdot$][ ⋅ ]: two (consistent) orientations are equivalent if their symmetric difference is finite. It is proved in [FLM18, Lemma 5.2] that the regular points are []delimited-[][$\cdot$][ ⋅ ]-minimal: the []delimited-[][$\cdot$][ ⋅ ]-class is a singleton. The fixed points of contracting isometry are regular and thus minimal.
An unpublished result of Bader–Guralnik says that the Roller boundary is homeomorphic to the horofunction boundary of $(\mathrm{Y}^{0},d)$ (see a proof [FLM18, Prop. 6.20]). Here is the main result relating the previous study of convergence boundary to Roller boundary.
Lemma 11.5.
The Roller boundary $\partial_{\mathrm{R}}\mathrm{Y}$ is a convergence boundary, with finite symmetric difference partition []delimited-[][$\cdot$][ ⋅ ] so that []delimited-[][$\cdot$][ ⋅ ] coincides with the partition of finite difference of horofunctions.
Moreover, assume that a non-elementary group $G$ acts essentially on $\mathrm{Y}$. Then the following hold.
  1. (1)
    Every contracting element satisfies the North–South dynamics with minimal fixed points.
  2. (2)
    $G$ admits a unique invariant closed subset, where the fixed point pairs of contracting elements are dense in the distinct pairs.
  3. (3)
    The Myrberg points of $G$ are regular, and in particular, are minimal.
Proof.
By [Gen20, Prop A.2], every point in $\partial_{\mathrm{R}}\mathrm{Y}$ is represented by a geodesic ray (called Buseman point), via the identification with the half-spaces into which the ray eventually enters. We first prove that if two horofunctions $b_{\xi},b_{\eta}\in\partial_{\mathrm{R}}\mathrm{Y}$ have finite difference at most $K$, then $\xi$ and $\eta$ are in the same symmetric difference []delimited-[][$\cdot$][ ⋅ ]-class.
Fix a basepoint $o\in\mathrm{Y}$. Let $\alpha,\beta$ be two (combinatorial) geodesic rays issuing from $o$ and ending at $\xi,\eta$. Recall that $\partial_{\mathrm{R}}\mathrm{Y}$ is the horofunction boundary with respect to the combinatorial metric $d$. We obtain similarly the same estimates on Gromov product as in the proof of Lemma 11.1:
$12\langle o,y_{n}\rangle_{x}=d(o,x)+d(o,y_{n})-d(x,y_{n})\leq K+$ 1
for any $x\in\alpha,y_{n}\in\beta$ with $n\gg 0$. As half-spaces are combinatorially convex, the geodesic rays $\alpha$ and $\beta$ share the same set of half-spaces containing the common initial point $o$. Thus, it suffices to consider the difference on the half-spaces whose bounding walls are duel to edges of $\alpha$ or $\beta$.
$o$$x$$y_{n}$$\alpha$$\beta$$m$$\xi$$\eta$$(K+1)/2\geq$
Figure 16. The walls due to $[m,x]$
Recall that the median $m$ of $(o,x,y_{n})$ is the unique common point on all three sides of some geodesic triangle with these vertices. Alternatively, $m$ is the intersection of all geodesic segments between pairs of points in $(o,x,y_{n})$. From the above inequality, we obtain
$\langle o,y_{n}\rangle_{x}=d(x,m)\leq(K+1)/2.$
Replacing the initial subpaths of $\alpha$ and $\beta$ with appropriate geodesics with the same endpoints, we may assume that the median point $m$ lies both on $\alpha$ and $\beta$ so that $[o,m]_{\alpha}=[o,m]_{\beta}$. Thus, the walls intersecting $[o,x]_{\alpha}$ but not $[o,m]_{\alpha}=[o,m]_{\beta}$ are due to the geodesic $[x,m]$, which are at most $d(x,m)\leq(K+1)/2$ walls. Note that $m$ depends on $x$. See Fig. 16.
Letting $x\to\xi$ and choosing $m$ accordingly, a limiting argument shows that the symmetric difference of the walls intersecting $\alpha$ and the ones intersecting $\beta$ has at most $(K+1)/2$ half-spaces. This shows that $\xi,\eta$ have finite difference on walls.
Conversely, if $\xi,\eta$ has finite symmetric difference on $K$ walls, the similar reasoning as above produces two unbounded sequence of points $x_{n}\in\alpha,m_{n}\in\beta$ with $d(x_{n},m_{n})\leq K/2$. Thus, $\|b_{\xi}-b_{\eta}\|_{\infty}\leq K/2$.
To complete the proof, it remains to show that a Myrberg point is regular: it contains an infinite sequence of a descending chain of pairwise strongly separated half-spaces. Once this is proved, the statements (1) and (2) are immediate consequences of Lemmas 3.27 and 3.30.
Let $\gamma$ be a geodesic ray ending at a Myrberg point $\xi$. By [Gen20, Theorem 3.9], for any contracting isometry $h$, the combinatorial geodesic axis is duel to a pair of strongly separated half-spaces with arbitrarily large distance. As each wall separates $\mathrm{Y}$ into two convex components, for any two pairs $(x,y)$ and $(z,w)$ with $d(x,z),d(y,w)\leq r$, the walls due to $[x,y]$ differ by at most $2r$ from the ones to $[z,w]$. By Corollary 4.17, for any large $L>0$, $\gamma$ contains a sequence of disjoint sub-segments of length $L$ within the $r$-neighborhood of a translated contracting axis of $h$. Hence, $\gamma$ has to be duel to infinitely many pairs of strongly separated walls. By strong separateness, the half-spaces bounded by those walls and into which $\gamma$ eventually enters form a descending chain, so $\xi$ is a regular point by definition. Thus the lemma is proved. ∎
11.3. Contracting boundary in conformal measure
We start by proving a general result for any convergence boundary $\partial{\mathrm{Y}}$. Recall that a point $\xi$ in $\partial{\mathrm{Y}}$ is called visual if there exists a geodesic ray $\gamma$ from any point in $\mathrm{Y}$ ending at $[\xi]$. If in addition, $\gamma$ is contracting, then we call $\xi\in\partial{\mathrm{Y}}$ a contracting point.
Theorem 11.6.
Suppose that $G\curvearrowright\mathrm{Y}$ as in (1) is a co-compact action. Let $\{\mu_{x}\}_{x\in\mathrm{Y}}$ be a $\omega_{G}$-dimensional quasi-conformal density on a convergence boundary $\partial{\mathrm{Y}}$. Then the set of contracting points in $\partial{\mathrm{Y}}$ is $\mu_{1}$-null if and only if $G$ is not a hyperbolic group.
Fix a basepoint $o\in\mathrm{Y}$. As the action $G\curvearrowright\mathrm{Y}$ is co-compact, let $M>0$ such that $N_{M}(Go)=\mathrm{Y}$. In the proof, we need the following result which is of independent interest.
Let $\Lambda\subseteq\partial{\mathrm{Y}}$ be a set of visual points. Denote by $\textrm{Hull}_{R}(\Lambda)$ the set of orbital points $v\in Go$ such that $v$ is $R$-closed to a geodesic from $o$ to $\xi\in\Lambda$ for $R\geq 0$.
Lemma 11.7.
Suppose $\Lambda\subseteq\partial{\mathrm{Y}}$ has positive $\mu_{1}$-measure. Then for any $R\gg M$, there exists $c=c(\Lambda,R)>0$ such that
$|\textrm{Hull}_{R}(\Lambda)\cap N(o,n)|\geq c\cdot e^{\omega_{G}n},$
for any $n>0$.
Proof.
Denote $A_{n}=\textrm{Hull}_{R}(\Lambda)\cap A(o,n,\Delta)$. Set $R>\max\{r_{0},M\}$, where $r_{0}>0$ is given by Lemma 6.4. By definition, we have
$\Lambda\subseteq\bigcup_{v\in A_{n}}[\Pi_{o}^{F}(v,R)]$
which yields
$\mu_{1}(\Lambda)\leq\sum_{h\in A_{n}}\mu_{1}([\Pi_{o}(v,R)]).$
A co-compact action must be of divergent type, so by Theorem 1.14, $\mu_{1}$ is supported on Myrberg set. By Lemma 5.3, any two horofunctions in the locus of a Myrberg point have uniformly bounded difference. The proof for the upper bound in Lemma 6.4 gives
$\mu_{1}([\Pi_{o}(v,R)])\prec_{R}e^{\omega_{G}d(o,v)}$
Hence, there exists a constant $c>0$ such that
$|A_{n}|\geq c\cdot e^{\omega_{G}n}$
completing the proof. ∎
We are now ready to give the proof.
Proof of Theorem 11.6.
If $G$ is a hyperbolic group, then the co-compact action implies that $\mathrm{Y}$ is hyperbolic and the action $G\curvearrowright\mathrm{Y}$ is of divergent type. By Lemma 4.7, any $(r,F)$-conical point is visual and thus contracting, because every geodesic ray in a hyperbolic space is contracting. By Theorem 1.14, the set of contracting points is $\mu_{1}$-full, hence the $\Longrightarrow$” direction follows.
Let us prove the other direction by assuming that $G$ is not hyperbolic. Let $\Lambda$ denote the set of contracting points, and assume $\mu(\Lambda)>0$ for a contradiction.
Let $\Lambda_{C}(x)$ be the set of $\xi\in\Lambda$ so that there exists a $C$-contracting geodesic ray at $\xi$ starting at $x\in\mathrm{Y}$. Then $\Lambda$ admits the following countable union:
$\Lambda=\bigcup_{C\in\mathbb{N}}\Bigg{(}\bigcup_{x\in\mathrm{Y}}\Lambda_{C}(x)\Bigg{)}$
where $\Lambda(gx)=g\Lambda(x)$ for any $g\in G,x\in\mathrm{Y}$. Since the conformal measures are almost $G$-equivariant and the $M$-neighborhood of a $G$-orbit $Go$ covers $\mathrm{Y}$, there exists $C>0$ such that $\mu_{1}(\Lambda_{C}(o))>0$ for some (or any) $o\in\mathrm{Y}$.
Let $U=\textrm{Hull}_{R}(\Lambda_{C}(o))$ be the hull of $\Lambda_{C}(o)$ for some $R\gg 0$. As $\mu_{1}(\Lambda_{C}(o))>0$, the set $U\subseteq Go$ has the growth rate $\omega_{G}$ by Lemma 11.7.
By definition, any segment $[o,v]$ for $v\in U$ has two endpoints $R$-close to a $C$-contracting geodesic ray. Recall that by [Yan20, Prop. 2.2], there exists a constant $D=D(C,R)$ such that any geodesic segment with two endpoints in an $2R$-neighborhood of a $C$-contracting geodesic is $D$-contracting. Thus, $[o,v]$ for $v\in U$ is $D$-contracting.
It is well-known that if every geodesic segment in $\mathrm{Y}$ is uniformly contracting then $\mathrm{Y}$ is Gromov-hyperbolic. Thus there exists a segment $\alpha$, so that $\alpha$ is not $D$-contracting. As $G$ acts co-compactly on $\mathrm{Y}$, we may assume without loss of generality that $\alpha=[o,ho]$ for a basepoint $o\in\mathrm{Y}$ and $h\in G$ is not $D$-contracting.
Observe that any $[o,v]$ for $v\in U$ cannot contain $\alpha$ in its $R$-neighborhood. Indeed, if not, $\alpha$ lies in $2R$-neighborhood of a $C$-contracting geodesic ray, so must be $D$-contracting. That is to say, any $[o,v]$ for $v\in U$ does not contain any $(R,h)$-barrier in a sense of [Yan19]. By [Yan19, Theorem C], the set $U$ is growth tight: $\omega_{U}<\omega_{G}$. This contradicts the equality obtained as above. Hence, it is proved that $\mu_{1}(\Lambda)=0$. ∎
12. Conformal density for mapping class groups: Case Study (III)
Let $\Sigma_{g}$ be an orientable closed surface of genus $g\geq 2$. Consider the mapping class group
$\textrm{Mod}(\Sigma_{g})=\pi_{0}(\textrm{Homeo}^{+}(\Sigma_{g}))$
i.e.: isotopy classes of orientation preserving homeomorphisms of $\Sigma_{g}$.
Uniformization theorem for surfaces sets up one-to-one correspondence between complex structures and hyperbolic metrics on $\Sigma_{g}$. The Teichmüller space, denoted by $\mathcal{T}_{g}$, of $\Sigma_{g}$ is the set of isotopy classes of hyperbolic metrics or of complex structures on $\Sigma_{g}$. We equip $\mathcal{T}_{g}$ with Teichmüller metric $d_{T}$, for which the isometry group is exactly the mapping class groups $\textrm{Mod}(\Sigma_{g})$. It is well-known that pseudo-Anosov mapping classes are contracting in Teichmüller metric [Min96], so the results in the previous subsections apply. In the sequel, we refer to [FM12, FLP79] for more details on the undefined notions.
12.1. Preliminaries on Thurston and Gardiner–Masur boundaries
Let $\mathscr{S}$ be the set of (isotopy classes of) essential simple closed curves on $\Sigma_{g}$. Given $x\in\mathcal{T}_{g}$, each element $c\in\mathscr{S}$ can be assigned either the length $\ell_{x}(c)$ of a closed geodesic representative on $\Sigma_{g}$ equipped with a constant curvature $-1$ of Riemannian metric or the extremal length $\mathrm{Ext}_{x}(c)$ on $\Sigma_{g}$ with a Riemann surface structure.
Denote by $\mathbb{R}_{>0}^{\mathscr{S}}$ the set of positive functions on $\mathscr{S}$. We then obtain the embedding of $\mathrm{Y}$ into $\mathbb{R}_{>0}^{\mathscr{S}}$ in two ways
(48) $x\in\mathcal{T}_{g}\mapsto\Big{(}\ell_{x}(c)\Big{)}_{c\in\mathscr{S}}\;\textrm{ and }\quad x\in\mathcal{T}_{g}\mapsto\Big{(}\sqrt{\mathrm{Ext}_{x}(c)}\Big{)}_{c\in\mathscr{S}}$
whose topological closures in the projectived space $\mathbb{P}\mathbb{R}_{>0}^{\mathscr{S}}$ give the corresponding Thurston and Gardiner–Masur compactifications of $\mathcal{T}_{g}$ denoted by $\mathcal{T}_{g}\cup\partial_{\mathrm{Th}}\mathcal{T}_{g}$ and $\mathcal{T}_{g}\cup\partial_{\mathrm{GM}}\mathcal{T}_{g}$ ([FLP79, GM91]). We first describe the Thurston boundary $\partial_{\mathrm{Th}}\mathcal{T}_{g}$ in more details.
Thurston boundary
Let $\mathscr{MF}$ be the set of measured foliations on $\Sigma_{g}$ with a natural topology, so that $\mathscr{MF}$ admits a topological embedding into $\mathbb{R}_{>0}^{\mathscr{S}}$, by taking the intersection number $I(\xi,c)$ of a measured foliation $\xi$ with a curve $c$ in $\mathscr{S}$. As the weighted multi-curves form a dense subset in $\mathscr{MF}$, their geometric intersection number extends continuously to a bi-linear intersection form
$I(\cdot,\cdot):\quad\mathscr{MF}\times\mathscr{MF}\to\mathbb{R}_{\geq 0}.$
This allows to understand an essential simple closed curves $c\in\mathcal{S}$ as a measured foliation, where the transversal measure is given by the geometric intersection number.
Let $\mathscr{MF}\to\mathscr{PMF}$ be the projection to the set $\mathscr{PMF}$ of projective measured foliations by positive reals scaling. Thurston proved that $\mathscr{PMF}$ is homeomorphic to the sphere of dimension $76g-$ 7, and compactifies $\mathcal{T}_{g}$ in the projective space $\mathbb{P}\mathbb{R}_{>0}^{\mathscr{S}}$ as a closed unit ball ([FLP79]). Then $\partial_{\mathrm{Th}}\mathcal{T}_{g}:=\mathscr{PMF}$ is the so-called Thurston boundary.
By abuse of language, recalling $I(\cdot,\cdot)$ is bi-linear, we shall write $I(\xi,\eta)=0$ for two projective measured foliations $\xi,\eta\in\mathscr{PMF}$ to mean $I(\check{\xi},\check{\eta})=0$ for any of their lifts $\check{\xi},\check{\eta}\in\mathscr{MF}$.
Gardiner–Masur boundary
In [GM91], Gardiner–Masur proved that the Gardiner–Masur boundary $\partial_{\mathrm{GM}}\mathcal{T}_{g}$ from the embedding (48) contains $\mathscr{PMF}$ as a proper subset. In [Miy13, Corollary 1], Miyachi proved that the identification between the Teichmüller space $\mathcal{T}_{g}$ extends continuously
(49) $\mathcal{T}_{g}\cup\mathscr{UE}\subseteq\mathcal{T}_{g}\cup\partial_{\mathrm{Th}}\mathcal{T}_{g}\quad\longmapsto\quad\mathcal{T}_{g}\cup\mathscr{UE}\subseteq\mathcal{T}_{g}\cup\partial_{\mathrm{GM}}\mathcal{T}_{g}$
to a homeomorphism between the subsets $\mathscr{UE}$ in both boundaries. Furthermore, the following result will be crucially used later on.
Lemma 12.1.
[Miy14b, Corollary 1] For any $o\in\mathcal{T}_{g}$, there exists a unique continuous extension of the Gromov product
$(x,y)\in\mathcal{T}_{g}\times\mathcal{T}_{g}\quad\longmapsto\quad\langle x,y\rangle_{o}=\frac{d(x,o)+d(y,o)-d(x,y)}{2}\in\mathbb{R}_{\geq 0}$
to the Gardiner–Masur boundary $\mathscr{PMF}\subseteq\partial_{\mathrm{GM}}\mathcal{T}_{g}$ so that for any $\xi,\eta\in\mathscr{PMF}$
(50) $e^{-2\langle\xi,\eta\rangle_{o}}=\frac{I(\check{\xi},\check{\eta})}{\mathrm{Ext}_{o}(\check{\xi})^{1/2}\mathrm{Ext}_{o}(\check{\eta})^{1/2}}$
where $\check{\xi},\check{\eta}$ are any lifts of $\xi,\eta$ in $\mathscr{MF}$.
Remark 12.2.
(extension in Thurston boundary) Thanks to (49), the equality (50) holds for uniquely ergodic points $\xi,\eta\in\partial_{\mathrm{Th}}\mathcal{T}_{g}$. We now explain it also holds for $\xi,\eta\in\mathcal{S}\subseteq\mathscr{PMF}=\partial_{\mathrm{Th}}\mathcal{T}_{g}$. By Hubbard-Masur theorem, for any basepoint $o\in\mathcal{T}_{g}$ and for any $c\in\mathcal{S}\subseteq\mathscr{PMF}$, there exists a quadratic differential $q$ on the Riemann surface $o$ whose vertical foliation $q^{+}$ realizes $c$. Recall that the Teichmüller geodesic ray issuing from $o$ directed by $q^{+}$ has the limit point $c$ in either Thurston or Gardiner–Masur boundary (cf. [Mas82b, Theorem], [Miy13, Theorem 2]). Let $x_{n},y_{n}$ be on the geodesic rays ending at $\xi\in\partial_{\mathrm{Th}}\mathcal{T}_{g}$ and $\eta\in\partial_{\mathrm{Th}}\mathcal{T}_{g}$ so that $x_{n}\to\xi$ and $y_{n}\to\eta$ in both compactifications. Thus, (50) holds for $\xi,\eta\in\mathcal{S}\subseteq\partial_{\mathrm{Th}}\mathcal{T}_{g}$ as the right side of (50) involves only the Teichmüller metric without caring about compactification. H. Miyachi also indicated to the author that this also follows from the discussion [Miy13, Sec. 5.3].
12.2. Kaimanovich–Masur partition on Thurston boundary
Following Kaimanovich–Masur [KM96], we shall define a partition on $\mathscr{PMF}$ as follows. Let $\mathscr{MIN}$ be the subset of projective minimal foliations $\xi\in\mathscr{PMF}$ so that $I(\xi,c)>0$ for any $c\in\mathscr{S}$. The []delimited-[][$\cdot$][ ⋅ ]-class of $\xi\in\mathscr{MIN}$ defined as
$[\xi]\;:=\;\{\eta\in\mathscr{MIN}:\;I(\xi,\eta)=0\}$
forms a partition of $\mathscr{MIN}$, as $I(\xi,\eta)=0$ for $\xi\in\mathscr{MIN}$ implies $\eta\in\mathscr{MIN}$ ([Ree81]). The []delimited-[][$\cdot$][ ⋅ ]-class of minimal foliations $\xi\in\mathscr{MIN}$ being singleton forms exactly the set $\mathscr{UE}$ of uniquely ergodic foliations. The remaining non-minimal foliations $\mathscr{MF}\setminus\mathscr{MIN}$ are then partitioned into countably many classes
$[\xi]\;:=\;\{\eta:\;\forall c\in\mathscr{S},\;I(\eta,c)=0\Longleftrightarrow I(\xi,c)=0\}$
according to whether they share the same disjoint set of curves in $\mathscr{S}$. It is proved in [KM96, Lemma 1.1.2] that the partition []delimited-[][$\cdot$][ ⋅ ], generated by a countable partition, is measurable in the sense of Rokhlin.
The following useful fact is proved in [KM96, Lemma 1.4.2] where $\xi$ lies in $\mathscr{UE}$ and the conclusion follows as $y_{n}\to\xi$. We provide some explanation on the changes.
Lemma 12.3.
Assume that $x_{n}\in\mathcal{T}_{g}$ tends to $\xi\in\mathscr{MIN}$. If a sequence of points $y_{n}\in\mathcal{T}_{g}$ satisfies
$d(o,y_{n})-d(x_{n},y_{n})\to+\infty$
then $y_{n}\to[\xi]$.
Proof.
In the proof [KM96, Lemma 1.4.1], a sequence of $\beta_{n}\in\mathscr{S}$ is produced such that $\mathrm{Ext}_{x_{n}}(\beta_{n})$ is uniformly bounded and
$I(\beta_{n},F_{n})\to 0$
where $F_{n}$ is the vertical measured foliation of the terminal quadratic differential of the Teichmüller map from $x_{0}$ to $x_{n}$. Taking convergence in $\mathscr{PMF}$ and passing to a subsequence, assume that
$\beta_{n}\stackrel{{\scriptstyle\mathscr{PMF}}}{{\Longrightarrow}}F,\quad F_{n}\stackrel{{\scriptstyle\mathscr{PMF}}}{{\Longrightarrow}}F_{\infty}$
for some $F,F_{\infty}\in\mathscr{MF}$. Then $I(F,F_{\infty})=0=I(F_{\infty},\xi)$ by [Kar05, Lemma 43]. As all the foliations in $[\xi]$ for $\xi\in\mathscr{MIN}$ is minimal ([Ree81]), we obtain $F\in[\xi]$.
The remaining proof from page 245 of [KM96, Lemma 1.4.1] shows that $y_{n}$ tends to $F\in\mathscr{PMF}$. This is what we wanted. ∎
The Thurston boundary with the above partition []delimited-[][$\cdot$][ ⋅ ] satisfies the following stronger version of Assump B, which also holds for CAT(0) and hyperbolic spaces (see Lemma 5.4). To be precise, the cones in the following definition are contained in the cones in Assump B, where $\|[o,y]\cap N_{C}(X_{n})\|\gg 0$ is required.
Lemma 12.4.
Let $X_{n}$ be an escaping sequence of contracting subsets in $\mathcal{T}_{g}$. Denote the cone $Y_{n}=\Omega_{o}(X_{n})=\{y\in\mathcal{T}_{g}:[o,y]\cap X_{n}\neq\emptyset\}$. Then there exist a subsequence of $Y_{n}$ and a boundary point $\xi$ so that the following holds. Any convergent sequence $y_{n}\in Y_{n}$ tends into $[\xi]$.
Proof.
Indeed, after passage to subsequence, let $X_{n}$ be an escaping sequence of $C$-contracting subsets so that $x_{n}\in\pi_{X_{n}}(o)$ tends to $\xi\in\mathscr{PMF}$. Take any sequence of points $y_{n}\in\Omega_{o}(X_{n})$ converging to some $\eta\in\mathscr{PMF}$ by compactness. Recall that $\Omega_{o}(X_{n})$ is the set of points $y\in\mathcal{T}_{g}$ with $[o,y]\cap X_{n}\neq\emptyset.$ The contracting property in Lemma 2.2 implies $d(\pi_{X_{n}}(o),[o,y_{n}])\leq C$, so we have
(51) $|d(o,y_{n})-d(o,x_{n})-d(x_{n},y_{n}|\leq 2C$
If $\xi\in\mathscr{MIN}$, then $\eta\in[\xi]$ follows by Lemma 12.3 as $d(o,x_{n})\to\infty$.
To finish the proof of Assump B, it thus remains to consider the case $\xi\in\mathscr{PMF}\setminus\mathscr{MIN}$, and prove $[\xi]=[\eta]$: they has the same set of disjoint curves in $\mathscr{S}$. That is to say, $I(\xi,c)=0$ for $c\in\mathscr{S}$ is equivalent to $I(\eta,c)=0$. The two directions are symmetric. We only prove that if $I(\xi,c)=0$ then $I(\eta,c)=0$.
Let $z_{n}$ be be a sequence of points on a geodesic ray ending at $c\in\mathscr{PMF}=\partial_{\mathrm{Th}}\mathcal{T}_{g}$, so $z_{n}\to c$ also takes place in $\partial_{\mathrm{GM}}\mathcal{T}_{g}$. By Lemma 12.1, as $I(\xi,c)=0$, we have $\langle x_{n},z_{n}\rangle_{o}\to\infty$. By definition of Gromov product,
[…]
so by Lemma 12.1 with the Remark after it, we have $I(\eta,c)=0$. ∎
We are now ready to prove the main result, Theorem 1.2, of this subsection.
Proposition 12.5.
The Thurston boundary $\partial_{\mathrm{Th}}\mathcal{T}_{g}$ with the above Kaimanovich–Masur partition []delimited-[][$\cdot$][ ⋅ ] is a convergence boundary of Teichmüller space $\mathcal{T}_{g}$, so that the uniquely ergodic points $\mathscr{UE}$ are non-pinched points.
Proof.
It is well-known that a contracting geodesic ray $X$ tends to a unique ergodic point (see [Cor17]). Any sequence $y_{n}$ with an escaping $x_{n}\in\pi_{X}(y_{n})$ satisfies the condition of Lemma 12.3, so Assump A follows. The non-pinched points in Assump C contain $\mathscr{UE}$, which follows by Lemma 12.1 and (49).
Assump B follows from Lemma 12.4. The proof is thus complete. ∎
In [MP89], McCarthy–Papadopoulos defined a notion of the limit set, denoted by $\Lambda_{\mathcal{MP}}(H)$, in Thurston boundary for a non-elementary subgroup $H<\textrm{Mod}(\Sigma_{g})$ with a pseudo-Anosov element, as the closure of fixed points of pseudo-Anosov elements. They further showed that the set $\Lambda(Ho)$ of accumulation points of a $H$-orbit in $\mathcal{T}_{g}$ is contained in the intersection locus of $\Lambda$ (those measured foliations with zero intersection with the ones in $\Lambda$). Their limit set $\Lambda_{\mathcal{MP}}(H)$ is exactly the set $\Lambda$ in Lemma 3.30, so we have $[\Lambda]=[\Lambda Go]$. Here []delimited-[][$\cdot$][ ⋅ ] denotes the KM partition, but not the intersection locus in [MP89]. Therefore, it would be interesting to examine the exact relation between the Kaimanovich–Masur partition []delimited-[][$\cdot$][ ⋅ ] and the intersection locus.
In [LS14], Liu–Su proved that Gardiner–Masur boundary is the horofunction boundary of Teichmüller space (with Teichmüller metric). We list a few consequences of the general theory, in view of the corresponding results on Thurston boundary [MP89].
Lemma 12.6.
$\partial_{\mathrm{GM}}\mathcal{T}_{g}$ is a convergence boundary with respect to finite difference relation []delimited-[][$\cdot$][ ⋅ ]. Moreover,
  1. (1)
    The finite difference relation []delimited-[][$\cdot$][ ⋅ ] restricts on $\mathscr{UE}\subseteq\partial_{\mathrm{GM}}\mathcal{T}_{g}$ as a maximal partition.
  2. (2)
    All $(r,F)$-conical points in Definition 4.3 are uniquely ergodic points.
  3. (3)
    Every pseudo-Anosov elements are non-pinched contracting elements with minimal fixed points and have the North–South dynamics on $\partial_{\mathrm{GM}}\mathcal{T}_{g}$.
  4. (4)
    Every non-elementary subgroup $G<\textrm{Mod}(\Sigma_{g})$ has a unique minimal $G$-invariant closed subset $\Lambda$, so that $[\Lambda(Go)]=[\Lambda]$ for any $o\in\mathcal{T}_{g}$.
Proof.
The assertions (3) and (4) are corollaries of Lemmas 3.27 and 3.30, provided that (1) and (2) are proved.
(1). Let $[\xi]=[\eta]$ with $\xi\in\mathscr{UE}$, so $\|b_{\xi}-b_{\eta}\|\leq K<\infty$. Take a geodesic ray $\gamma=[o,\xi]$ and $x_{n}\in\gamma$ tends to $\xi$. Let $y_{n}\to\eta$. As $|b_{\xi}(x_{n})-b_{y_{n}}(x_{n})|=|d(o,x_{n})+d(o,y_{n})-d(x_{n},y_{n})|\leq K$, then $y_{n}\to\xi$ by Lemma 12.3, so $\xi=\eta$ is proved.
(2). Projecting to the quotient $\mathcal{T}_{g}/{\textrm{Mod}(\Sigma_{g})}$, any geodesic ray ending at an $(r,F)$-conical point $\xi\in\partial_{\mathrm{GM}}\mathcal{T}_{g}$ is recurrent into the $r$-neighborhood of closed geodesics corresponding to pseudo-Anosov elements in $F$. Thus, through (49), $\xi$ is uniquely ergodic by a criterion of Masur [Mas80]. ∎
12.3. Conformal density on Thurston and Gardiner–Masur boundaries
We first introduce the (formal) conformal density defined in [ABEM12] on $\partial_{\mathrm{Th}}\mathcal{T}_{g}$, and then clarify the relation with the conformal density obtained from the Patterson’s construction.
The train track parametrization endows the set $\mathscr{MF}$ of measured foliations with a Piecewise Linear structure. Let $\mu_{\mathrm{Th}}$ be the resulted $\textrm{Mod}(\Sigma_{g})$-invariant Thurston measure on $\mathscr{MF}$, in the Lebesgue measure class, pasting from the Lebesgue measure on the train track charts on $\mathscr{MF}$. By Masur’s ergodicity theorem, $\mu_{\mathrm{Th}}$ is a $\textrm{Mod}(\Sigma_{g})$-invariant ergodic measure, which is thus unique up to scaling.
According to [ABEM12], a family of conformal measures are constructed via $\mu_{\mathrm{Th}}$ as follows. Let $U$ be a Borel subset in $\mathscr{PMF}$. Define for any $x\in\mathcal{T}_{g}$,
$\lambda_{x}(U)\;:=\;\mu_{\mathrm{Th}}(\{\xi\in\mathscr{MF}:\check{\xi}\in U,\;\mathrm{Ext}_{x}(\xi)\leq 1\}).$
where $\check{\xi}$ denotes the image of $\xi$ in $\mathscr{PMF}$.
By definition, the family of $\textrm{Mod}(\Sigma_{g})$-equivariant measures $\{\lambda_{x}\}_{x\in\mathcal{T}_{g}}$ are mutually absolutely continuous, and are related as follows
$\mu_{y}-\textrm{a.e. }\xi\in\mathscr{PMF}:\quad\frac{d\lambda_{x}}{d\lambda_{y}}(\xi)=e^{-\omega_{G}B_{\xi}(x,y)}$
where $\omega_{G}=6g-6$ and the cocycle $B_{\xi}:\mathcal{T}_{g}\times\mathcal{T}_{g}\to\mathbb{R}$ is is given by
(52) $\forall x,y\in\mathcal{T}_{g},\quad B_{\xi}(x,y):=\frac{1}{2}\cdot\log\frac{{\mathrm{Ext}_{x}(\hat{\xi})}}{{\mathrm{Ext}_{y}(\hat{\xi})}}.$
where $\hat{\xi}\in\mathscr{MF}$ is any lift of $\xi$. By [Wal19, Theorem 1], one can verify that for $\xi\in\mathscr{UE}$, $B_{\xi}(\cdot,\cdot)$ is indeed the Buseman cocycle at $\xi$ in the horofunction boundary $\partial_{\mathrm{GM}}\mathcal{T}_{g}$ (see [TY21] for details). By (49), the Buseman cocycles extend continuously to $\mathscr{UE}\subseteq\partial_{\mathrm{Th}}\mathcal{T}_{g}$, so Assump D is satisfied with $\epsilon=0$.
It is well-known that $\lambda_{x}$ is supported on $\mathscr{UE}$ ([Mas82a, Vee82]). In terms of Definition 6.1, $\{\lambda_{x}\}_{x\in\mathcal{T}_{g}}$ is a $(6g-6)$-dimensional $\textrm{Mod}(\Sigma_{g})$-equivariant conformal density on $\partial_{\mathrm{Th}}\mathcal{T}_{g}$. The shadow lemma 6.4 and principle 6.11 thus hold in this context, where the former has been already obtained in [TY21] via a different method.
We emphasize the above construction is not obtained via the Patterson’s construction. However, the action $\textrm{Mod}(\Sigma_{g})\curvearrowright\mathcal{T}_{g}$ has purely exponential growth ([ABEM12, Thm 1.2]), hence it is of divergent type. By Theorem 1.16, there exists a unique $(6g-6)$-dimensional conformal density on $\partial_{\mathrm{Th}}\mathcal{T}_{g}$, which was first proved in [Ham09a, LM08]. Thus, the conformal density obtained by Patterson’s construction coincides with $\lambda_{x}$.
We record the following result for further counting applications.
Lemma 12.7.
Let $G<\textrm{Mod}(\Sigma_{g})$ be a non-elementary subgroup with pseudo-Anosove elements. Assume that the limit set $\Lambda G$ of $G$ is a proper subset of $\mathscr{PMF}$. Then its Poincaré series $\mathcal{P}_{G}(s,o,o)$ is convergent at $66g-$ 6.
In particular, the Thurston measure of the conical limit set $\Lambda_{c}(G)$ defined in (6) is zero.
Remark 12.8.
Analogous to the Ahlfors area theorem in Kleinian groups, it would be interesting to study which subgroups of $\textrm{Mod}(\Sigma_{g})$ have limit set of zero Lebesgue measure. If $G$ is a convex-cocompact subgroup, then the conical limit set $\Lambda_{c}(G)$ is the whole limit set, so we recover [IL07, Corollary 4.8]. The limit set of geometrically finite Kleinian groups consists of conical points and countable many parabolic points. If the Lebesgue measure has no charge on conical limit points, then the whole limit set is null. We expect the “in particular” statement helpful to solve the zero measure problem for certain limit sets.
Proof.
A non-elementary subgroup contains at least two independent pseudo-Anosov elements, so is sufficiently large in terms of [MP89]. With respect to Kaimanovich–Masur []delimited-[][$\cdot$][ ⋅ ]-partition, any pseudo-Anosov element $h$ has minimal fixed points $[h^{-}]=\{\xi\}$ for some $\xi\in\mathscr{UE}$. By [MP89], $G$ possesses a unique $G$-invariant minimal closed subset $\Lambda\subseteq\mathscr{PMF}$ defined as the closure of all fixed points of pseudo-Anosov elements (see also Lemma 3.30). Let $Z(\Lambda):=\{\eta\in\mathscr{PMF}:I(\xi,\eta)=0\text{ for some }\xi\in\Lambda\}$. Then $\Lambda(Go)\subseteq Z(\Lambda)$ is proved by [MP89, Prop 8.1].
By [MP89, Prop 6.1, Thm 6.16], if $\Lambda\neq\mathscr{PMF}$, then $Z(\Lambda)\neq\mathscr{PMF}$ is closed and $G$ acts properly discontinuously on the complement $\Omega:=\mathscr{PMF}\setminus Z(\Lambda)$. By Lemma 3.15, the fixed point pairs of all pseudo-Anosov elements are dense in distinct pairs of $\mathscr{PMF}$. Choose a pseudo-Anosov element $h\in\textrm{Mod}(\Sigma_{g})$ and a closed neighborhood $U$ of $h^{\pm}$ such that $U$ is contained in $\Omega$. By the proper action of $G\curvearrowright\Omega$, the set $A:=\{1\neq g\in G:U\cap gU\neq\emptyset\}$ is finite. As $G<\textrm{Mod}(\Sigma_{g})$ is residually finite, let us choose a finite index subgroup $\dot{G}<G$ such that $\dot{G}\cap A=\emptyset.$ Thus, any nontrivial element $g\in\dot{G}$ sends $U$ into $V:=\mathscr{PMF}\setminus U$.
By the Nouth-South dynamics, there exists a high power of $h$ (with the same notation) such that $h^{n}(U)\subseteq V$ for any $n\neq 0$. Thus, $\langle h\rangle$ and $\dot{G}$ are ping-pong players on $(U,V)$, so generate a free product $\langle h\rangle\star\dot{G}$. By [Yan19, Lemma 2.23], the Poincaré series of $\dot{G}$ converges at $\omega_{\textrm{Mod}(\Sigma_{g})}=6g-6$. If $G\curvearrowright\mathcal{T}_{g}$ is of divergent type, we have $\omega_{G}=\omega_{\dot{G}}<6g-6.$
By Theorem 1.22, we have the Thurston measure of the conical limit set of $G$ is null. ∎
For further reference, we record the following corollary of Theorem 1.16 for Gardiner–Masur boundary.
Lemma 12.9.
Up to scaling, the Gardiner–Masur boundary $\partial_{\mathrm{GM}}\mathcal{T}_{g}$ admits a unique and ergodic $\textrm{Mod}(\Sigma_{g})$-equivariant conformal density of dimension $66g-$ 6, which is supported on the Myrberg set (a subset of uniquely ergodic points).
12.4. Bowen–Margulis–Sullivan measure on Teichmüller geodesic flow
Let $G$ be a non-elementary subgroup of $\mathrm{Mod}(\Sigma_{g})$ with pseudo-Anosov elements, so that $\mathcal{M}:=\mathcal{T}_{g}/G$ is a (branched) cover of the moduli space of $\Sigma_{g}$. Let $\{\mu_{x}\}_{x\in\mathcal{T}_{g}}$ be a $\omega_{G}$-dimensional conformal density on $\partial_{\mathrm{Th}}\mathcal{T}_{g}=\mathscr{PMF}$ in Theorem 1.22.
By Theorem 1.14, it suffices to prove the direction “(3) $\Longleftrightarrow$ (4)” of Theorem 1.22. This requires to run the well-known Hopf’s argument which almost follows verbatim the case for CAT(-1) spaces [Hop71, Nic89, BM96]. What follows sets up the necessary background, and then points out the required ingredients.
Let $\pi:\mathcal{Q}\mathcal{T}_{g}\to\mathcal{T}_{g}$ be the vector bundle of quadratic differentials over Teichmüller space. We first explain a $G$-invariant geodesic flow on the sub-bundle $\mathcal{Q}^{1}\mathcal{T}_{g}$ of unit area quadratic differentials, and then construct a flow invariant measure called Bowen–Margulis–Sullivan measure. This construction is due to Sullivan in hyperbolic spaces.
Teichmüller geodesics $\gamma$ are uniquely determined by a pair of projective classes $(\check{\xi},\check{\eta})$ of transversal measured foliations $(\xi,\eta)$, and vice versa. Such pairs $(\xi,\eta)$ form exactly the complement of the following big diagonal
$\Delta\;:=\;\{(\xi,\eta)\in\mathscr{MF}\times\mathscr{MF}:\quad\exists c\in\mathscr{S},\;I(c,\xi)+I(c,\eta)=0\}$
According to Hubbard-Masur theorem [HM79], the transversal pair $(\xi,\eta)$ is realized as the corresponding vertical and horizontal foliations $(q^{+},q^{-})$ of a unique quadratic differential $q\in\mathcal{Q}\mathcal{T}_{g}$. According to [LM08], we consider the following $G$-equivariant homeomorphism
$\begin{array}[]{rl}\mathcal{Q}\mathcal{T}_{g}&\longrightarrow\quad\mathscr{MF}\times\mathscr{MF}\setminus\Delta\\ q&\longmapsto\quad(q^{+},q^{-}).\end{array}$
Remark 12.10.
By abuse of language, we write $\gamma=(\check{\xi},\check{\eta})=(q^{+},q^{-})$ and $\gamma^{+}=\check{\xi},\gamma^{-}=\check{\eta}$. Indeed, the two half rays of $\gamma$ do not necessarily converge to $\check{\xi},\check{\eta}$ in $\mathcal{T}_{g}\cup\partial_{\mathrm{Th}}\mathcal{T}_{g}$, which however do converge for $\xi,\eta\in\mathscr{UE}$.
The geodesic flow $(\mathcal{G}^{t})_{t\in\mathbb{R}}$ on $\mathcal{Q}^{1}\mathcal{T}_{g}$ is defined as
$\mathcal{G}^{t}(q):=(e^{t}q^{+},e^{-t}q^{-})$
for any $q\in\mathcal{Q}^{1}\mathcal{T}_{g}$ and $t\in\mathbb{R}$. The flow line $t\mapsto q_{t}:=\mathcal{G}^{t}(q)$ gives the lift of the Teichmüller geodesic $\gamma=(q_{0}^{+},q_{0}^{-})$ in the moduli space $\mathcal{Q}\mathcal{T}_{g}$ of quadratic differentials. Endow $\mathcal{Q}^{1}\mathcal{T}_{g}$ with a $\textrm{Mod}(\Sigma_{g})$-invariant distance $d_{Q}$ as follows:
[…]
which descends to the quotient metric still denoted by $d_{Q}$ on $\mathcal{Q}^{1}\mathcal{M}_{g}:=\mathcal{Q}^{1}\mathcal{T}_{g}/G$.
Hopf parameterization of Teichmüller flows
Set
$\mathcal{T}_{g}\Join\mathcal{T}_{g}:=\Big{(}(\mathscr{PMF}\times\mathscr{PMF})\setminus\Delta\Big{)}.$
Fix a basepoint $o\in\mathcal{T}_{g}$, the Hopf parameterization of $\mathcal{Q}^{1}\mathcal{T}_{g}$ is given by the following $G$-equivariant homeomorphism
$\begin{array}[]{ccc}\mathcal{Q}^{1}\mathcal{T}_{g}&\longrightarrow&\mathcal{T}_{g}\Join\mathcal{T}_{g}\times\mathbb{R}\\ q&\longmapsto&(\xi,\eta,B_{\xi}(o,\pi(q)))\end{array}$
where $B_{\xi}(\cdot,\cdot)$ is the cocycle in (52). The $G$-action on the target is given by
$\forall g\in G:\;g(q^{-},q^{+},s)=(gq^{-},gq^{+},s+B_{\xi}(g^{-1}o,o)).$
As a consequence, the geodesic flow $(\mathcal{G}^{t})_{t\in\mathbb{R}}$ on $\mathcal{Q}^{1}\mathcal{T}_{g}$ is conjugate to the additive action on the time factor of $\mathcal{T}_{g}\Join\mathcal{T}_{g}\times\mathbb{R}$ as follows:
$(\xi,\eta,s)\longmapsto(\xi,\eta,t+s).$
Following [ABEM12, Section 2.3.1], define
$\beta(x,\xi,\eta):=\big{[}e^{B_{\xi}(x,p)+B_{\eta}(x,p)}\big{]}^{6g-6}={\left[\frac{\sqrt{\mathrm{Ext}_{x}(\hat{\xi})\mathrm{Ext}_{x}(\hat{\eta})}}{I(\hat{\xi},\hat{\eta})}\right]}^{6g-6}$
for any $p\in[\xi,\eta]$, where $\hat{\xi},\hat{\eta}\in\mathscr{MF}$ are lifts of $\xi,\eta$, and $\mathrm{Ext}_{p}(\check{\xi})\mathrm{Ext}_{p}(\check{\eta})=I(\check{\xi},\check{\eta})^{2}$ by [GM91, Thm 5.1]. Define the Bowen–Margulis–Sullivan measure $m$ on $\mathcal{T}_{g}\Join\mathcal{T}_{g}\times\mathbb{R}$ for any $x\in\mathcal{T}_{g}$,
$m:=\beta(x,\xi,\eta)^{-1}\cdot\mu_{x}\times\mu_{x}\times\textrm{Leb}$
which is independent of $x$. As $m$ is $G$-invariant, it descends to the flow invariant measure $\mathbf{m}$ on $\mathcal{Q}^{1}\mathcal{M}_{g}$.
12.5. Hopf–Tsuji–Sullivan dichotomy for geodesic flows
We first give a brief introduction to Hopf decomposition of a flow dynamical system. This is classical fact, for which we follow the exposition from [Lin18] and refer the reader to the more references therein.
Let $\Omega$ be a locally compact and $\sigma$-compact Hausdorff topological space. A flow $\phi$ on $\Omega$ is a continuous map $\phi(t,\omega):\mathbb{R}\times\Omega\to\Omega$ such that $\phi(0,\omega)=\omega$ and $\phi(t,\phi(s,\omega))=\phi(t+s,\omega)$. Write $\phi^{t}(\omega)=\phi(t,\omega)$.
Assume that $\mathbf{m}$ is a $\phi$-invariant Borel measure on $\Omega$. By Hopf decomposition theorem, $\Omega$ admits a unique decomposition ($\mathbf{m}$-mod 0) of $\phi$-invariant Borel subsets called conservative and dissipative components, denoted by $\Omega_{C}$ and $\Omega_{D}$, as follows:
  • $\Omega_{C}$ admits no wandering set: there exists no Borel subset $W$ such that $\phi^{n}(W)\cap W=\emptyset$ for any $n\in\mathbb{Z}\setminus 0$.
  • $\Omega_{C}$ admits a measurable fundamental set $W$ so that $\cup_{n\in\mathbb{Z}}\phi^{n}(W)=\Omega$.
According to Poincaré’s recurrence theorem, $\mathbf{m}$-almost every point $\omega$ of $\Omega_{C}$ is positively recurrent: there exists $t_{n}\in\mathbb{R}\to\infty+$ so that $\phi^{t_{n}}(\omega)\to\omega$. On the other hand, by Hopf’s divergence theorem, $\mathbf{m}$-almost every point $\omega$ of $\Omega_{D}$ is positively divergent: for every compact $K\subset\Omega$, there exists $T$ so that $\phi^{t}(\omega)\notin K$ for all $t>T$.
Let us return to our setup: the geodesic flow $(\mathcal{Q}^{1}\mathcal{M}_{g},\mathcal{G}^{t},\textbf{m})$ is a measure preserving dynamical system, so the above Hopf decomposition holds. The point of Theorem 1.22 says that a dichotomy of Hopf decomposition holds according to whether the conical points have positive measure or not.
Conservativity $\Longleftrightarrow$ Full measure on conical points
If $(\mathcal{Q}^{1}\mathcal{M}_{g},\mathcal{G}^{t},\textbf{m})$ is conservative, then there is no m-positive wandering set $W$ so that $\{\mathcal{G}^{n}W:n\in\mathbb{Z}\}$ is pairwise disjoint. By Poincaré recurrence theorem, almost every point $q\in\mathcal{Q}^{1}\mathcal{M}_{g}$ is positively recurrent: $\mathcal{G}^{t_{n}}q\to q$ for an unbounded sequence $t_{n}>0$.
Let $\gamma$ be a Teichmüller geodesic ray ending at (the projective class of) $q^{+}$. As $\mathcal{Q}^{1}\mathcal{M}_{g}$ is a sphere bundle over $\mathcal{T}_{g}/G$, we see that $\gamma$ returns infinitely often to a compact subset of $\mathcal{T}_{g}/G$. In other words, $q^{+}\in\Lambda_{c}(G)$ is a conical point as defined in (6).
By the dichotomy of Theorem 1.14, the $\mu_{x}$-measure of conical points is either null or full, so $\mu_{x}$ is fully supported on the set of conical points and thus on the Myrberg set.
For the converse direction, by Hopf decomposition, $(\mathcal{Q}^{1}\mathcal{M}_{g},\mathcal{G}^{t},\mathbf{m})$ consists of the conservative and dissipative components, uniquely up to m-null sets. The flow lines in the dissipative component exit every compact sets, so if the conical points are of full measure, then the geodesic flow on $\mathcal{Q}^{1}\mathcal{M}_{g}$ must be conservative.
Conservativity $\Longleftrightarrow$ Ergodicity of flows
If $(\mathcal{G}^{t})_{t\in\mathbb{R}}$ on $\mathcal{Q}^{1}\mathcal{T}_{g}$ is ergodic without atoms, then it is of course conservative. The converse direction requires to run the Hopf argument.
To that end, we need a positive $\mathbf{m}$-integrable function $\Phi(x)$ on $\mathcal{Q}^{1}\mathcal{M}_{g}$. If $\mathbf{m}(\mathcal{Q}^{1}\mathcal{M}_{g})$ is finite, we can choose $\Phi=1$. If $G=\textrm{Mod}(\Sigma_{g})$, $\mathbf{m}$ is the Masur-Veech measure which is finite. The proof of ergodicity via Hopf argument is given in [Mas82a, Theorem 4].
In general, define $\Phi(x):=d_{Q}(x,\mathbf{o})$ for some $\mathbf{o}\in\mathcal{Q}^{1}\mathcal{M}_{g}$. The same argument as in [Nic89, Lemma 8.3.1] proves that $\Phi(x)$ is $\mathbf{m}$-integrable function and for some $c>0$ and all $x,y\in\mathcal{Q}^{1}\mathcal{M}_{g}$ with $d_{Q}(x,y)<1$
$\frac{|\Phi(x)-\Phi(y)|}{\Phi(y)}<c\cdot d_{Q}(x,y)$
where $d_{Q}$ is a metric on $\mathcal{Q}^{1}\mathcal{M}_{g}$ defined in (53).
Recall that $\mu_{x}$ is supported on the Myrberg set $\Lambda_{m}{(Go)}$, which consists of uniquely ergodic points. By [Mas80], any two geodesic rays $\alpha,\beta$ in $\mathcal{T}_{g}$ ending at a common unique ergodic point are asymptotic: there exists $a\in\mathbb{R}$ such that $d_{T}(\alpha(t+a),\beta(a))\to 0$ and thus for two $p,q\in\mathcal{Q}^{1}\mathcal{M}_{g}$ with the same vertical foliation,
$d_{Q}(\mathcal{G}^{t+a}(p),\mathcal{G}^{t}(q))\to 0,\quad t\to\infty.$
With these ingredients, the ergodicity of the geodesic flow and thus product measures $\mu_{x}\times\mu_{x}$ follows verbatim the Hopf’s argument (see [Nic89, Thm 8.3.2] for details). So Theorem 1.22 is proved.